### v2

# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

Abstract
The moduli space of stable Higgs bundles of degree 00 is equipped with the hyperkähler metric, called the Hitchin metric. On the locus where the spectral curves are smooth, there is the hyperkähler metric called the semi-flat metric, associated with the algebraic integrable systems with the Hitchin section. We prove the exponentially rapid decay of the difference between the Hitchin metric and the semi-flat metric along the ray $(E,t\theta)$ as $t\to\infty$.
MSC: 53C07, 58E15, 14D21, 81T13.
Keywords: harmonic bundle, moduli space of Higgs bundles, Hitchin metric, semi-flat metric
1 Introduction
1.1 The Hitchin metric and the semi-flat metric of the moduli space of Higgs bundles
1.1.1 Hitchin metric
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. In this paper, a Higgs bundle on $X$ is a holomorphic vector bundle $E$ with a holomorphic section $\theta$ of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$, where $K_{X}$ denotes the canonical bundle of $X$. The study of the moduli space of Higgs bundles was initiated by Hitchin in [14]. Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 00 and rank $n$ on $X$. It is naturally equipped with the integrable almost complex structure $I$. Moreover, $\mathcal{M}_{H}$ is equipped with the naturally defined hyperkähler structure, i.e., the Riemannian metric $g_{H,{\mathbb{R}}}$ and the three integrable almost complex structures $I$, $J$ and $K$ such that (i) the associated $2$-forms $\omega_{T}(\cdot,\cdot)=g_{H,{\mathbb{R}}}(\cdot,T\cdot)$ $(T=I,J,K)$ are closed, (ii) $IJK=-1$. Let $g_{H}$ denote the associated Kähler metric of $(\mathcal{M}_{H},I)$, i.e., $g_{H}=g_{H,{\mathbb{R}}}+\sqrt{-1}\omega_{I}$. The Riemannian metric $g_{H,{\mathbb{R}}}$ and the Kähler metric $g_{H}$ are called the Hitchin metric.
By the theorem of Hitchin and Simpson [14, 25], because $(E,\theta)\in\mathcal{M}_{H}$ is stable of degree 00, it carries a harmonic metric $h$, i.e., $h$ is a Hermitian metric of $E$ satisfying
$R(h)+[\theta,\theta^{\dagger}_{h}]=0.$
Here, $R(h)$ denotes the curvature of the Chern connection of $(E,h)$, and $\theta^{\dagger}_{h}$ denotes the adjoint of $\theta$ with respect to $h$. The induced Hermitian metric on $\mathop{\rm End}\nolimits(E)$ is also denoted by $h$. The tangent space $T_{(E,\theta)}\mathcal{M}$ is identified with the space of harmonic $1$-forms of the harmonic bundle $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$. The Hermitian metric $(g_{H})_{|(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms. (For example, see §4.2.3.) Namely, for harmonic $1$-forms $\tau$ of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$, let $[\tau]\in T_{(E,\theta)}\mathcal{M}_{H}$ denote the corresponding element, and then we have
$g_{H|(E,\theta)}([\tau],[\tau])=2\sqrt{-1}\int_{X}\Bigl{(}h(\tau^{1,0},\tau^{1,0})-h(\tau^{0,1},\tau^{0,1})\Bigr{)}=2\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\Bigl{(}\tau^{1,0}\wedge(\tau^{1,0})^{\dagger}_{h}-\tau^{0,1}\wedge(\tau^{0,1})^{\dagger}_{h}\Bigr{)}.$
It is attractive and challenging to study the asymptotic behaviour of $g_{H}$ around infinity of the non-compact space $\mathcal{M}_{H}$.
1.1.2 Semi-flat metric
We set $\mathcal{A}_{H}=\bigoplus_{j=1}^{n}H^{0}(X,K_{X}^{j})$. Let $\pi_{0}:T^{\ast}X\to X$ denote the projection. Let $\eta$ denote the tautological section of $\pi_{0}^{\ast}(K_{X})$. For any $a=(a_{j})\in\mathcal{A}_{H}$, we obtain the section $\xi_{a}=\eta^{n}+\sum_{j=1}^{n}(-1)a_{j}\eta^{n-j}$ of $\pi_{0}^{\ast}(K_{X}^{\otimes n})$. Let $\Sigma_{a}$ denote the 00-scheme of the section $\xi_{a}$. Such a curve is called a spectral curve. Let $\mathcal{A}_{H}^{\prime}\subset\mathcal{A}_{H}$ denote the Zariski open subset of $a\in\mathcal{A}_{H}$ such that $\Sigma_{a}$ are smooth. It is known that $\mathcal{A}_{H}^{\prime}$ is non-empty.
For any Higgs bundle $(E,\theta)$ of rank $n$, we obtain $\Phi_{H}(E,\theta)=(a_{j}(E,\theta)\,|\,j=1,\ldots,n)\in\mathcal{A}_{H}$ determined by
$\det(\eta\mathop{\rm id}\nolimits_{E}-\theta)=\eta^{n}+\sum_{j=1}^{n}(-1)^{j}a_{j}(E,\theta)\eta^{n-j}.$
It induces the morphism $\Phi_{H}:\mathcal{M}_{H}\to\mathcal{A}_{H}$ called the Hitchin fibration. There exists the Hitchin section of $\Phi_{H}$ depending on a root $K_{X}^{1/2}$ of $K_{X}$. For any $(E,\theta)\in\mathcal{M}_{H}$, the curve $\Sigma_{\theta}=\Sigma_{\Phi_{H}(E,\theta)}$ is called the spectral curve of $(E,\theta)$.
Let $\mathcal{M}_{H}^{\prime}$ denote the Zariski open subset of $(E,\theta)\in\mathcal{M}_{H}$ such that the spectral curves $\Sigma_{\theta}$ are smooth, i.e., $\mathcal{M}_{H}^{\prime}=\Phi_{H}^{-1}(\mathcal{A}_{H}^{\prime})$. We obtain the induced smooth fibration $\Phi_{H}^{\prime}:\mathcal{M}_{H}^{\prime}\to\mathcal{A}_{H}^{\prime}$ which is an algebraic integrable system in the sense of [8]. Each fiber $(\Phi_{H}^{\prime})^{-1}(a)$ $(a\in\mathcal{A}_{H}^{\prime})$ is naturally polarized. According to [8, Theorem 3.8], $\mathcal{M}_{H}^{\prime}$ carries a canonical hyperkähler structure, i.e., the Riemannian metric $g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}$ with the integrable almost complex structures $I,J^{\prime},K^{\prime}$ satisfying the axioms of hyperkähler structure. We set $g_{\mathop{\rm sf}\nolimits}=g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}+\sqrt{-1}\omega_{I}$. The Riemannian metric $g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}$ and the Kähler metric $g_{\mathop{\rm sf}\nolimits}$ are called the semi-flat metric.
Let us recall how $g_{\mathop{\rm sf}\nolimits}$ is constructed. For $a=(a_{i})\in\mathcal{A}_{H,n}^{\prime}$, let $U_{a}$ be a simply connected neighbourhood of $a$ in $\mathcal{A}_{H,n}^{\prime}$. We may regard $\Phi^{\prime}_{H|U_{a}}:\Phi_{H}^{-1}(U_{a})\to U_{a}$ as a principal torus bundle. By using a Hitchin section, we define the integrable connection of the principal torus bundle, which induces the decomposition of the tangent space $T_{(E,\theta)}\mathcal{M}_{H}^{\prime}$ into the horizontal direction and the vertical direction:
$T_{(E,\theta)}\mathcal{M}_{H}^{\prime}=(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}\oplus(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}.$
It is independent of the choice of a Hitchin section. Because the fiber $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is identified with a connected component of the Picard variety of $\Sigma_{\theta}$, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ is naturally identified with $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. The moduli space $\mathcal{M}_{H}^{\prime}$ is equipped with a complex symplectic structure. Because $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is Lagrangian, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ and $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ are mutually dual. Hence, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ is naturally identified with $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. The space $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ is identified with the space of harmonic $1$-forms on $\Sigma_{\theta}$. The Hermitian metric $(g_{\mathop{\rm sf}\nolimits})_{(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms on $\Sigma_{\theta}$. (See Proposition 4.6.) Namely, for harmonic $1$-forms $\tau$ on $\Sigma_{\theta}$, let $[\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ denote the corresponding element, and then we have
$g_{\mathop{\rm sf}\nolimits|(E,\theta)}([\tau],[\tau])=2\sqrt{-1}\int_{\Sigma_{\theta}}\Bigl{(}\tau^{1,0}\wedge\overline{\tau^{1,0}}-\tau^{0,1}\wedge\overline{\tau^{0,1}}\Bigr{)}.$
1.1.3 A weak version of a conjecture of Gaiotto-Moore-Neitzke
A conjecture of Gaiotto-Moore-Neitzke [9, 10] implies the following estimate for any $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ with respect to $(g_{\mathop{\rm sf}\nolimits})_{(E,t\theta)}$ as $t\to\infty$:
$(g_{H}-g_{\mathop{\rm sf}\nolimits})_{(E,t\theta)}=O(e^{-\beta t}).$ (1)
Here, $\beta$ is a positive constant depending on $(E,\theta)$. It is our purpose in this paper to prove this weak version of GMN-conjecture. Note that their original conjecture is much more precise.
A previous work was done by Fredrickson in [5] on the basis of her another work [6], the works of Mazzeo-Swoboda-Weiss-Witt [18, 19] and interesting ideas of Dumas-Neitzke [4]. More precisely, for each $(E,\theta)\in\mathcal{M}_{H}^{\prime}$, let $\mathop{\rm Crit}\nolimits(\theta)\subset\Sigma_{\theta}$ denote the set of the critical points of the projection $\pi:\Sigma_{\theta}\to X$. For each $Q\in\mathop{\rm Crit}\nolimits(\theta)$, let $r_{Q}$ denote the ramification index of $\pi$ at $Q$. We obtain the Zariski open subset $\mathcal{M}_{H}^{\prime\prime}$ of $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ such that $r_{Q}\leq 2$ for any $Q\in\mathop{\rm Crit}\nolimits(\theta)$. We have $\mathcal{M}^{\prime}_{H}=\mathcal{M}_{H}^{\prime\prime}$ if $n=2$ but $\mathcal{M}_{H}^{\prime}\neq\mathcal{M}_{H}^{\prime\prime}$ if $n\geq 3$, in general. The conjectural estimate (1) has been studied for any $(E,\theta)\in\mathcal{M}_{H}^{\prime\prime}$ in [5]. We should also note that a similar estimate has been studied by Fredrickson-Mazzeo-Swoboda-Weiss [7] in the case of parabolic Higgs bundles of rank $2$. Chen-Li [2] studied the rank $2$ but more singular case by using the methods in [5, 7]. In [15], Holdt studied the original approach of [9, 10] in the case that a Higgs bundle has a non-trivial parabolic structure. Note that because it is based on Fock-Goncharov coordinates, and it is not known whether the approach can be applied to the non-singular case.
We shall prove the conjectural estimate (1) for any $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ by using a different method. More precisely, let $s$ be the automorphism of the complex vector bundle $T\mathcal{M}_{H}^{\prime}$ determined by $g_{H}=g_{\mathop{\rm sf}\nolimits}\cdot s$.
Theorem 1.1 (Theorem 4.7)
For any compact subset $K\subset\mathcal{M}_{H}^{\prime}$, there exist positive constants $B_{i}$ $(i=1,2)$ such that
$|(s-\mathop{\rm id}\nolimits)_{(E,t\theta)}|_{g_{\mathop{\rm sf}\nolimits}}\leq B_{1}\exp(-B_{2}t)$
for any $(E,\theta)\in K$.
Remark 1.2
Regrettably, to the best of the author’s understanding, some more explanations or modifications are necessary for the arguments in [5, 7, 19]  [^1] The author thanks Laura Fredrickson for sincerely replying to the author’s questions. The author also thanks Rafe Mazzeo for answering to the inquiry regarding this remark..
Remark 1.3
In [22], we study the constant $B_{2}$ for the rank $2$ case.
1.2 Large-scale solutions of the Hitchin equation
1.2.1 Large-scale solutions
Let $(E,\theta)\in\mathcal{M}_{H}^{\prime}$. We fix a flat metric $h_{\det(E)}$ of $\det(E)$. For any $t>0$, we obtain harmonic metrics $h_{t}$ of $(E,\theta)$ such that $\det(h_{t})=h_{\det(E)}$. For large $t$, they are also called large-scale solutions of the Hitchin equation. To study the asymptotic behaviour of $g_{H}$ along the ray $(E,t\theta)$, it is basic to study the asymptotic behaviour of large-scale solutions $h_{t}$ as $t\to\infty$.
1.2.2 Limiting configuration
Let $\pi:\Sigma_{\theta}\to X$ denote the projection. According to Hitchin [14] and Beauville-Narasimhan-Ramanan [1], there exists a line bundle $L$ on $\Sigma_{\theta}$ with an isomorphism
$(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta),$
where $\pi_{\ast}(\eta)$ is the Higgs field induced by the tautological $1$-form $\eta$ on $T^{\ast}X$ as the push-forward of $\eta:L\to L\otimes\pi^{\ast}(K_{X})$. We set $D(\theta)=\pi(\mathop{\rm Crit}\nolimits(\theta))$ and $X^{\circ}=X\setminus D(\theta)$. For any subset $S\subset X$, we set $\Sigma_{\theta|K}:=\pi^{-1}(S)$. There exists a Hermitian metric $h_{L}$ of $L_{|\Sigma_{\theta|X^{\circ}}}$ such that the following holds.
  • $h_{L}$ is flat, i.e., $R(h_{L})=0$.
  • For each $Q\in\pi^{-1}(D(\theta))$, there exist a local holomorphic coordinate $\zeta_{Q}$ and a local frame $v_{Q}$ around $Q$ such that $\zeta_{Q}(Q)=0$ and $h_{L}(v_{Q},v_{Q})=|\zeta_{Q}|^{-r_{Q}+1}$.
By taking the push-forward of $h_{L}$ with respect to $\pi$, we obtain a Hermitian metric $h_{\infty}$ of $E_{|X^{\circ}}$. It is a decoupled harmonic metric in the sense
$R(h_{\infty})=[\theta,\theta^{\dagger}_{h_{\infty}}]=0.$
We also note that $\det(h_{\infty})$ induces a $C^{\infty}$-metric of $\det(E)$ on $X$ whose curvature is 00, and hence we may assume that $\det(h_{\infty})=h_{\det(E)}$ by adjusting a positive constant. It is called the limiting configuration. It was introduced by Mazzeo-Swoboda-Weiss-Witt [18] in the rank $2$ case, and by Fredrickson [6] in the higher rank case.
1.2.3 Convergence
In [18], Mazzeo-Swoboda-Weiss-Witt proved the exponentially rapid convergence of the sequence $h_{t}$ to $h_{\infty}$ as $t\to\infty$ locally on $X^{\circ}$. It is generalized to the higher rank case by Fredrickson in [6] under some mild assumption on the ramification of the projection $\Sigma_{\theta}\to X$ (see [23, Remark 1.2]), and by Szabo and the author in [23] more generally.
We recall a very rough strategy in [18]. For simplicity, we assume $\mathop{\rm tr}\nolimits(\theta)=0$, $\det(E)=\mathcal{O}_{X}$ and $h_{\det(E)}=1$. For any point $P\in D(\theta)$, let $(X_{P},z_{P})$ be a holomorphic coordinate neighbourhood around $P$ such that (i) $z_{P}(P)=0$, (ii) $X_{P}\simeq\{|z_{P}|<2\}$, (iii) $X_{P}\cap D(\theta)=\{P\}$. We set $(E_{P},\theta_{P}):=(E,\theta)_{|X_{P}}$. For any $0<r\leq 2$, we set $X_{P}(r)=\{|z_{P}|<r\}$.
Step 1
We construct harmonic metrics $h_{P,t}$ $(t>0)$ of $(E_{P},t\theta_{P})$ such that (i) $\det(h_{P,t})=h_{\det(E)|X_{P}}$, (ii) $h_{P,t}-h_{\infty}$ and their derivatives are $O(e^{-\epsilon t})$ on $X(1)\setminus X(1/2)$ for some $\epsilon>0$.
Step 2
We patch $h_{\infty}$ and $h_{P,t}$ $(P\in D(\theta))$ by using a partition of unity to construct a family of Hermitian metrics $\widetilde{h}_{t}$ of $E$ such that (i) $\det(\widetilde{h}_{t})=h_{\det(E)}$, (ii) $\widetilde{h}_{t}-h_{\infty}=O(e^{-\epsilon^{\prime}t})$ on $X\setminus\bigcup_{P\in D(\theta)}X_{P}(1/2)$, (iii) $\widetilde{h}_{t}-h_{P,t}=O(e^{-\epsilon^{\prime}t})$ on $X_{P}$, (iv) $R(\widetilde{h}_{t})+t^{2}[\theta,\theta^{\dagger}_{\widetilde{h}_{t}}]=O(e^{-\epsilon^{\prime}t})$ holds on $X$ for some $\epsilon^{\prime}>0$. In particular, we have the exponentially rapid convergence of the sequence $\widetilde{h}_{t}$ to $h_{\infty}$ locally on $X^{\circ}$.
Step 3
We show $\widetilde{h}_{t}-h_{t}=O(e^{-\delta t})$ for some $\delta>0$ on $X$.
In [18], the so called fiducial solutions play important roles for the construction of $h_{P,t}$ in Step 1. In the rank $2$ case, the Higgs bundle $(E_{P},\theta_{P})$ is essentially uniquely determined, and it is highly symmetric. Namely, we may normalize the coordinate $z_{P}$ such that $\det(\theta_{P})=-z_{P}(dz_{P})^{2}$, and there exists a frame $v_{1},v_{2}$ of $E_{P}$ such that
$\theta_{P}(v_{1})=v_{2}\,dz_{P},\quad\theta_{P}(v_{2})=v_{1}\,z_{P}\,dz_{P},\quad v_{1}\wedge v_{2}=1.$
There exists a harmonic metric $h_{P,t}$ of $(E_{P},\theta_{P})$ such that (i) $h_{P,t}(v_{i},v_{i})$ depends only on $|z_{P}|$, (ii) $h_{P,t}(v_{i},v_{j})=0$ $(i\neq j)$, (iii) $\det(h_{P,t})=1$. We can find such a harmonic metric by solving a non-linear ordinary differential equation, which is a special case of Painlevé III equation, or by rescaling a wild harmonic bundle on ${\mathbb{C}}$. The solutions are called fiducial solutions. We can adjust $h_{P,t}$ so that the conditions in Step 1 are satisfied. This method can be also useful in the higher rank case if $r_{Q}\leq 2$ for any $Q\in\mathop{\rm Crit}\nolimits(\theta)$, and it was efficiently used in [5, 6]. However, it is not clear whether the method is so useful for the study of large-scale solutions in the higher rank case without any assumption on the ramification.
1.2.4 Locally defined symmetric pairings and a new method
In [23], we introduced another method for the construction of $h_{P,t}$ for $(E,t\theta)$ as in Step 1, which is available in a more general case. There exists a non-degenerate symmetric pairing $C_{P}$ of $E_{P}$ such that the following holds (see [23]):
  • $C_{P}$ is a non-degenerate symmetric pairing of $(E_{P},\theta_{P})$, i.e., $\theta_{P}$ is self-adjoint with respect to $C_{P}$.
  • $C_{P|X_{P}\setminus\{P\}}$ is compatible with $h_{\infty|X_{P}\setminus\{P\}}$ in the sense that the induced morphism $\Psi_{C_{P}}:E_{P}\to E_{P}^{\lor}$ is isometric with respect to $h_{\infty}$ and the induced metric $h_{\infty}^{\lor}$ on $X_{P}\setminus\{P\}$.
By the theorem of Donaldson on the Dirichlet problem for harmonic metrics [3], there exists a harmonic metric $h_{P,t}$ such that $h_{P,t|\partial X_{P}}=h_{\infty|\partial X_{P}}$. It is easy to observe that $C_{P}$ and $h_{P,t}$ are compatible (see [17]). By an estimate in [23] on the basis of a variant of Simpson’s main estimate in [20], which goes back to [26], the conditions in Step 1 are automatically satisfied.
We remark that the argument for Step 3 in [23] is also different from those in [6, 18]. In the latter, after the detailed analysis of the fiducial solutions, they closely study the family of Green operators associated with $(E,t\theta,\widetilde{h}_{t})$, and eventually they proved for sufficiently large $t$, that there exists $h_{t}$ in an exponentially small neighbourhood of $\widetilde{h}_{t}$ in the space of Hermitian metrics by using the inverse function theorem. The argument in [23] is based on the maximum principle of subharmonic functions.
The study [6, 18] were motivated by the conjectural estimate (1), and actually applied in [5, 19]. We would like to check whether our result and method in [23] are useful enough for the study of (1), that is our original purpose in this paper.
We note that our route to (1) is apparently different from those in [5, 19], where they first show that $g_{\mathop{\rm sf}\nolimits}$ is identified with the natural $L^{2}$-metric $g_{L^{2}}(\mathcal{M}^{\prime}_{\infty})$ on the moduli space $\mathcal{M}_{\infty}^{\prime}$ of Higgs bundles with the limiting configurations, and they show that $g_{H}-g_{L^{2}}(\mathcal{M}^{\prime}_{\infty})=O(e^{-\beta t})$ along the ray $(E,t\theta)$ as $t\to\infty$. (See [5] for more precise explanations.) We shall compare $g_{H}$ and $g_{\mathop{\rm sf}\nolimits}$ more directly. An outline is explained in the next subsection.
1.3 Outline
1.3.1 The horizontal direction and the vertical direction
Let $(E,\theta)\in\mathcal{M}_{H}^{\prime}$. Let $\mathop{\rm Def}\nolimits(E,\theta)$ be the complex of sheaves
[…]
Here, the first term sits in the degree 00. The hypercohomology group of the complex $\mathop{\rm Def}\nolimits(E,\theta)$ is denoted as $H^{\ast}(X,\mathop{\rm Def}\nolimits(E,\theta))$. It is well known that there exists the following natural isomorphism (see §2.2 for an explanation):
$T_{(E,\theta)}\mathcal{M}_{H}^{\prime}\simeq H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta)).$
We shall identify them by the isomorphism. There exist natural isomorphisms (see Proposition 2.4):
$\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))\simeq\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}},\quad\quad\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))\simeq\pi_{\ast}K_{\Sigma_{\theta}}.$
We obtain the following exact sequence:
[…]
It is easy to observe that $\iota^{{\tt v}}$ induces the isomorphism of $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and the tangent space of $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))\subset\mathcal{M}^{\prime}_{H}$, i.e., $(T_{(E,\theta)}\mathcal{M}^{\prime}_{H})^{{\tt v}}$. There uniquely exists the splitting $\iota^{{\tt h}}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ such that the image $\iota^{{\tt h}}$ equals $(T_{(E,\theta)}\mathcal{M}^{\prime}_{H})^{{\tt h}}$.
1.3.2 $C^{\infty}$-representatives of the horizontal infinitesimal deformations
Let $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ denote the complex obtained as the space of global sections of the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta)$. There exists the line bundle $L$ on $\Sigma_{\theta}$ with an isomorphism $(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta)$. For any $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, the multiplication of $\nu$ on $L$ induces $F_{\nu}\in H^{0}(X,\mathop{\rm End}\nolimits(E)\otimes K_{X}(D))$. We would like to modify $F_{\nu}$ to a $1$-cocycle $\tau$ in $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ such that the cohomology class $[\tau]$ equals $\iota^{{\tt h}}(\nu)$. We remark that any $1$-cocycle $\tau$ in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ naturally induces a holomorphic section $\rho(\tau)$ of $\pi_{\ast}(K_{\Sigma_{\theta}})$ such that $\Upsilon_{1}([\tau])=\rho(\tau)$. (See §2.2.6.)
Let $\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})$ (resp. $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P})$) denote the sheaf of holomorphic endomorphisms of $E_{P}$ which are symmetric (resp. anti-symmetric) with respect to $C_{P}$. We obtain the following subcomplex $\mathop{\rm Def}\nolimits(E_{P},\theta_{P};C_{P})$ of $\mathop{\rm Def}\nolimits(E_{P},\theta_{P})$:
$\begin{CD}\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P})@>{\mathop{\rm ad}\nolimits\theta_{P}}>{}>\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})\otimes K_{X_{P}}.\end{CD}$
Let $A^{\bullet}(\mathop{\rm Def}\nolimits(E_{P},\theta_{P};C_{P}))$ denote the complex obtained as the space of the global sections of the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E_{P},\theta_{P};C_{P})$. Then, we can obtain the following sufficient condition for a representative of $\iota^{{\tt h}}(\nu)$ in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$.
Proposition 1.4 (Corollary 3.32)
A $1$-cocycle $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfies $[\tau]=\iota^{{\tt h}}(\nu)$ if the following holds.
  • There exist a relatively compact open neighbourhood $X_{P}^{\prime}$ of $P$ in $X_{P}$ $(P\in D(\theta))$ such that $\tau_{|X\setminus\bigcup X_{P}^{\prime}}=F_{\nu|X\setminus\bigcup X_{P}^{\prime}}$.
  • $\tau_{|X_{P}}\in A^{1}(\mathop{\rm Def}\nolimits(E_{P},\theta_{P};C_{P}))$.
1.3.3 Harmonic representatives for horizontal infinitesimal deformations
Let $h_{t}$ be the harmonic metric of $(E,t\theta)$ such that $\det(h_{t})=h_{\det(E)}$. Let $g_{X}$ be a conformal metric of $X$. Let ¯$\overline{\partial}_{E}$ denote the Cauchy operator of $E$. The induced Cauchy operator of $\mathop{\rm End}\nolimits(E)$ is also denoted by ¯$\overline{\partial}_{E}$. Let $\iota^{{\tt h}}_{t}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$ denote the splitting morphism for $(E,t\theta)$, where we identify $\Sigma_{t\theta}$ and $\Sigma_{\theta}$ in a natural way. To study the harmonic representative ${\tt H}(\nu,t)$ of $\iota^{{\tt h}}_{t}(\nu)$, we construct a $1$-cocycle ${\tt H}^{\prime}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that the following holds.
  • (i)
    $\rho({\tt H}^{\prime}(\nu,t))=\pi_{\ast}(\nu)$.
  • (ii)
    ${\tt H}^{\prime}(\nu,t)$ satisfies the conditions in Proposition 1.4. In particular, $[{\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$.
  • (iii)
    (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t)=O(e^{-\beta t})\|\nu\|$ for some $\beta>0$, where $\|\nu\|$ denotes the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$.
  • (iv)
    $\mathop{\rm tr}\nolimits\bigl{(}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t)\bigr{)}=0$.
Let $P\in D(\theta)$. There exists a holomorphic function $\alpha_{P}$ on $\Sigma_{\theta|X_{P}}$ such that $d\alpha_{P}=\nu_{|\Sigma_{\theta|X_{P}}}$ and that $\alpha_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. The multiplication of $\alpha_{P}$ on $L_{|\Sigma_{\theta|X_{P}}}$ induces an endomorphism $F_{\alpha_{P}}$ of $E_{P}$ such that $[\theta,F_{\alpha_{P}}]=0$. Let $h_{P,t}$ be the harmonic metric of $(E_{P},t\theta_{P})$ such that $h_{P,t|\partial X_{P}}=h_{\infty|\partial X_{P}}$. Let $\nabla_{h_{P,t}}=\overline{\partial}_{E}+\partial_{E_{P},h_{P,t}}$ denote the Chern connection of $(E,h_{P,t})$. Let $\theta^{\dagger}_{h_{P,t}}$ denote the adjoint of $\theta$ with respect to $h_{P,t}$. Then, it is easy to see that
${\tt H}_{P}(\nu,t)=(\partial_{E,h_{P,t}}+\mathop{\rm ad}\nolimits t\theta^{\dagger}_{h_{P,t}})F_{\alpha_{P}}\in A^{1}(\mathop{\rm Def}\nolimits(E_{P},t\theta_{P}))$
is a harmonic $1$-form for $(E_{P},t\theta_{P},h_{P,t})$. (See Proposition 2.26.) Moreover, it satisfies $\rho({\tt H}_{P}(\nu,t))=\pi_{\ast}(\nu)_{|X_{P}}$ and ${\tt H}_{P}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E_{P},t\theta_{P};C_{P}))$.
By the asymptotic orthogonality in [20], we obtain ${\tt H}_{P}(\nu,t)-F_{\nu}=O(e^{-\beta t})\|\nu\|$ for some $\beta>0$ on $X_{P}(1)\setminus X_{P}(1/2)$. There exists a unique $C^{\infty}$-section $\sigma_{P}(\nu,t)$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C_{P})$ on $X_{P}(1)\setminus X_{P}(1/2)$ such that ${\tt H}_{P}(\nu,t)-F_{\nu}=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\sigma_{P}(\nu,t)$ on $X_{P}(1)\setminus X_{P}(1/2)$ and that $|\sigma_{P}(\nu,t)|_{h_{P,t}}=O(e^{-\epsilon t})\|\nu\|$ for some $\epsilon>0$. We patch $F_{\nu}$ and ${\tt H}_{P}(\nu,t)$ $(P\in D(\theta))$ by using $\sigma_{P}(\nu,t)$ and a partition of unity, and we obtain a $1$-cocycle ${\tt H}^{\prime}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfying the conditions (i–iv). Note that we can check the condition (iii) by using $h_{P,t}-h_{t}=O(e^{-\beta t})$ on $X_{P}(1)\setminus X_{P}(1/2)$ for some $\beta>0$ in [23].
For each $t>0$, there uniquely exists $\gamma(\nu,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ such that $\mathop{\rm tr}\nolimits\gamma(\nu,t)=0$ and
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
Then, ${\tt H}(\nu,t)={\tt H}^{\prime}(\nu,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)$ are the harmonic representatives of $\iota^{{\tt h}}_{t}(\nu)$. By using the estimate for the first eigenvalues of the Laplacian operators (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.42) and the regularity for solutions of the Poisson equation (see Proposition 2.36), we obtain $|\gamma(\nu,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)|_{h_{t},g_{X}}=O(e^{-\epsilon t})$ for some $\epsilon>0$.
1.3.4 Estimate of the pairings in the horizontal direction
As the restriction of the estimate (1) in the horizontal direction, we would like to prove that
$\Bigl{|}g_{H}(\iota^{{\tt h}}_{t}(\nu),\iota^{{\tt h}}_{t}(\nu))-g_{\mathop{\rm sf}\nolimits}(\iota^{{\tt h}}_{t}(\nu),\iota^{{\tt h}}_{t}(\nu))\Bigr{|}=O(e^{-\epsilon t})\|\nu\|^{2}$ (2)
for some $\epsilon>0$. It is equivalent to
$2\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt H}(\nu,t)^{1,0},{\tt H}(\nu,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt H}(\nu,t)^{0,1},{\tt H}(\nu,t)^{0,1}\bigr{)}\Bigr{)}-2\sqrt{-1}\int_{\Sigma_{\theta}}\nu\wedge\overline{\nu}=O\bigl{(}e^{-\epsilon t}\bigr{)}\|\nu\|^{2}.$ (3)
Let us explain an outline. We set $X_{1}=X\setminus\bigcup_{P}X_{P}(1)$. We obtain (3) from the following two estimates:
$2\sqrt{-1}\int_{X_{1}}h_{t}\bigl{(}F_{\nu},F_{\nu}\bigr{)}-2\sqrt{-1}\int_{\Sigma_{\theta|X_{1}}}\nu\wedge\overline{\nu}=O\bigl{(}e^{-\epsilon t}\bigr{)}\|\nu\|^{2},$ (4)
$2\sqrt{-1}\int_{X_{P}(1)}\Bigl{(}h_{P,t}\bigl{(}{\tt H}_{P}(\nu,t)^{1,0},{\tt H}_{P}(\nu,t)^{1,0}\bigr{)}-h_{P,t}\bigl{(}{\tt H}_{P}(\nu,t)^{0,1},{\tt H}_{P}(\nu,t)^{0,1}\bigr{)}\Bigr{)}-2\sqrt{-1}\int_{\Sigma_{\theta|X_{P}(1)}}\nu\wedge\overline{\nu}\\ =O\bigl{(}e^{-\epsilon t}\bigr{)}\|\nu\|^{2}.$ (5)
Let $F_{\overline{\nu}}$ denote the $C^{\infty}$-section of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X^{\circ}$ induced by the multiplication of $\overline{\nu}$ on $L$. We have
$\int_{X_{1}}h_{t}\bigl{(}F_{\nu},F_{\nu}\bigr{)}-\int_{\Sigma_{\theta|X_{1}}}\nu\wedge\overline{\nu}=\int_{X_{1}}\mathop{\rm tr}\nolimits\Bigl{(}F_{\nu}\cdot\bigl{(}(F_{\nu})^{\dagger}_{h_{t}}-F_{\overline{\nu}}\bigr{)}\Bigr{)}.$
By using the asymptotic orthogonality of $h_{t}$ on $X_{1}$ [20], we obtain
$|F_{\nu}|_{h_{t},g_{X}}\leq B_{0}\|\nu\|,\quad\quad\bigl{|}(F_{\nu})^{\dagger}_{h_{t}}-F_{\overline{\nu}}\bigr{|}_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\nu\|$
for some positive constants $B_{i}$ $(i=0,1,2)$ on $X_{1}$. Hence, we obtain the estimate (4).
To obtain the estimate (5), we use a variant of the key idea due to Dumas-Neitzke [4] efficiently applied by Fredrickson in [5]. By applying the Stokes formula to harmonic $1$-forms, we obtain the following equality (Proposition 2.27):
$\sqrt{-1}\int_{X_{P}(1)}\Bigl{(}h_{P,t}\bigl{(}{\tt H}_{P}(\nu,t)^{1,0},{\tt H}_{P}(\nu,t)^{1,0}\bigr{)}-h_{P,t}\bigl{(}{\tt H}_{P}(\nu,t)^{0,1},{\tt H}_{P}(\nu,t)^{0,1}\bigr{)}\Bigr{)}=\\ \sqrt{-1}\int_{\partial X_{P}(1)}\Bigl{(}h_{P,t}\bigl{(}F_{\alpha_{P}},{\tt H}_{P}(\nu,t)^{1,0}\bigr{)}-h_{P,t}\bigl{(}F_{\alpha_{P}},{\tt H}_{P}(\nu,t)^{0,1}\bigr{)}\Bigr{)}=\\ \sqrt{-1}\int_{\partial X_{P}(1)}\mathop{\rm tr}\nolimits\Bigl{(}F_{\alpha_{P}}\cdot\bigl{(}{\tt H}_{P}(\nu,t)^{1,0}\bigr{)}^{\dagger}_{h_{P,t}}-F_{\alpha_{P}}\cdot\bigl{(}{\tt H}_{P}(\nu,t)^{0,1}\bigr{)}^{\dagger}_{h_{P,t}}\Bigr{)}.$ (6)
We also have
$\sqrt{-1}\int_{\Sigma_{\theta|X_{P}(1)}}\nu\wedge\overline{\nu}=\sqrt{-1}\int_{\partial\Sigma_{\theta|X_{P}(1)}}\alpha_{P}\wedge\overline{\nu}=\sqrt{-1}\int_{\partial X_{P}}\mathop{\rm tr}\nolimits\bigl{(}F_{\alpha_{P}}\cdot F_{\overline{\nu}}\bigr{)}.$
On a neighbourhood of $\partial X_{P}(1)$ we obtain
$|F_{\alpha_{P}}|_{h_{P,t}}\leq B_{0}\|\nu\|,\quad\quad\bigl{|}({\tt H}_{P}(\nu,t)^{1,0})^{\dagger}_{h_{P,t}}-F_{\overline{\nu}}\bigr{|}_{h_{P,t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\nu\|,$
$|{\tt H}_{P}(\nu,t)^{0,1}|_{h_{P,t}g_{X}}\leq B_{3}\exp(-B_{4}t)\|\nu\|.$
Thus, we obtain the estimate (5).
Remark 1.5
A special case of Proposition 2.27 seems to be closely related with the exactness in [4, §9] though the author has not studied the precise relation.
1.3.5 Harmonic representatives of vertical infinitesimal deformations
Let $\tau\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. It is uniquely expressed by a harmonic $(0,1)$-form on $\Sigma_{\theta}$, which is also denoted by $\tau$. Under the natural identification $\Sigma_{t\theta}$ and $\Sigma_{\theta}$, we have $\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}})=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,t\theta))$. Hence, we obtain $\iota_{t}^{{\tt v}}(\tau)\in H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$. There exists the harmonic representative ${\tt V}(\tau,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ of $\iota_{t}^{{\tt v}}(\tau)$.
The multiplication of $\tau$ on $L$ induces a $C^{\infty}$-section $F_{\tau}$ of End($\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}_{X^{\circ}}$. It is helpful to prove that there exists $B_{1},B_{2}>0$ such that the following holds on $X\setminus\bigcup_{P\in D(\theta)}X_{P}(1/2)$:
$\bigl{|}{\tt V}(\tau,t)-F_{\tau}\bigr{|}_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|.$ (7)
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exist the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
We have the holomorphic vector bundles $E^{\dagger}_{t}=(E,\partial_{E,h_{t}})$ with the Higgs field $\theta_{h_{t}}^{\dagger}$ on $X^{\dagger}$. The spectral curve of $(E^{\dagger}_{t},t\theta_{h_{t}}^{\dagger})$ equals $(\Sigma_{t\theta})^{\dagger}\subset(T^{\ast}X)^{\dagger}=T^{\ast}(X^{\dagger})$. Let $\pi^{\dagger}_{t}:(\Sigma_{t\theta})^{\dagger}\to X^{\dagger}$ denote the projection. There exists a holomorphic line bundle $L^{\dagger}_{t}$ on $(\Sigma_{t\theta})^{\dagger}$ with an isomorphism $(\pi_{t}^{\dagger})_{\ast}L^{\dagger}_{t}\simeq E^{\dagger}_{t}$.
Let $P\in D(\theta)$. There exists an anti-holomorphic function $\beta_{P}$ on $\Sigma_{\theta|X(P)}$ such that ¯$\overline{\partial}\beta_{P}=\tau_{|\Sigma_{\theta|X_{P}}}$ and $\beta_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. We may regard $\beta_{P}$ as a holomorphic function on $(\Sigma^{\dagger}_{t\theta})_{|X_{P}^{\dagger}}$. The multiplication of $\beta_{P}$ on $L^{\dagger}_{t}$ induces the endomorphism $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}$ of $E^{\dagger}_{t}$. It satisfies
$\bigl{(}\partial_{E,h_{t}}+\mathop{\rm ad}\nolimits t\theta^{\dagger}_{h_{t}}\bigr{)}F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}=0.$
We obtain the following harmonic $1$-form for $(E,t\theta,h_{t})$ on $X_{P}$:
${\tt V}_{P}(\tau,t):=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}.$
By the asymptotic orthogonality in [20], we obtain ${\tt V}_{P}(\tau,t)-F_{\tau}=O\bigl{(}e^{-\beta t}\bigr{)}$ on $X_{P}(1)\setminus X_{P}(1/2)$ for some $\beta>0$.
Remark 1.6
We have $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}=(F_{\overline{\beta}_{P}})^{\dagger}_{h_{t}}$, where $F_{\overline{\beta}_{P}}$ is the holomorphic endomorphism of $(E,\theta)_{|X_{P}}$ induced by the holomorphic function $\beta_{P}$ on $\Sigma_{\theta|X_{P}}$.
For each $P\in D(\theta)$, let $\chi_{P}:X\to\{0\leq a\leq 1\}$ be a $C^{\infty}$-function such that $\chi_{P}=1$ on $X_{P}(1/2)$ and $\chi_{P}=0$ on $X\setminus X_{P}(1)$. We define the $C^{\infty}$ $(0,1)$-form $\tau^{\prime}$ on $\Sigma_{\theta}$ as follows:
$\tau^{\prime}=\tau-\sum_{P\in D(\theta)}d\bigl{(}\chi_{P}\beta_{P}\bigr{)}.$
We obtain the $C^{\infty}$-section $F_{\tau^{\prime}}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ such that $[F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}(\tau)$. We set
${\tt V}^{\prime}(\tau,t)=F_{\tau^{\prime}}+\sum_{P\in D(\theta)}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\bigl{(}\chi_{P}\cdot F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{P}}\bigr{)}.$
Then, ${\tt V}^{\prime}(\tau,t)$ satisfies (7) for some $B_{i}>0$. We have (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)=O\bigl{(}e^{-\epsilon t}\bigr{)}\|\tau\|$ on $X$ for some $\epsilon>0$. Moreover, $\mathop{\rm tr}\nolimits\bigl{(}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr{)}=0$ holds. Then, for each $t>0$, there uniquely exists $\gamma(\tau,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ such that $\mathop{\rm tr}\nolimits\gamma(\tau,t)=0$ and
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
Again, by using the estimate for the first eigenvalues of the Laplacian operators (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.42) and the regularity for solutions of the Poisson equation (see Proposition 2.36), we obtain $|\gamma(\tau,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ for some $B_{i}>0$. Because ${\tt V}(\tau,t)={\tt V}^{\prime}(\tau,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)$, we obtain the estimate (7) for ${\tt V}(\tau,t)$.
1.3.6 Estimate of pairings with vertical infinitesimal deformations
To prove the estimate (1), we would like to prove the following estimates for $\tau\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$:
$\Bigl{|}g_{H}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt v}}_{t}(\tau))-g_{\mathop{\rm sf}\nolimits}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt v}}_{t}(\tau))\Bigr{|}=O(e^{-\epsilon t})\|\tau\|^{2},$ (8)
$\Bigl{|}g_{H}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt h}}_{t}(\nu))-g_{\mathop{\rm sf}\nolimits}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt h}}_{t}(\nu))\Bigr{|}=O(e^{-\epsilon t})\|\tau\|\cdot\|\nu\|$ (9)
for some $\epsilon>0$. They are equivalent to the following estimates, respectively:
$2\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1}\bigr{)}\Bigr{)}-(-2\sqrt{-1})\int_{\Sigma_{\theta}}\tau\overline{\tau}=O(e^{-\epsilon t})\|\tau\|^{2},$ (10)
$2\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt H}(\nu,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt H}(\nu,t)^{0,1}\bigr{)}\Bigr{)}=O(e^{-\epsilon t})\|\tau\|\cdot\|\nu\|.$ (11)
Let us explain an outline of the proof of the estimate (10). Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we obtain
$\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1}\bigr{)}\Bigr{)}=-\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}F_{\tau^{\prime}},{\tt V}(\tau,t)^{0,1}\bigr{)}\Bigr{)}\\ =-\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\Bigl{(}F_{\tau^{\prime}}\cdot\bigl{(}{\tt V}(\tau,t)^{0,1}\bigr{)}^{\dagger}_{h_{t}}\Bigr{)}.$ (12)
We also have
$\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}=\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl{(}F_{\tau^{\prime}}F_{\overline{\tau}}\bigr{)}.$
By the estimate (7) and the asymptotic orthogonality in [20], the following holds on $X\setminus\bigcup X_{P}(1/2)$ for some $B_{3},B_{4}>0$:
$\bigl{|}\bigl{(}{\tt V}(\tau,t)^{0,1}\bigr{)}^{\dagger}_{h_{t}}-F_{\overline{\tau}}\bigr{|}_{h_{t},g_{X}}\leq B_{3}\exp(-B_{4}t)\|\tau\|^{2}.$
Then, we obtain the estimate (10). We can obtain the estimate (11) similarly.
1.3.6.1 Acknowledgements
This study grew out of my efforts to understand the attractive works [4, 5, 19]. I hope that this would be useful in the further study of this subject. I thank Qiongling Li and Szilard Szabo for discussions and collaborations. This study is partially based on the joint works [17, 23]. I am grateful to Gao Chen, Laura Fredrickson, Nianzi Li, Rafe Mazzeo and Siqi He for some discussions. In particular, I appreciate Fredrickson and Mazzeo for sincerely replying to my inquiries regarding Remark 1.2. I am grateful to Fredrickson, Mazzeo and Szabo for helpful comments to this manuscript. I thank Hitoshi Fujioka for discussions. I thank Yoshifumi Tsuchimoto and Akira Ishii for their constant encouragements.
I gave talks on the related topics in the conferences “The Hitchin system, Langlands duality and mirror symmetry” in April 2023 in ICMAT, and “Minimal Surfaces in Symmetric Spaces” in May 2023 in IMAG. The preparation of the talks at these conferences were useful in this research. I thank the organizers for giving me the opportunities of the talks. I gave a series of lectures on this topic in University of Science and Technology of China in October 2024. I thank Bin Xu for the opportunity of lectures, and all the participants for many stimulating questions.
I am partially supported by the Grant-in-Aid for Scientific Research (A) (No. 21H04429), the Grant-in-Aid for Scientific Research (A) (No. 22H00094), the Grant-in-Aid for Scientific Research (A) (No. 23H00083), and the Grant-in-Aid for Scientific Research (C) (No. 20K03609), Japan Society for the Promotion of Science. I am also partially supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
2 Preliminaries
2.1 Some notation
Let $X$ be any Riemann surface. We use the notation $K_{X}$ to denote the sheaf of holomorphic $1$-forms in this section. The cotangent bundle is denoted by $T^{\ast}X$. Let $\pi_{0}:T^{\ast}X\to X$ denote the projection. Let $\eta$ denote the tautological section of $\pi_{0}^{\ast}K_{X}$, i.e., $\eta$ is the section induced by the diagonal morphism $T^{\ast}X\to T^{\ast}X\times_{X}T^{\ast}X$.
Let $(E,\theta)$ be a Higgs bundle on $X$, i.e., $E$ denotes a holomorphic vector bundle on $X$, and $\theta$ denotes a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$. The spectral curve of the Higgs bundle $(E,\theta)$ is denoted by $\Sigma_{\theta}$, i.e., $\Sigma_{\theta}$ denotes the 00-scheme of the section $\det(\eta\mathop{\rm id}\nolimits_{E}-\theta)$ of $\pi_{0}^{\ast}\bigl{(}K_{X}^{\otimes\mathop{\rm rank}\nolimits(E)}\bigr{)}$. For any subset $U\subset X$, we set $\Sigma_{\theta|U}=\Sigma_{\theta}\times_{X}U$. We set
$D(\theta):=\bigl{\{}P\in X\,\big{|}\,|T_{P}^{\ast}X\cap\Sigma_{\theta}|<\mathop{\rm rank}\nolimits(E)\bigr{\}}.$
The Higgs bundle is called regular semisimple if $D(\theta)=\emptyset$. The Higgs bundle is called generically regular semisimple if $D(\theta)\neq X$. In this case, $D(\theta)$ is discrete in $X$. We set $X^{\circ}:=X\setminus D(\theta)$. The induced Higgs field $\mathop{\rm ad}\nolimits(\theta)$ on $\mathop{\rm End}\nolimits(E)$ is often denoted by $\theta$ if there is no risk of confusion.
2.1.1 Local decomposition
Suppose that $(E,\theta)$ is generically regular semisimple. For any $P\in X$, let $(X_{P},z_{P})$ denote a holomorphic coordinate neighbourhood of $P$ in $X$ such that $z_{P}(P)=0$. We set $X_{P}^{\circ}=X_{P}\setminus\{P\}$. We assume that $X_{P}$ is simply connected and that $X_{P}^{\circ}\cap D(\theta)=\emptyset$. There exists the decomposition into the connected components
$\Sigma_{\theta|X_{P}}=\coprod_{Q\in\Sigma_{\theta|P}}(\Sigma_{\theta})_{Q},$ (13)
where $(\Sigma_{\theta})_{Q}\cap T_{P}^{\ast}X=\{Q\}$. There exists the decomposition
$(E,\theta)_{|X_{P}}=\bigoplus_{Q\in T_{P}^{\ast}X\cap\Sigma_{\theta}}(E_{Q},\theta_{Q})$ (14)
such that $\Sigma_{\theta_{Q}}=(\Sigma_{\theta})_{Q}$. If $P\in X^{\circ}$, we have $\mathop{\rm rank}\nolimits E_{Q}=1$ for any $Q\in\Sigma_{\theta|P}$.
2.1.2 Filtrations
There exists a decreasing filtration $\mathcal{F}^{j}(E_{Q|P})$ $(j=0,\ldots,r_{Q}-1)$ obtained as the images of
$\pi_{\ast}\bigl{(}L_{|(\Sigma_{\theta})_{Q}}(-jQ)\bigr{)}_{|P}\to\pi_{\ast}(L_{|(\Sigma_{\theta})_{Q}})_{|P}.$
We also set $\mathcal{F}^{r_{Q}}(E_{Q|P})=0$. We set $\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{j}(E_{Q|P})=\mathcal{F}^{j}(E_{Q|P})\big{/}\mathcal{F}^{j+1}(E_{Q|P})$ for $j=0,\ldots,r_{Q}-1$. We have $\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{j}(E_{Q|P})=1$.
2.2 Deformation complexes of Higgs bundles
2.2.1 Deformation complexes of Higgs bundles
The deformation complex $\mathop{\rm Def}\nolimits(E,\theta)$ of $(E,\theta)$ is the following complex of sheaves of holomorphic sections:
$\mathop{\rm End}\nolimits(E)\stackrel{{\scriptstyle\mathop{\rm ad}\nolimits\theta}}{{\longrightarrow}}\mathop{\rm End}\nolimits(E)\otimes K_{X}.$
Here, the first term sits in the degree 00.
When the Higgs bundle $(E,\theta)$ is equipped with a non-degenerate symmetric pairing $C$, let $\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)$ (resp. $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$) denote the sheaf of sections $f$ of $\mathop{\rm End}\nolimits(E)$ such that $f$ is self-adjoint (resp. anti-self-adjoint) with respect to $C$. We obtain the subcomplex $\mathop{\rm Def}\nolimits(E,\theta;C)$ of $\mathop{\rm Def}\nolimits(E,\theta)$ given by
$\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\stackrel{{\scriptstyle\mathop{\rm ad}\nolimits\theta}}{{\longrightarrow}}\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}.$
2.2.2 The morphism $\mathop{\rm ad}\nolimits(\theta)$ in the case where the spectral curve is smooth
Suppose that $(E,\theta)$ is generically regular semisimple and that $\Sigma_{\theta}$ is smooth.
Lemma 2.1
The following holds.
  • The kernel and the cokernel of the morphism $\mathop{\rm ad}\nolimits(\theta):\mathop{\rm End}\nolimits(E)\to\mathop{\rm End}\nolimits(E)\otimes K_{X}$ are locally free $\mathcal{O}_{X}$-modules of rank $\mathop{\rm rank}\nolimits(E)$.
  • Let $U$ be any open subset of $X$ with a holomorphic coordinate $z$. Let $f$ be the endomorphism of $E_{|U}$ determined by $\theta=f\,dz$. Then, for any $P\in U$, $f_{|P}\in\mathop{\rm End}\nolimits(E_{|P})$ is regular in the following sense:
    $\bigl{\{}g\in\mathop{\rm End}\nolimits(E_{|P})\,\big{|}\,[f_{|P},g]=0\bigr{\}}=\bigoplus_{j=0}^{\mathop{\rm rank}\nolimits(E)-1}{\mathbb{C}}f_{|P}^{j}.$ (15)
    In particular, we obtain $\dim\bigl{\{}g\in\mathop{\rm End}\nolimits(E_{|P})\,\big{|}\,[f_{|P},g]=0\bigr{\}}=\mathop{\rm rank}\nolimits(E)$. As a result,
    $\mathop{\rm Ker}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|U}=\bigoplus_{j=0}^{\mathop{\rm rank}\nolimits(E)-1}\mathcal{O}_{U}f^{j}.$
Proof   Let us prove (15). Let $P\in X$ with a holomorphic coordinate neighbourhood $(X_{P},z_{P})$ as in §2.1.1. We obtain the coordinate $(z,\xi)$ on $T^{\ast}X_{P}$. There exists the decomposition (14) of $(E,\theta)_{|X_{P}}$. Each connected component $(\Sigma_{\theta})_{Q}$ is smooth. The eigenvalues of $f_{Q|P}$ are $\xi(Q)$. Hence, it is enough to consider the case where $|\Sigma_{\theta|P}|=1$ and $X=X_{P}$. Moreover, we may assume that $\Sigma_{\theta|P}=\{0\}$ in $T_{P}^{\ast}X$ by replacing $\theta$ with $\theta-\frac{1}{r}\mathop{\rm tr}\nolimits(\theta)\mathop{\rm id}\nolimits_{E}$. We set $\widetilde{X}=\Sigma_{\theta}$. Let $\pi:\widetilde{X}\to X$ denote the projection. We identify $T^{\ast}X=X\times{\mathbb{C}}$ by using the coordinate $z_{P}$. It induces a holomorphic coordinate of $\widetilde{X}$. There exists an isomorphism $E\simeq\pi_{\ast}(\mathcal{O}_{\widetilde{X}})$ under which the endomorphism $f$ is induced by the multiplication of $\xi$ on $\mathcal{O}_{\widetilde{X}}$. We obtain the frame $v_{j}=\pi_{\ast}(\xi^{j})$ $(j=0,\ldots,r-1)$ of $E$. Because $f_{P}(v_{j|P})=v_{j+1|P}$ $(j<r-1)$ and $f_{P}(v_{r-1|P})=0$, we obtain (15). The other claims follow from (15).
Corollary 2.2
Let $N_{Q}$ denote the nilpotent part of $f_{Q|P}$. Then, we have $N_{Q}(\mathcal{F}^{j}(E_{P|Q}))\subset\mathcal{F}^{j+1}(E_{P|Q})$ for $j=0,\ldots,r_{Q}-1$. Moreover, $N_{Q}$ induces an isomorphism $\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j}(E_{Q|P})\simeq\mathop{\rm Gr}\nolimits^{\mathcal{F}}_{j+1}(E_{Q|P})$ for $j=0,\ldots,r_{Q}-1$.
There exists the natural inclusion $\mathop{\rm Ker}\nolimits(\mathop{\rm ad}\nolimits\theta)\otimes K_{X}\subset\mathop{\rm End}\nolimits(E)\otimes K_{X}$. It induces a morphism
$\mathop{\rm Ker}\nolimits(\mathop{\rm ad}\nolimits\theta)\otimes K_{X}\longrightarrow\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta).$ (16)
The following lemma is obvious.
Lemma 2.3
The morphism (16) induces an isomorphism
$(\mathop{\rm Ker}\nolimits(\mathop{\rm ad}\nolimits\theta)\otimes K_{X})_{|X\setminus D(\theta)}\simeq\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X\setminus D(\theta)}.$ (17)
As a result, we obtain a splitting
$\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X\setminus D(\theta)}\to\bigl{(}\mathop{\rm End}\nolimits(E)\otimes K_{X}\bigr{)}_{|X\setminus D(\theta)}.$ (18)
2.2.3 Deformation complex in the case where $\Sigma_{\theta}$ is smooth
We continue to assume that $(E,\theta)$ is generically regular semisimple and that $\Sigma_{\theta}$ is smooth. Let $\pi:\Sigma_{\theta}\to X$ denote the projection. Let $\eta\in H^{0}(T^{\ast}X,\pi_{0}^{\ast}(K_{X}))$ denote the tautological $1$-form. The restriction to $\Sigma_{\theta}$ is also denoted by $\eta$. There exists a holomorphic line bundle $L$ on $\Sigma_{\theta}$ with $(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta)$, where $\pi_{\ast}(\eta)$ is the Higgs field of $\pi_{\ast}(L)$ induced by $\eta:L\to L\otimes\pi^{\ast}(K_{X})$. We obtain the following refinement of Lemma 2.1.
Proposition 2.4
There are natural isomorphisms:
$\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))=\mathop{\rm Ker}\nolimits\mathop{\rm ad}\nolimits(\theta)\simeq\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}},\quad\quad\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))=\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))\simeq\pi_{\ast}K_{\Sigma_{\theta}}.$
Proof   Because $L$ is an $\mathcal{O}_{\Sigma_{\theta}}$-module, $E=\pi_{\ast}L$ is naturally a $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}$-module. Hence, we obtain the induced morphism
$\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}})\to\mathop{\rm Ker}\nolimits\mathop{\rm ad}\nolimits\theta\subset\mathop{\rm End}\nolimits(E).$
We can easily check that it is an isomorphism by using Lemma 2.1. Thus, we obtain $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}\simeq\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))$.
Let $P\in X^{\circ}$. There exist a coordinate neighbourhood $(X_{P},z_{P})$ and a decomposition of the Higgs bundle (14) as in §2.1.1. Let $\psi$ be any holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ on $X_{P}$. We have the decomposition $\psi=\sum\psi_{Q,Q^{\prime}}$, where $\psi_{Q,Q^{\prime}}\in\mathop{\rm Hom}\nolimits(E_{Q^{\prime}},E_{Q})\otimes K_{X}$. We may naturally regard $\sum_{Q}\psi_{Q,Q}$ as a section of $\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. This induces an isomorphism $\rho_{P}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X_{P}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X_{P}}$. By considering the isomorphisms $\rho_{P}$ for any $P\in X^{\circ}$, we obtain the isomorphism $\rho_{X^{\circ}}:\mathop{\rm Cok}\nolimits\bigl{(}\mathop{\rm ad}\nolimits\theta\bigr{)}_{|X^{\circ}}\simeq\pi_{\ast}(K_{\Sigma_{\theta}})_{|X^{\circ}}$. Let us prove that $\rho_{X^{\circ}}$ extends to an isomorphism $\rho_{X}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)\simeq\pi_{\ast}(K_{\Sigma_{\theta}})$. It is enough to consider the case where $X$ is a neighbourhood $U$ of 00 in ${\mathbb{C}}$ and $\Sigma_{\theta|0}=\{0\}$. Let $f$ be the endomorphism of $E$ determined by $\theta=f\,dz$. We set $U^{\circ}=U\setminus\{0\}$ and $\widetilde{U}=\Sigma_{\theta}$.
Lemma 2.5
The isomorphism $\rho_{U^{\circ}}$ extends to a morphism $\rho_{U}:\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)\to\pi_{\ast}(K_{\widetilde{U}})$.
Proof   We set $r:=\mathop{\rm rank}\nolimits(E)$. Let $\varphi_{r}:{\mathbb{C}}\to{\mathbb{C}}$ be given by $\varphi_{r}(\zeta)=\zeta^{r}$. We set $U^{(r)}:=\varphi_{r}^{-1}(U)$. The induced morphism $U^{(r)}\to U$ is also denoted by $\varphi_{r}$. There exists the isomorphism $\gamma:U^{(r)}\simeq\widetilde{U}$ such that $\varphi_{r}=\pi\circ\gamma$.
There exists the line bundle $L$ on $\widetilde{U}$ with an isomorphism $(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta)$. We obtain the line bundle $L^{\prime}=\gamma^{\ast}(L)$ on $U^{(r)}$. Let $e$ be a frame of $L^{\prime}$ on $U^{(r)}$. By using the isomorphism $E\simeq\varphi_{r\ast}(L^{\prime})$, we obtain the frame $u_{j}=\varphi_{r\ast}(\zeta^{j}e)$ $(j=0,\ldots,r-1)$ of $E$. Let $\omega=\exp(2\pi\sqrt{-1}/r)$. We set
$v_{k}=\sum_{j=0}^{r-1}\omega^{-jk}\zeta^{-j}\varphi_{r}^{\ast}(u_{j}).$
They induce a frame of $\varphi_{r}^{\ast}(E)(\ast 0)$.
Let $\alpha$ be a holomorphic function $\alpha$ on $U^{(r)}$. The multiplication of $\alpha$ to $L^{\prime}$ induces an endomorphism $F_{\alpha}$ of $E$. We have
$\varphi_{r}^{\ast}(F_{\alpha})(v_{k})=\alpha(\omega^{k}\zeta)v_{k}.$
By using the identification $T^{\ast}U=U\times{\mathbb{C}}=\{(z,\xi)\in U\times{\mathbb{C}}\}$, the endomorphism $f$ is identified with $F_{\gamma^{\ast}\xi}$. Hence, the decomposition
$\varphi_{r}^{\ast}E_{|U^{(r)}\setminus\{0\}}=\bigoplus_{i=1}^{r}\mathcal{O}_{U^{(r)}\setminus\{0\}}v_{i}$
is the eigen decomposition of $\varphi_{r}^{\ast}(f)$.
Let ${\boldsymbol{u}}=(u_{0},\ldots,u_{r-1})$ and ${\boldsymbol{v}}=(v_{0},\ldots,v_{r-1})$. For any holomorphic section $g$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$ on $U$, we obtain the matrix valued $1$-form $G$ on $U$ determined by $g({\boldsymbol{u}})={\boldsymbol{u}}G$. Let $B$ be the matrix valued holomorphic function on $U^{(r)}$ determined by ${\boldsymbol{v}}=\varphi_{r}^{\ast}({\boldsymbol{u}})B$. We have
$\varphi_{r}^{\ast}(g)({\boldsymbol{v}})={\boldsymbol{v}}(B^{-1}\varphi_{r}^{\ast}(G)B).$
We can check that the entries of $B^{-1}\varphi_{r}^{\ast}(G)B$ are holomorphic $1$-forms on $U^{(r)}$ by using
$B_{j,k}=\omega^{-jk}\zeta^{-j},\quad\quad(B^{-1})_{k,j}=\frac{1}{r}\omega^{jk}\zeta^{j},\quad\quad\varphi_{r}^{\ast}(dz)=r\zeta^{r-1}d\zeta,\quad\quad(0\leq j,k\leq r-1).$
Then, the claim of Lemma 2.5 follows.
Lemma 2.6
Let $V$ be an $r$-dimensional complex vector space. Let $N$ be a nilpotent endomorphism such that $N^{r-1}\neq 0$. For any $H\in\sum_{j=1}^{r-1}{\mathbb{C}}N^{j}$, there exists $G\in\mathop{\rm End}\nolimits(V)$ such that $[N,G]=H$.
Proof   Let $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ denote the space of $(2\times 2)$-matrices with 00. Let
$h=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),\quad e=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right),\quad f=\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right).$
There exists a representation $\kappa$ of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$ on $V$ such that $\kappa(f)=N$. There exists the decomposition $\mathop{\rm End}\nolimits(V)={\mathbb{C}}\cdot\mathop{\rm id}\nolimits_{V}\oplus\mathop{\rm End}\nolimits_{0}(V)$, where $\mathop{\rm End}\nolimits_{0}(V)$ denotes the space of the trace-free endomorphisms of $V$. It is compatible with the induced $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$-action. It is well known that $\mathop{\rm End}\nolimits_{0}(V)$ does not contain a trivial representation of $\mathfrak{s}\mathfrak{l}(2,{\mathbb{C}})$, and $N^{j}$ $(j=1,\ldots,r-1)$ are contained in $\mathop{\rm End}\nolimits_{0}(V)$. Then, the claim is clear.
We obtain Proposition 2.4 from the following lemma.
Lemma 2.7
$\pi_{\ast}(K_{\widetilde{U}})$ equals the image of $\rho_{U}$.
Proof   Let $\nu$ be a holomorphic $1$-form on $\widetilde{U}$. The multiplication of $\nu$ induces a section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes K_{U}(\{0\})$. We have $[f,F_{\nu}]=0$. It implies $[f_{|0},\mathop{\rm Res}\nolimits_{0}(F_{\nu})]=0$. It is easy to see that $\mathop{\rm Res}\nolimits_{0}(F_{\nu})\in\sum_{j=1}^{r-1}{\mathbb{C}}f_{|0}^{j}$. By Lemma 2.6, there exists $g_{0}\in\mathop{\rm End}\nolimits(E_{|0})$ such that $[f,g_{0}]=\mathop{\rm Res}\nolimits_{0}(F_{\nu})$. Let $g$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)$ such that $g_{|0}=g_{0}$. Then, $\widetilde{F}=F_{\nu}-[\theta,z^{-1}g]$ is a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{U}$, and $\rho_{U}(\widetilde{F})=\pi_{\ast}(\nu)$. Thus, we obtain Lemma 2.7 and Proposition 2.4.
Corollary 2.8
If moreover $(E,\theta)$ is equipped with a non-degenerate symmetric pairing $C$, then we obtain
$\mathcal{H}^{0}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta;C)\bigr{)}=0,\quad\quad\mathcal{H}^{1}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta;C)\bigr{)}\simeq\pi_{\ast}K_{\Sigma_{\theta}}.$
Proof   Let $U$ be any open subset of $X$ with a holomorphic coordinate $z$. Let $f$ be the endomorphism of $E$ determined by $\theta=f\,dz$. Note that $f$ is self-adjoint with respect to $C$. Because $\mathop{\rm Ker}\nolimits(\mathop{\rm ad}\nolimits\theta)=\bigoplus_{j=0}^{r-1}\mathcal{O}_{U}f^{j}$, we obtain $\mathcal{H}^{0}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta;C)\bigr{)}_{|U}=0$.
Let $\tau$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ such that $\tau$ is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))\simeq\pi_{\ast}(K_{\Sigma_{\theta}}^{1})$. There exists a holomorphic section $s$ of $\mathop{\rm End}\nolimits(E)$ such that $\tau=[\theta,s]$. Let $\tau^{\lor}$ (resp. $s^{\lor}$) denote the adjoint of $\tau$ (resp. $s$) with respect to $C$. Because $\theta^{\lor}=\theta$, we have $\tau^{\lor}=-[\theta,s^{\lor}]$. If $\tau=\pm\tau^{\lor}$, we have $\tau=\frac{1}{2}[\theta,s\mp s^{\lor}]$. We also note that any holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\otimes K_{X}$ is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Then, we obtain $\mathcal{H}^{1}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta;C)\bigr{)}\simeq\pi_{\ast}K_{\Sigma_{\theta}}$.
2.2.4 Trace morphisms
As the composition of $\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}})=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))\to\mathop{\rm End}\nolimits(E)$ and $\mathop{\rm tr}\nolimits:\mathop{\rm End}\nolimits(E)\to\mathcal{O}_{X}$, we obtain the morphism
$\mathop{\rm tr}\nolimits:\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}})\to\mathcal{O}_{X}.$
It equals the standard trace map. Similarly, for any discrete subset $S\subset X$, we obtain
$\mathop{\rm tr}\nolimits:\pi_{\ast}\Bigl{(}\mathcal{O}_{\Sigma_{\theta}}\bigl{(}\ast\pi^{-1}(S)\bigr{)}\Bigr{)}\to\mathcal{O}_{X}(\ast S).$
We obtain the morphism $\mathop{\rm End}\nolimits(E)\otimes K_{X}\to K_{X}$ by taking the trace. It induces a morphism of complexes $\mathop{\rm Def}\nolimits(E,\theta)\to K_{X}[1]$. The induced morphism $\pi_{\ast}(K_{\Sigma_{\theta}})=\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))\longrightarrow K_{X}$ equals the trace morphism
$\mathop{\rm tr}\nolimits:\pi_{\ast}(K_{\Sigma_{\theta}})\to K_{X}$
defined as follows. We obtain $\pi_{\ast}(K_{\Sigma_{\theta}})_{|X^{\circ}}\to\bigl{(}\mathop{\rm End}\nolimits(\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}}))\otimes K_{X}\bigr{)}_{|X^{\circ}}$ as in §2.2.7 below. It extends to
$\pi_{\ast}(K_{\Sigma_{\theta}})\to\mathop{\rm End}\nolimits(\pi_{\ast}(\mathcal{O}_{\Sigma_{\theta}}))\otimes K_{X}(D(\theta)).$
Hence, we obtain $\mathop{\rm tr}\nolimits:\pi_{\ast}(K_{\Sigma_{\theta}})\to K_{X}(D(\theta))$. It is easy to check that the image is contained in $K_{X}$, i.e, $\mathop{\rm tr}\nolimits$ induces a morphism $\mathop{\rm tr}\nolimits:\pi_{\ast}(K_{\Sigma_{\theta}})\to K_{X}$.
2.2.5 Locally defined non-degenerate symmetric pairings
Let $P\in D(\theta)$. There exists a decomposition of the Higgs bundle (14). Let $r_{Q}:=\mathop{\rm rank}\nolimits(E_{Q})$, which equals the ramification index of $\pi$ at $Q$. Let $C_{L,Q}:L_{|(\Sigma_{\theta})_{Q}}\otimes L_{|(\Sigma_{\theta})_{Q}}\simeq\mathcal{O}_{(\Sigma_{\theta})_{Q}}\bigl{(}-(r_{Q}-1)Q\bigr{)}$ be a symmetric pairing. Let us recall that it induces a non-degenerate symmetric pairing of the Higgs bundle $(E_{Q},\theta_{Q})$ as studied in [23]. Indeed, we obtain a symmetric pairing $C_{Q}:E_{Q}\otimes E_{Q}\to\mathcal{O}_{X}$ as the composition of the natural morphisms
[…]
Here, the morphism $a$ is induced by the trace map $\mathop{\rm tr}\nolimits$.
Lemma 2.9
The restriction of $C_{Q|P}$ to $\mathcal{F}^{j}(E_{Q|P})\otimes\mathcal{F}^{k}(E_{Q|P})$ is 00 if $k+j>r_{Q}-1$. The induced pairings $\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{j}(E_{Q|P})\otimes\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{r_{Q}-1-j}(E_{Q|P})\longrightarrow{\mathbb{C}}$ are perfect. In particular, $C_{Q|P}$ is non-degenerate.
We set $C_{P}=\bigoplus C_{Q}$ and $C_{P|P}=\bigoplus C_{Q|P}$. Let $\nu$ be a holomorphic $1$-form on $\Sigma_{\theta}$. The restriction $\nu^{\circ}=\nu_{|X^{\circ}}$ induces a holomorphic section $F_{\nu^{\circ}}$ of $(\mathop{\rm End}\nolimits(E)\otimes K_{X})_{|X^{\circ}}$ as in §2.2.7 below. It induces a holomorphic section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes K_{X}(D)$. Because $F_{\nu}$ is self-adjoint with respect to $C_{P}$, $\mathop{\rm Res}\nolimits_{P}(F_{\nu})$ is self-adjoint with respect to $C_{P|P}$. We also note that $\mathop{\rm Res}\nolimits_{P}(F_{\nu})$ is nilpotent, and that there exists the decomposition $\mathop{\rm Res}\nolimits_{P}(F_{\nu})=\sum\mathop{\rm Res}\nolimits_{P}(F_{\nu})_{Q}$, where $\mathop{\rm Res}\nolimits_{P}(F_{\nu})_{Q}$ are endomorphisms of $E_{Q|P}$. Let $f_{P}$ be the endomorphism of $E_{|X_{P}}$ determined by $\theta_{|X_{P}}=f_{P}\,dz_{P}$. By using Corollary 2.8, we obtain the following lemma.
Lemma 2.10
There uniquely exists $g_{P}\in\mathop{\rm End}\nolimits(E_{|P})$ such that (i) $[f_{P|P},g_{P}]=\mathop{\rm Res}\nolimits_{P}(F_{\nu})$, (ii) $g_{P}$ is anti-self-adjoint with respect to $C_{P|P}$.
2.2.6 Dolbeault resolutions
We continue to assume that $(E,\theta)$ is generically regular semisimple and that $\Sigma_{\theta}$ is smooth. Let $\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ denote the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta)$. The space of global sections of $\mathcal{A}^{j}(\mathop{\rm Def}\nolimits(E,\theta))$ is denoted by $A^{j}(\mathop{\rm End}\nolimits(E))$ which is naturally identified with the space of $C^{\infty}$ $\mathop{\rm End}\nolimits(E)$-valued $j$-forms. If moreover $(E,\theta)$ is equipped with a non-degenerate symmetric pairing $C$, let $\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta;C))$ denote the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta;C)$, and the space of global sections of $\mathcal{A}^{j}(\mathop{\rm Def}\nolimits(E,\theta;C))$ is denoted by $A^{j}(\mathop{\rm End}\nolimits(E),\theta;C)$.
The Cauchy operator of $E$ is denoted by ¯$\overline{\partial}_{E}$. The induced Cauchy operators of $\mathop{\rm End}\nolimits(E)$ is also denoted by ¯$\overline{\partial}_{E}$.
Lemma 2.11
Suppose that $\tau\in A^{1}(\mathop{\rm End}\nolimits(E))$ is a $1$-cocycle, i.e., (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$. Then, $\tau$ induces a holomorphic section of $\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))\simeq\pi_{\ast}K_{\Sigma_{\theta}}$. The induced holomorphic section is denoted by $\rho(\tau)$.
Proof   There exists the natural morphism $\mathop{\rm Def}\nolimits(E,\theta)\to\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))[-1]$. By considering the induced morphism of the Dolbeault resolutions, we obtain the claim of the lemma.
For any section $\tau$ of $A^{j}(\mathop{\rm Def}\nolimits(E,\theta))$, let $\mathop{\rm Supp}\nolimits(\tau)$ denote the support of $\tau$.
Proposition 2.12
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau)$.
Proof   Let us begin with a special case.
Lemma 2.13
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (i) (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$, (ii) $\mathop{\rm Supp}\nolimits\tau\subset X^{\circ}$, (iii) $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau)$.
Proof   By the conditions, there exists a unique $C^{\infty}$-section $\sigma$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that (i) $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, (ii) $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau^{1,0})$. Because $\mathop{\rm ad}\nolimits(\theta):\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\to\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ is a monomorphism on $X^{\circ}$, it is easy to check that $\tau^{0,1}-\overline{\partial}_{E}\sigma=0$.
Let $U$ be any non-compact connected open subset of $X$. Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$.
Lemma 2.14
There exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
Proof   Because $U$ is non-compact, there exists a $C^{\infty}$-section $\sigma_{1}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that ¯$\overline{\partial}_{E}\sigma_{1}=\tau^{0,1}$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}$. Then, $\tau_{1}$ is a holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ which is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Because $U$ is non-compact, there exists a unique holomorphic section $\sigma_{2}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that $\mathop{\rm ad}\nolimits(\theta)(\sigma_{2})=\tau_{1}$. We set $\sigma=\sigma_{1}+\sigma_{2}$, and then (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ holds. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$ we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
Let $N$ be an open neighbourhood of $D$ such that each connected component is non-compact. By applying Lemma 2.14 to $\tau_{|N}$, there exists $\sigma_{N}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|N})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{N}=\tau_{|N}$. Let $N^{\prime}$ be a neighbourhood of $D$ in $N$ such that the closure of $N^{\prime}$ in $X$ is contained in $N$. Let $\chi:X\to[0,1]$ be a $C^{\infty}$-function such that $\chi=1$ on $N^{\prime}$ and $\chi=0$ on $X\setminus N$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)(\chi\sigma_{N})$. Because $\mathop{\rm Supp}\nolimits(\tau_{1})\subset X^{\circ}$, there exists $\sigma_{1}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}=\tau_{1}$ and that $\mathop{\rm Supp}\nolimits(\sigma_{1})\subset\mathop{\rm Supp}\nolimits(\tau_{1})$. We set $\sigma=\chi_{N}\sigma_{N}+\sigma_{1}$. Then, we have (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
2.2.7 Appendix
Assume that $(E,\theta)$ is regular semisimple. There exists a line bundle $L$ on $\Sigma_{\theta}$ with an isomorphism $(\pi_{\ast}L,\pi_{\ast}\eta)\simeq(E,\theta)$. For any simply connected open subset $U\subset X$, there exists the decomposition into Higgs bundles of rank $1$:
$(E,\theta)_{|U}=\bigoplus_{i=1}^{\mathop{\rm rank}\nolimits E}(E_{U,i},\theta_{U,i}).$ (19)
Let $\nu$ be a $(p,q)$-form on $\Sigma_{\theta}$. The multiplication of $\nu$ on $L$ induces a $C^{\infty}$-section $F_{\nu}$ of $A^{p,q}(\mathop{\rm End}\nolimits(E))$.
Lemma 2.15
  • $F_{\nu}$ is diagonal with respect to the decomposition (19).
  • If $\nu$ is a holomorphic function, then $F_{\nu}$ satisfies (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)F_{\nu}=0$. In particular, $(\partial_{E,h}+\mathop{\rm ad}\nolimits\theta_{h}^{\dagger})F_{\nu}$ is a harmonic $1$-form of a harmonic bundle $(E,\theta,h)$. (See Proposition 2.26 below.)
  • If $\nu$ is a holomorphic $(1,0)$-form, then ¯$\overline{\partial}_{E}F_{\nu}=0$. If $\nu$ is a $(0,1)$-form, then $\mathop{\rm ad}\nolimits\theta(F_{\nu})=0$.
2.3 Infinitesimal deformations and relative Higgs bundles
2.3.1 Infinitesimal deformations of a Higgs bundle
For any open subset $U\subset X$, we set $\mathcal{O}_{U^{[1]}}=\mathcal{O}_{U}[\epsilon]/\epsilon^{2}$. We obtain the ringed space $U^{[1]}=(U,\mathcal{O}_{U^{[1]}})$. Let $\iota_{U}:U\to U^{[1]}$ be the morphism of the ringed spaces defined by the identity map on $U$ and $\iota_{U}^{\ast}(f_{0}+f_{1}\epsilon)=f_{0}$. Let $q_{U}:U^{[1]}\to U$ be the morphism of the ringed spaces defined by the identity map on $U$ and $q_{U}^{\ast}(f)=f$.
An infinitesimal deformation of a Higgs bundle $(E,\theta)$ on $X$ is a locally free $\mathcal{O}_{X^{[1]}}$-module $\widetilde{E}$ with a morphism $\widetilde{\theta}:\widetilde{E}\to\widetilde{E}\otimes q^{\ast}K_{X}$ such that $\iota^{\ast}(\widetilde{E},\widetilde{\theta})=(E,\theta)$. The following lemma is well known. (See [24, 28].)
Lemma 2.16
$H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ equals the space of the isomorphism classes of infinitesimal deformations of $(E,\theta)$.
Proof   We just recall how an infinitesimal deformation of $(E,\theta)$ induces an element of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. We set $n=\mathop{\rm rank}\nolimits E$. Let $X=\bigcup_{i=0}^{m}U_{i}$ be an open covering of $X$ such that $U_{i}$ are Stein. Recall that any non-compact Riemann surface is Stein.
Let $(\widetilde{E},\widetilde{\theta})$ be an infinitesimal deformation of $(E,\theta)$. There exist isomorphisms $a_{i}:q^{\ast}(E_{|U_{i}})\simeq\widetilde{E}_{|U_{i}}$. Let $\psi_{i,j}$ denote the composition of the following isomorphisms:
$q_{U_{i}\cap U_{j}}^{\ast}(E_{|U_{i}\cap U_{j}})\simeq q_{U_{j}}^{\ast}(E_{|U_{j}})_{|(U_{i}\cap U_{j})^{[1]}}\stackrel{{\scriptstyle a_{j}}}{{\simeq}}\widetilde{E}_{|(U_{i}\cap U_{j})^{[1]}}\stackrel{{\scriptstyle a_{i}^{-1}}}{{\simeq}}q_{U_{i}}^{\ast}(E_{|U_{i}})_{|(U_{i}\cap U_{j})^{[1]}}\simeq q_{U_{i}\cap U_{j}}^{\ast}(E_{|U_{i}\cap U_{j}}).$
There exists $\rho_{i,j}\in H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))$ such that $\psi_{i,j}=\mathop{\rm id}\nolimits+\epsilon\rho_{i,j}$. We have $\rho_{j,k}-\rho_{i,k}+\rho_{i,j}=0$ on $U_{i}\cap U_{j}\cap U_{k}$. We also obtain $\kappa_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$ such that $a_{i}^{-1}\circ\widetilde{\theta}_{|U_{i}^{[1]}}\circ a_{i}=q_{U_{i}}^{\ast}(\theta)+\epsilon\kappa_{i}$. We have $\kappa_{j}-\kappa_{i}-\mathop{\rm ad}\nolimits(\theta)\rho_{i,j}=0$ on $U_{i}\cap U_{j}$. Then, the tuple
($(\{\rho_{i,j}\},\{\kappa_{i}\})\in\bigoplus_{i,j}H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))\oplus\bigoplus_{i}H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$
gives a $1$-cocycle of the Čech cohomology of $\mathop{\rm Def}\nolimits(E,\theta)$.
Let ($(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is a $1$-cocycle induced by different trivializations $\widetilde{a}_{i}$. There exists $b_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E))$ such that $\widetilde{a}_{i}=a_{i}\circ(1+\epsilon b_{i})$. Then, we have $\widetilde{\rho}_{i,j}=\rho_{i,j}+b_{j}-b_{i}$ on $U_{i}\cap U_{j}$ and $\widetilde{\kappa}_{i}=\kappa_{i}+\mathop{\rm ad}\nolimits(\theta)b_{i}$ on $U_{i}$. Hence, ($(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is cohomologous to ($(\{\rho_{i,j}\},\{\kappa_{i}\})$.
Remark 2.17
Let $(\mathcal{F}^{\bullet},d_{\mathcal{F}})$ be a complex of sheaves. For each $\mathcal{F}^{j}$, we obtain the Čech complex $\check{C}^{\bullet}(\mathcal{F}^{j})$ whose differential is denoted by $d_{\check{C}}$. (See [16, §B] for the signature of Čech complex.) We obtain a morphism of complexes $d_{\mathcal{F}}:\check{C}^{\bullet}(\mathcal{F}^{j})\to\check{C}^{\bullet}(\mathcal{F}^{j+1})$ satisfying $d_{\mathcal{F}}\circ d_{\mathcal{F}}=0$. Thus, we obtain a double complex $\check{C}^{\bullet}(\mathcal{F}^{\bullet})$. The differential of the total complex $\mathop{\rm Tot}\nolimits\check{C}^{\bullet}(\mathcal{F}^{\bullet})$ is induced by $d_{\check{C}}:\check{C}^{i}(\mathcal{F}^{j})\longrightarrow\check{C}^{i+1}(\mathcal{F}^{j})$ and $(-1)^{i}d_{\mathcal{F}}:\check{C}^{i}(\mathcal{F}^{j})\longrightarrow\check{C}^{i}(\mathcal{F}^{j+1})$.
2.3.2 Relative Higgs bundles and the induced infinitesimal deformations
Let $S$ be a complex manifold. Let $\Omega^{1}_{S\times X/S}$ denote the pull back of $K_{X}$ to $S\times X$ by the projection. Let $\mathcal{E}$ be a holomorphic vector bundle on $S\times X$. A morphism $\theta_{\mathcal{E}}:\mathcal{E}\to\mathcal{E}\otimes\Omega^{1}_{S\times X/S}$ is called a relative Higgs field of $\mathcal{E}$. Such a pair $(\mathcal{E},\theta_{\mathcal{E}})$ is called a relative Higgs bundle.
Let $s_{0}\in S$. Let $\iota_{s_{0}}:X\to S\times X$ denote the inclusion $X=\{s_{0}\}\times X\subset S\times X$. Let $\mathcal{O}_{S,s_{0}}$ denote the stalk of $\mathcal{O}_{S}$ at $s_{0}$ which is a local ring. A morphism of local rings $\mathcal{O}_{S,s_{0}}\to{\mathbb{C}}[\epsilon]/\epsilon^{2}$ is equivalent to a tangent vector $v\in T_{s_{0}}S$. It induces a morphism of ringed spaces $\iota_{v}:X^{[1]}\to S\times X$ such that $\iota_{s_{0}}=\iota_{v}\circ\iota$. We obtain an infinitesimal deformation $\iota_{v}^{\ast}(\mathcal{E},\theta_{\mathcal{E}})$ of $\iota_{s_{0}}^{\ast}(\mathcal{E},\theta_{\mathcal{E}})$.
2.3.3 Appendix: $C^{\infty}$-description of the induced infinitesimal deformation
For any complex submanifold $S^{\prime}$ of $S$, let $(\mathcal{E}_{S^{\prime}},\theta_{\mathcal{E},S^{\prime}})$ denote the restriction of $(\mathcal{E},\theta_{\mathcal{E}})$ to $S^{\prime}\times X$. Let $q_{S^{\prime}}:S^{\prime}\times X\to X$ denote the projection. Let $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S^{\prime}}$ denote the sheaf of $C^{\infty}$-functions $f$ on $S^{\prime}\times X$ such that ¯$\overline{\partial}_{S^{\prime}}f=0$.
Lemma 2.18
There exist a neighbourhood $S_{0}$ of $s_{0}$ in $S$ and a $C^{\infty}$-isomorphism $\varphi:q_{S_{0}}^{\ast}(\mathcal{E}_{s_{0}})\simeq\mathcal{E}_{S_{0}}$ such that $\varphi$ is holomorphic in the $S_{0}$-direction, i.e., $\varphi$ induces a morphism of $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S^{\prime}}$-modules.
Proof   We set $n=\mathop{\rm rank}\nolimits\mathcal{E}$. Let $X=\bigcup_{i=0}^{m}U_{i}$ be an open covering of $X$ such that $U_{i}\cap U_{j}=\emptyset$ $(1\leq i\neq j\leq m)$, i.e., $U_{i}$ $(1\leq i\leq m)$ has the non-empty intersection with only $U_{0}$. There exist a neighbourhood $S_{1}$ of $s_{0}$ in $S$ and ${\boldsymbol{v}}_{i}$ be frames of $\mathcal{E}_{|S_{1}\times U_{i}}$. We obtain holomorphic maps $A_{i}(s,x):S_{1}\times(U_{0}\cap U_{i})\to\mathop{\rm GL}\nolimits(n,{\mathbb{C}})$ determined by ${\boldsymbol{v}}_{0|S_{1}\times(U_{0}\cap U_{i})}={\boldsymbol{v}}_{i|S_{1}\times(U_{0}\cap U_{i})}A_{i}$. We may assume that $A_{i}$ extends to a holomorphic map on a neighbourhood of $S_{1}\times\overline{U_{0}\cap U_{i}}$. Let $I_{n}\in\mathop{\rm GL}\nolimits(n,{\mathbb{C}})$ denote the identity matrix. Let $\mathcal{U}$ be a convex open neighbourhood of $I_{n}$ in $\mathop{\rm GL}\nolimits(n,{\mathbb{C}})$. There exists a neighbourhood $S_{2}$ of $s_{0}$ in $S_{1}$ such that $A_{i}(s_{0},x)^{-1}A_{i}(s,x)\in\mathcal{U}$ for any $x\in\overline{U_{0}\cap U_{i}}$.
Let $U_{i}^{\prime}$ be a relatively compact neighbourhood of $(X\setminus U_{0})\cap U_{i}$ in $U_{i}$. Let $U_{i}^{\prime\prime}$ be a relatively compact neighbourhood of $\overline{U_{i}}^{\prime}$ in $U_{i}$. We assume that $\overline{U_{i}}^{\prime\prime}\cap\overline{U}_{j}=\emptyset$ for any $j\neq 0,i$. Let $\chi_{i}:X\to[0,1]$ be a $C^{\infty}$-function such that $\chi_{i}=0$ on $U_{i}^{\prime}$ and $\chi_{i}=1$ on $X\setminus U_{i}^{\prime\prime}$.
Let $\varphi_{U_{0}}$ be the isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{0}}\simeq\mathcal{E}_{|S_{2}\times U_{0}}$ defined by $\varphi_{U_{0}}(q_{S_{2}}^{\ast}({\boldsymbol{v}}_{0|\{s_{0}\}\times U_{0}}))={\boldsymbol{v}}_{0|SS_{2}\times U_{0}}$. Let $\varphi_{U_{i}}$ be the isomorphism $q_{S_{2}}^{\ast}(\mathcal{E})_{|S_{2}\times U_{i}}\simeq\mathcal{E}_{|S_{2}\times U_{i}}$ defined by
$\varphi_{U_{i}}(q_{S_{2}}^{\ast}({\boldsymbol{v}}_{i|\{s_{0}\}\times U_{0}}))={\boldsymbol{v}}_{i|S_{2}\times U_{0}}\Bigl{(}I_{n}+\chi_{i}\cdot\bigl{(}A_{i}(s,x)A_{i}(s_{0},x)^{-1}-I_{n}\bigr{)}\Bigr{)}.$
We obtain a $C^{\infty}$-isomorphism $\varphi:q_{S_{2}}^{\ast}(\mathcal{E}_{s_{0}})\simeq\mathcal{E}_{S_{2}}$ from $\varphi_{U_{i}}$ $(1\leq i\leq m)$ and the restriction of $\varphi_{0}$ to $X\setminus\bigcup_{i=1}^{m}\overline{U}_{i}^{\prime\prime}$. It is holomorphic in the $S_{2}$-direction.
We set $(E,\theta):=(\mathcal{E}_{s_{0}},\theta_{\mathcal{E},s_{0}})$. We obtain a $C^{\infty}$-section $B^{0,1}$ of $q_{S_{0}}^{-1}(\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}_{X})$ by
$\varphi^{\ast}(\overline{\partial}_{\mathcal{E}_{S_{0}}})=\overline{\partial}_{q_{S_{0}}^{\ast}(E)}+B^{0,1}.$
We obtain a $C^{\infty}$-section $B^{1,0}$ of $q_{S_{0}}^{-1}(\mathop{\rm End}\nolimits(E)\otimes\Omega^{1,0}_{X})$ by
$\varphi^{\ast}(\theta_{\mathcal{E},S_{0}})=q_{S_{0}}^{\ast}(\theta)+B^{1,0}.$
Because $\varphi$ is holomorphic in the $S_{0}$-direction, $B^{1,0}$ and $B^{0,1}$ are holomorphic in the $S_{0}$-direction.
For any tangent vector $v\in T_{s_{0}}S$, we obtain the section $\widetilde{v}$ of $T_{s_{0}}S\times X\subset\iota_{0}^{\ast}(T(S\times X))$. By using a $C^{\infty}$-isomorphism $\varphi:q_{S_{0}}^{\ast}(E)\simeq\mathcal{E}$, we define the derivative $\widetilde{v}(B^{1,0})$ and $\widetilde{v}(B^{0,1})$. We obtain a section $\tau=\widetilde{v}(B^{1,0})+\widetilde{v}(B^{0,1})$ of $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$.
Lemma 2.19
$\tau$ is a closed $1$-form in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$, and its cohomology class is $\iota_{v}(\mathcal{E},\theta_{\mathcal{E}})$.
Proof   Because ¯$\overline{\partial}_{q_{S_{0}}^{\ast}(E)}(B^{1,0})+\bigl{[}q_{S_{0}}^{\ast}(\theta),B^{0,1}\bigr{]}=0$, we obtain (¯$(\overline{\partial}+\mathop{\rm ad}\nolimits\theta)(\tau)=\overline{\partial}_{E}(\widetilde{v}(B^{1,0}))+[\theta,\widetilde{v}(B^{0,1})]=0$.
There exists the following naturally defined quasi-isomorphisms:
$A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\longrightarrow\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr{)}\longleftarrow\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta)\bigr{)}.$
Here, $\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ is the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta)$. See Remark 2.17 for the construction of $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}(\mathcal{F}^{\bullet})$ from a complex of sheaves $\mathcal{F}^{\bullet}$.
Let us compare the cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E},\theta_{\mathcal{E}})$ in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr{)}$. In the following argument, we shall identify $\mathcal{E}$ and $q_{S_{0}}^{-1}(E)$ by $\varphi$, and we omit to denote $\varphi$. Let $X=\bigcup_{i=0}^{m}U_{i}$ be an open covering. We may assume that there exist holomorphic frames ${\boldsymbol{v}}_{i}$ of $\mathcal{E}_{|S_{0}\times U_{i}}$. We obtain the matrix valued holomorphic function $D_{i,j}$ on $S_{0}\times(U_{i}\cap U_{j})$ determined by ${\boldsymbol{v}}_{j|S_{0}\times(U_{i}\cap U_{j})}={\boldsymbol{v}}_{i|S_{0}\times(U_{i}\cap U_{j})}D_{ij}$. We also obtain the matrix valued holomorphic $1$-form $\Theta_{i}$ on $S_{0}\times U_{i}$ determined by $\theta_{\mathcal{E}}{\boldsymbol{v}}_{i}={\boldsymbol{v}}_{i}\Theta_{i}$. The restriction of ${\boldsymbol{v}}_{i}$ to $\{s_{0}\}\times U_{i}$ are denoted by ${\boldsymbol{v}}_{i|s_{0}}$. Let $\kappa_{i}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes\Omega_{U_{i}}^{1}$ determined by $\kappa_{i}{\boldsymbol{v}}_{i|s_{0}}={\boldsymbol{v}}_{i|s_{0}}\widetilde{v}(\Theta_{i})$. Let $\rho_{i,j}$ be the holomorphic section of $\mathop{\rm End}\nolimits(E)_{U_{i}\cap U_{j}}$ determined by $\rho_{i,j}{\boldsymbol{v}}_{j}={\boldsymbol{v}}_{i}\widetilde{v}(D_{i,j})$. By definition, ($(\{\rho_{i,j}\},\{\kappa_{i}\})$ is the $1$-cocycle of $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta)\bigr{)}$ induced by $\iota_{v}(\mathcal{E})$. Let $M_{i}$ be the matrix valued $C^{\infty}$-function determined by ${\boldsymbol{v}}_{i}=q^{-1}_{S_{0}}({\boldsymbol{v}}_{i|s_{0}})M_{i}$. The entries of $M_{i}$ are contained in $\mathcal{C}^{\infty}_{X}\mathcal{O}_{S_{0}}$. Because ¯$\overline{\partial}_{\mathcal{E}}(q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}}))=q_{S_{0}}^{-1}({\boldsymbol{v}}_{i|s_{0}})(-\overline{\partial}(M_{i})M_{i}^{-1})$, we obtain
$B^{0,1}q^{-1}({\boldsymbol{v}}_{i|s_{0}})=q^{-1}({\boldsymbol{v}}_{i|s_{0}})(-\overline{\partial}(M_{i})M_{i}^{-1}).$
Hence, we obtain
$\widetilde{v}(B^{0,1})q^{-1}({\boldsymbol{v}}_{i|s_{0}})=q^{-1}({\boldsymbol{v}}_{i|s_{0}})\bigl{(}-\overline{\partial}\widetilde{v}(M_{i})\bigr{)}.$ (20)
Because $q_{S_{0}}^{-1}(\theta)q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})=q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})M_{i}\Theta_{i}M_{i}^{-1}$, we obtain
$B^{1,0}q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})=q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})\bigl{(}M_{i}\Theta_{i}M_{i}^{-1}-q_{S_{0}}^{\ast}(\Theta_{i|s_{0}})\bigr{)}.$
Here, $\Theta_{i|s_{0}}$ denotes the restriction of $\Theta_{i}$ to $\{s_{0}\}\times U_{i}$. We obtain
$\widetilde{v}(B^{1,0})q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})=q_{S_{0}}^{\ast}({\boldsymbol{v}}_{i|s_{0}})\bigl{(}\widetilde{v}(\Theta_{i})-[\Theta_{i|s_{0}},\widetilde{v}(M_{i})]\bigr{)}.$ (21)
Let $D_{ij|s_{0}}$ denote the restriction of $D_{ij}$ to $\{s_{0}\}\times(U_{i}\cap U_{j})$. Because $M_{i}D_{ij}=q_{S_{0}}^{\ast}(D_{ij|s_{0}})M_{j}$, we obtain
$\widetilde{v}(D_{ij})=D_{ij|s_{0}}\widetilde{v}(M_{j})-\widetilde{v}(M_{i})D_{ij|s_{0}}.$ (22)
Let $\sigma_{i}$ be the $C^{\infty}$-section of $\mathop{\rm End}\nolimits(E)$ on $U_{i}$ determined by $\sigma_{i}{\boldsymbol{v}}_{i}={\boldsymbol{v}}\bigl{(}\widetilde{v}(M_{i})\bigr{)}$. By (20), (21) and (22), we obtain
$\kappa_{i}-(B^{1,0}+B^{0,1})_{|U_{i}}=(\overline{\partial}+\mathop{\rm ad}\nolimits\theta)\sigma_{i},\quad\rho_{i,j}=\bigl{(}\sigma_{j|U_{i}\cap U_{j}}-\sigma_{i|U_{i}\cap U_{j}}\bigr{)}.$
Hence, the $1$-cocycles induced by $\tau$ and $\iota_{v}(\mathcal{E})$ are cohomologue in the complex $\mathop{\rm Tot}\nolimits^{\bullet}\check{C}^{\bullet}\bigl{(}\mathcal{A}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))\bigr{)}$.
Remark 2.20
For a different $C^{\infty}$-isomorphism in Lemma 2.18, we obtain another $1$-cocycle $\tau^{\prime}$ of $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The difference $\tau-\tau^{\prime}$ is a $1$-coboundary. It follows from Lemma 2.19. We can also check it directly.
2.4 Harmonic $1$-forms
2.4.1 Harmonic $1$-forms of $E$
Let $h$ be a harmonic metric of $(E,\theta)$. Let =¯$\nabla=\overline{\partial}_{E}+\partial_{E,h}$ be the Chern connection of $(E,\theta)$. Let $\theta^{\dagger}_{h}$ denote the adjoint of $\theta$ with respect to $h$. We fix a conformal metric $g_{X}$ of $X$. Let (¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}:A^{i+1}(E)\to A^{i}(E)$ denote the formal adjoint of ¯$\overline{\partial}_{E}+\theta$ with respect to $h$ and $g_{X}$.
Definition 2.21
$\tau\in A^{1}(E)$ is called a harmonic $1$-form of $(E,\theta,h)$ if (¯$(\overline{\partial}_{E}+\theta)\tau=0$ and (¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=0$. Here, (¯$(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}$ denotes adjoint of ¯$\overline{\partial}_{E}+\theta$ with respect to $h$ and $g_{X}$.
Remark 2.22
Let $\Delta_{h,g_{X}}=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}(\overline{\partial}_{E}+\theta)+(\overline{\partial}_{E}+\theta)(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}$. When $X$ is compact, we have $\Delta_{h,g_{X}}\tau=0$ if and only if (¯$(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)_{h,g_{X}}^{\ast}\tau=0$.
Lemma 2.23
$\tau$ is a harmonic $1$-form if and only if (¯$(\overline{\partial}_{E}+\theta)\tau=(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
Proof   Let $\Lambda_{X}$ denote the adjoint of the multiplication of the Kähler form associated with $g_{X}$. According to [25, Lemma 3.1], we have
(¯$(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{X}}\tau=-\sqrt{-1}[\Lambda_{X},\partial_{E,h}+\theta_{h}^{\dagger}]\tau=-\sqrt{-1}\Lambda_{X}((\partial_{E,h}+\theta_{h}^{\dagger})\tau).$
It is 00 if and only if $(\partial_{E,h}+\theta_{h}^{\dagger})\tau=0$.
For any complex number $\lambda$, we obtain $\mathbb{D}^{\lambda}=\overline{\partial}_{E}+\theta+\lambda(\partial_{E,h}+\theta^{\dagger}_{h}):A^{i}(E)\to A^{i+1}(E)$. Let $\mathbb{D}^{\lambda\ast}_{h,g_{X}}:A^{i+1}(E)\to A^{i}(E)$ denote the formal adjoint of $\mathbb{D}^{\lambda}$ with respect to $h$ and $g_{X}$.
Corollary 2.24
For a harmonic $1$-form $\tau$, we have $\mathbb{D}^{\lambda}\tau=0$ and $\mathbb{D}^{\lambda\ast}_{h,g_{X}}\tau=0$ for any $\lambda$.
Let $X^{\dagger}$ denote the conjugate of $X$. We obtain the Higgs bundle $(E,\partial_{E,h},\theta^{\dagger}_{h})$ on $X^{\dagger}$, and $h$ is a harmonic metric of $(E,\partial_{E,h},\theta^{\dagger}_{h})$.
Corollary 2.25
$\tau\in A^{1}(E)$ is a harmonic $1$-form for $(E,\overline{\partial}_{E},\theta,h)$ if and only if it is a harmonic $1$-form for $(E,\partial_{E,h},\theta^{\dagger}_{h},h)$.
The following proposition is useful to construct harmonic $1$-forms.
Proposition 2.26
Let $s$ be a section of $E$ such that (¯$(\overline{\partial}_{E}+\theta)s=0$. Then, $\tau=(\partial_{E,h}+\theta_{h}^{\dagger})s$ is a harmonic $1$-form of $(E,\theta,h)$. Similarly, if $s^{\prime}$ is a section of $E$ such that $(\partial_{E,h}+\theta^{\dagger}_{h})s^{\prime}=0$, then (¯$(\overline{\partial}_{E}+\theta)s^{\prime}$ is a harmonic $1$-form of $(E,\theta,h)$.
Proof   We have $(\partial_{E,h}+\theta_{h}^{\dagger})\tau=(\partial_{E,h}+\theta_{h}^{\dagger})^{2}s=0$. We also have
(¯$(\overline{\partial}_{E}+\theta)\tau=(\overline{\partial}_{E}+\theta)(\partial_{E,h}+\theta_{h}^{\dagger})s=-(\partial_{E,h}+\theta_{h}^{\dagger})(\overline{\partial}_{E}+\theta)s=0.$
Thus, we are done.
The following proposition is useful in the study of pairings of harmonic $1$-forms.
Proposition 2.27
Let $\tau_{i}$ $(i=1,2)$ be harmonic $1$-forms of $(E,\theta,h)$ on $X$. Let $U$ be a simply connected relatively compact open subset of $X$ such that $\partial U$ is smooth. Note that, because $\mathbb{D}^{1}\tau_{1}=0$, there exists a section $\sigma_{1}$ of $E$ on $U$ such that $\tau_{1|U}=\mathbb{D}^{1}\sigma_{1}$. Then, we obtain
$\int_{U}\Bigl{(}\sqrt{-1}h(\tau_{1}^{1,0},\tau_{2}^{1,0})-\sqrt{-1}h(\tau_{1}^{0,1},\tau_{2}^{0,1})\Bigr{)}=\sqrt{-1}\int_{\partial U}\bigl{(}h(\sigma_{1},\tau_{2}^{1,0})-h(\sigma_{1},\tau_{2}^{0,1})\bigr{)}.$ (23)
Proof   It follows from $\mathbb{D}_{h,g_{X}}^{1\ast}\tau_{2}=0$.
Note that the left hand side of (23) is the Hermitian product of the space of $L^{2}$-sections of $\mathcal{A}^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ on $U$.
Corollary 2.28
Suppose that there exists positive constants $B_{i}$ $(i=1,2)$ such that $|\tau_{2}|_{h,g_{X}}\leq B_{1}$ and $|\sigma|_{h}\leq B_{2}$ on a neighbourhood of $\partial U$. Then, there exists $B_{3}>0$, depending only on $B_{i}$ $(i=1,2)$ such that
$\|\tau_{|U}\|_{L^{2},h}\leq B_{3}.$
Here, $\|\tau_{|U}\|_{L^{2},h}$ denotes the $L^{2}$-norm of $\tau_{|U}$ with respect to $h$ and $g_{X}$. Note that it is independent of the choice of a conformal metric $g_{X}$.
Remark 2.29
Proposition 2.27 may be related with the exactness in [4, Lemma 12].
2.4.2 Harmonic $1$-forms of $\mathop{\rm End}\nolimits(E)$
Suppose that $(E,\theta)$ is generically regular semisimple and that $\Sigma_{\theta}$ is smooth. Let $C$ be a non-degenerate symmetric pairing of $(E,\theta)$. Let $h$ be a harmonic metric of $(E,\theta)$ compatible with $C$. The Chern connection of $\mathop{\rm End}\nolimits(E)$ is also denoted by $\nabla_{h}=\partial_{E,h}+\overline{\partial}_{E}$.
Let $s$ be a holomorphic section of $\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))$. We obtain the harmonic $1$-form $\tau=(\partial_{E,h}+\mathop{\rm ad}\nolimits\theta_{h}^{\dagger})s$ of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$ by Proposition 2.26. There exists a holomorphic function $\alpha$ on $\Sigma_{\theta}$ such that $s=F_{\alpha}$ by Proposition 2.4.
Proposition 2.30
  • $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$. Let $\rho(\tau)$ denote the induced section of $\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))\simeq\pi_{\ast}\Omega^{1}_{\Sigma_{\theta}}$.
  • $\rho(\tau)=\pi_{\ast}d\alpha$.
Proof   Let $P\in X^{\circ}$. Let $(X_{P},z_{P})$ be a simply connected coordinate neighbourhood around $P$ such that $X_{P}\cap D(\theta)=\emptyset$. There exists a decomposition $(E,\theta)_{|X_{P}}=\bigoplus_{i=1}^{\mathop{\rm rank}\nolimits(E)}(E_{P,i},\theta_{P,i})$ into Higgs bundles of rank $1$. For each $i$, let $v_{P,i}$ be a holomorphic frame of $E_{P,i}$ such that $C(v_{P,i},v_{P,i})=1$. We obtain a frame ${\boldsymbol{v}}_{P}=(v_{P,i})$. Let $B$ be the matrix determined by $\partial_{E,h}{\boldsymbol{v}}={\boldsymbol{v}}\cdot B\,dz$. Because $\nabla_{h}=\overline{\partial}_{E}+\partial_{E,h}$ is compatible with $C$, we obtain that the matrix $B$ is anti-symmetric. Then, it is easy to check the claims.
Remark 2.31
Let $X^{\dagger}$ denote the conjugate of $X$, and let $E^{\dagger}_{h}$ denote the holomorphic vector bundle $(E,\partial_{E,h})$ on $X^{\dagger}$. Then, $\theta^{\dagger}_{h}$ is a Higgs field of $E^{\dagger}_{h}$, and $h$ is a harmonic metric of $(E^{\dagger}_{h},\theta^{\dagger}_{h})$. Because $\partial_{E,h}(s^{\dagger}_{h})=(\overline{\partial}_{E}s)^{\dagger}_{h}=0$ and $[\theta^{\dagger}_{h},s^{\dagger}_{h}]=-[\theta,s]^{\dagger}_{h}=0$, $s^{\dagger}_{h}$ is a holomorphic section of $\mathcal{H}^{0}(E^{\dagger},\theta^{\dagger}_{h})$. Hence, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta^{\dagger}_{h})s^{\dagger}_{h}$ is also a harmonic $1$-form of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$.
2.5 Some estimates for harmonic bundles
2.5.1 Uhlenbeck theorem
For $r>0$, we set $\mathcal{B}(r)=\{|z|<r\}$. For a matrix valued $C^{\infty}$-function $s$ on $\mathcal{B}(r)$, and for $0<r_{1}\leq r$, let $\|s_{|\mathcal{B}(r_{1})}\|_{L_{\ell}^{p},g_{0}}$ denote the $L_{\ell}^{p}$-norm of $s_{|\mathcal{B}(r_{1})}$ with respect to the Euclidean metric $g_{0}$ and an appropriate norm of the space of matrices. If $r_{1}=r$, $\|s_{|\mathcal{B}(r)}\|_{L_{\ell}^{p},g_{0}}$ is denoted by $\|s\|_{L_{\ell}^{p},g_{0}}$.
Let $V$ be a holomorphic vector bundle on $\mathcal{B}(1)$ equipped with a Hermitian metric $h$. Let $\nabla_{h}$ denote the Chern connection, and let $R(h)=\mathfrak{R}\,dz\,d\overline{z}$ denote the curvature. For a $C^{\infty}$-section $s$ of $V$ or $\mathop{\rm End}\nolimits(V)$ and for $10<r_{1}<$ 1, let $\|s_{|\mathcal{B}(r_{1})}\|_{L_{\ell}^{p},h,g_{0}}$ denote the $L_{\ell}^{p}$-norm of $s_{|\mathcal{B}(r_{1})}$ with respect to $\nabla_{h}$, $h$ and $g_{0}$. For a frame ${\boldsymbol{v}}$ of $V$, let $\mathcal{A}^{{\boldsymbol{v}}}$ denote the connection form of $\nabla_{h}$ with respect to ${\boldsymbol{v}}$. If $r_{1}=1$, $\|s_{|\mathcal{B}(1)}\|_{L_{\ell}^{p},h,g_{0}}$ is denoted by $\|s\|_{L_{\ell}^{p},h,g_{0}}$. We recall the following theorem of Uhlenbeck. (See [27, Theorem 1.3] for more precise statement.)
Theorem 2.32 (Uhlenbeck)
Let $p>2$. There exist positive constants $\epsilon_{0}$ and $C_{0}$, which are independent of $(V,h)$, such that the following holds:
  • Suppose that $|\mathfrak{R}|_{h}\leq\epsilon_{0}$ on $\mathcal{B}(1)$. Then, there exists an orthonormal frame ${\boldsymbol{v}}$ of $V$ such that $\|\mathcal{A}^{{\boldsymbol{v}}}\|_{L_{1}^{p},g_{0}}\leq C_{0}\|\mathfrak{R}\|_{L^{p},h,g_{0}}$ and $d^{\ast}\mathcal{A}^{{\boldsymbol{v}}}=0$. Here, $d^{\ast}$ denotes the formal adjoint of the exterior derivative $d$.
The following lemma is also well known.
Lemma 2.33
Let $\ell\in{\mathbb{Z}}_{\geq 0}$ and $p>2$. Let $10<r_{1}<r_{2}<$ 1. There exists a positive constant $C_{1}(\ell,p,r_{1},r_{2})$ depending only on $\ell,p,r_{1},r_{2}$ such that the following holds.
  • Suppose that $|\mathfrak{R}|_{h}\leq\epsilon_{0}$ on $\mathcal{B}(1)$. Let ${\boldsymbol{v}}$ be a frame as in Theorem 2.32. Then,
    $\bigl{\|}\mathcal{A}^{{\boldsymbol{v}}}_{|\mathcal{B}(r_{1})}\bigr{\|}_{L_{\ell}^{p},g_{0}}\leq C_{1}(\ell,p,r_{1},r_{2})\|\mathfrak{R}_{|\mathcal{B}(r_{2})}\|_{L_{\ell}^{p},h,g_{0}}.$ (24)
Proof   Let $\mathcal{R}^{{\boldsymbol{v}}}$ be the matrix valued function determined by $\mathfrak{R}{\boldsymbol{v}}={\boldsymbol{v}}\cdot\mathcal{R}^{{\boldsymbol{v}}}$. The following holds.
$d(\mathcal{A}^{{\boldsymbol{v}}})+\mathcal{A}^{{\boldsymbol{v}}}\wedge\mathcal{A}^{{\boldsymbol{v}}}=\mathcal{R}^{{\boldsymbol{v}}}\,dz\,d\overline{z},\quad d^{\ast}(\mathcal{A}^{{\boldsymbol{v}}})=0$
We can obtain (24) by using the elliptic regularity and a standard bootstrapping argument.
2.5.2 Universal boundedness of higher derivatives of Higgs fields
Let $(E,\theta,h)$ be a harmonic bundle on $\mathcal{B}(r_{0})$ for some $r_{0}>1$. Let $f$ be the endomorphism of $E$ determined by $\theta=f\,dz$. Suppose that there exists $B_{1}>0$ such that the following holds.
  • For any $z\in\mathcal{B}(r_{0})$, any eigenvalue $\alpha$ of $f_{|z}$ satisfies $|\alpha|\leq B_{1}$.
Let $\nabla_{h}$ denote the Chern connection of $(E,h)$. Let $R(h)=\mathfrak{R}\,dz\,d\overline{z}$ denote the curvature of $\nabla_{h}$. Let $g_{0}=dz\,d\overline{z}$ be the Euclidean metric. For a section $\alpha$ of $\mathop{\rm End}\nolimits(E)$, and for $\ell\in{\mathbb{Z}}_{\geq 0}$ and $p\geq 1$, let $\|\alpha_{|\mathcal{B}(r)}\|_{L_{\ell}^{p},h,g_{0}}$ denote the $L_{\ell}^{p}$-norm of $\alpha$ on $\mathcal{B}(r)$ with respect to $h$ and $g_{0}$ and the derivatives induced by $\nabla_{h}$.
Let $0<r_{1}<r_{0}$. We set $r_{2}=\frac{r_{0}+r_{1}}{2}$. By a variant of Simpson’s main estimate (see [20, Proposition 2.1]), we obtain the following estimate in [23, Proposition 4.1].
Proposition 2.34
There exists $B_{2}>0$ such that
$\sup|f_{|\mathcal{B}(r_{2})}|_{h}+\sup|f^{\dagger}_{h|\mathcal{B}(r_{2})}|_{h}+\sup|\mathfrak{R}_{|\mathcal{B}(r_{2})}|_{h}\leq B_{2}.$
Moreover, for any $\ell\in{\mathbb{Z}}_{\geq 0}$ and $p\geq 1$, there exists $B_{3}(\ell,p)>0$, depending only on $r_{0}-r_{1}$, $B_{1}$ and $(\ell,p)$ such that
$\|f_{|\mathcal{B}(r_{2})}\|_{L_{\ell}^{p},h,g_{0}}+\|f^{\dagger}_{h|\mathcal{B}(r_{2})}\|_{L_{\ell}^{p},h,g_{0}}+\|\mathfrak{R}_{|\mathcal{B}(r_{2})}\|_{L_{\ell}^{p},h,g_{0}}\leq B_{3}(\ell,p).$
2.5.3 The sup norm and the $L^{2}$-norm of harmonic $1$-forms
Let $\tau\in A^{1}(E)$ be a harmonic $1$-form of $(E,\theta,h)$ on $\mathcal{B}(r_{0})$. Let $\|\tau\|_{L^{2},h}$ be the $L^{2}$-norm of $\tau$ with respect to $h$ and $g_{0}$. Note that $\|\tau\|_{L^{2},h}$ is independent of the choice of a conformal metric of $\mathcal{B}(r_{0})$.
Proposition 2.35
There exists $B_{4}>0$, depending only on $r_{0}-r_{1}$ and $B_{1}$, such that the following holds:
$\sup_{\mathcal{B}(r_{1})}|\tau|_{h,g_{0}}\leq B_{4}\|\tau\|_{L^{2},h}.$
Proof   Let $\epsilon_{0}$ be as in Theorem 2.32. Let $B_{2}$ be as in Proposition 2.34. We set $\epsilon_{1}:=\min\{2^{-1}(r_{2}-r_{1}),B_{2}^{-1/2}\epsilon_{0}^{1/2}\}$, which satisfies $0<\epsilon_{1}<r_{2}-r_{1}$ and $B_{2}\epsilon_{1}^{2}\leq\epsilon_{0}$. For any $z_{0}\in\mathcal{B}(r_{1})$, let $\psi_{z_{0}}:\mathcal{B}(1)\longrightarrow\mathcal{B}(r_{2})$ be defined by $\psi_{z_{0}}(\zeta)=z_{0}+\epsilon_{1}\zeta$. We obtain $(\widehat{E}_{z_{0}},\widehat{\theta}_{z_{0}},\widehat{h}_{z_{0}})=\psi_{z_{0}}^{\ast}(E,\theta,h)$ and $\widehat{\tau}_{z_{0}}=\psi_{z_{0}}^{\ast}(\tau)$ on $\mathcal{B}(1)$. Let $\nabla_{\widehat{h}_{z_{0}}}$ denote the Chern connection of $(\widehat{E}_{z_{0}},\widehat{h}_{z_{0}})$, and let $R(\widehat{h}_{z_{0}})$ denote the curvature of $\nabla_{\widehat{h}_{z_{0}}}$. Let $\|\widehat{\tau}_{z_{0}}\|_{L^{2},\widehat{h}_{z_{0}}}$ denote the $L^{2}$-norm of $\widehat{\tau}_{z_{0}}$ with respect to $\widehat{h}_{z_{0}}$ and $g_{0}$. We obtain
$|\widehat{\theta}_{z_{0}}|_{\widehat{h}_{z_{0}}}\leq B_{2}\epsilon_{1},\quad|R(\widehat{h}_{z_{0}})|_{\widehat{h}_{z_{0}}}\leq\epsilon_{0},\quad\|\widehat{\tau}_{z_{0}}\|_{L^{2},\widehat{h}_{z_{0}}}\leq\|\tau\|_{L^{2},h}.$
By using Theorem 2.32, Lemma 2.33, and the bootstrapping argument, we can show that there exists $B_{5}>0$, depending only on $B_{2}$, such that $|\widehat{\tau}_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}\leq B_{5}\|\tau\|_{L^{2},h}$ on $\{|\zeta|<1/4\}$. It implies that $|\tau|_{h,g_{0}}\leq\epsilon_{1}^{-1}B_{5}\|\tau\|_{L^{2},h}$ on $\mathcal{B}(r_{1})$.
2.5.4 Poisson equation
Let $s$ be a $C^{\infty}$-section of $E$ on $\mathcal{B}(r_{0})$.
Proposition 2.36
There exists $B_{10}>0$ depending only on $B_{1}$ and $r_{0}-r_{1}$ such that the following holds:
$\sup_{\mathcal{B}(r_{1})}\bigl{|}s\bigr{|}_{h}+\sup_{\mathcal{B}(r_{1})}\bigl{|}(\overline{\partial}_{E}+\theta)s\bigr{|}_{h,g_{0}}\leq B_{10}\|s\|_{L^{2},h,g_{0}}+B_{10}\sup_{\mathcal{B}(r_{2})}\bigl{|}(\overline{\partial}_{E}+\theta)^{\ast}_{h,g_{0}}(\overline{\partial}_{E}+\theta)s\bigr{|}_{h}.$ (25)
Proof   We use the notation in the proof of Proposition 2.35. We obtain the section $\widehat{s}_{z_{0}}=\psi_{z_{0}}^{-1}(s)$. By using Theorem 2.32, Lemma 2.33, Proposition 2.34 and the bootstrapping argument, we obtain the estimates for $|\widehat{s}_{z_{0}}|_{\widehat{h}_{z_{0}}}$ and |(¯$|(\overline{\partial}_{\widehat{E}_{z_{0}}}+\widehat{\theta}_{z_{0}})\widehat{s}_{z_{0}}|_{\widehat{h}_{z_{0}},g_{0}}$. Then, we obtain (25).
2.6 Real bilinear forms on complex vector spaces
This is a preliminary for §2.7.
2.6.1 Bilinear forms on real vector spaces
Let $V_{{\mathbb{R}}}$ be a finite dimensional ${\mathbb{R}}$-vector space. We set $V_{{\mathbb{R}}}^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{R}}}(V_{{\mathbb{R}}},{\mathbb{R}})$. Let $g_{{\mathbb{R}}}$ be a non-degenerate symmetric bilinear form on $V_{{\mathbb{R}}}$. It induces the isomorphism $\Psi_{g_{{\mathbb{R}}}}:V_{{\mathbb{R}}}\to V_{{\mathbb{R}}}^{\lor}$ by $\Psi_{g_{{\mathbb{R}}}}(u)(v)=g_{{\mathbb{R}}}(u,v)$. We obtain the induced symmetric form $g_{{\mathbb{R}}}^{\lor}$ on $V_{{\mathbb{R}}}^{\lor}$ by
$g_{{\mathbb{R}}}^{\lor}(a,b)=g_{{\mathbb{R}}}\bigl{(}\Psi_{g_{{\mathbb{R}}}}^{-1}(a),\Psi_{g_{{\mathbb{R}}}}^{-1}(b)\bigr{)}.$
Lemma 2.37
We have $\Psi_{g_{{\mathbb{R}}}}^{\lor}\circ\Psi_{g_{{\mathbb{R}}}}=\mathop{\rm id}\nolimits_{V_{{\mathbb{R}}}}$ and $(g_{{\mathbb{R}}}^{\lor})^{\lor}=g_{{\mathbb{R}}}$. If $g_{{\mathbb{R}}}$ is positive definite, $g_{{\mathbb{R}}}^{\lor}$ is also positive definite.
Let $\omega$ be a symplectic form on $V_{{\mathbb{R}}}$. It induces the isomorphism $\Psi_{\omega}:V_{{\mathbb{R}}}\to V_{{\mathbb{R}}}^{\lor}$ by $\Psi_{\omega}(u)(v)=\omega(u,v)$. We obtain the induced symplectic form $\omega^{\lor}$ on $V_{{\mathbb{R}}}^{\lor}$ by
$\omega^{\lor}(a,b)=\omega\bigl{(}\Psi_{\omega}^{-1}(a),\Psi_{\omega}^{-1}(b)\bigr{)}.$
Lemma 2.38
We have $\Psi_{\omega^{\lor}}\circ\Psi_{\omega}=-\mathop{\rm id}\nolimits_{V}$ and $(\omega^{\lor})^{\lor}=\omega$.
2.6.2 Real bilinear forms on complex vector spaces
Let $V$ be a finite dimensional ${\mathbb{C}}$-vector space. Let $V_{{\mathbb{R}}}$ denote the underlying ${\mathbb{R}}$-vector space. We set $V^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{C}}}(V,{\mathbb{C}})$ and $V_{{\mathbb{R}}}^{\lor}=\mathop{\rm Hom}\nolimits_{{\mathbb{R}}}(V_{{\mathbb{R}}},{\mathbb{R}})$. The multiplication of $\sqrt{-1}$ on $V$ induces the ${\mathbb{R}}$-automorphism $I$ on $V_{{\mathbb{R}}}$ such that $I^{2}=-\mathop{\rm id}\nolimits_{V_{{\mathbb{R}}}}$. It induces an ${\mathbb{R}}$-linear automorphism $I^{\lor}$ on $V_{{\mathbb{R}}}^{\lor}$ by $I^{\lor}(f)(v)=f(Iv)$.
Let us recall the identification of $V_{{\mathbb{R}}}^{\lor}$ and $V^{\lor}$. There exists the decomposition
$V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}=\mathop{\rm Hom}\nolimits_{{\mathbb{R}}}(V_{{\mathbb{R}}},{\mathbb{C}})=V^{\lor}\oplus\overline{V^{\lor}},$
where $V^{\lor}$ is the $\sqrt{-1}$-eigen space of $I^{\lor}$, and $\overline{V^{\lor}}$ is the $-\sqrt{-1}$-eigen space of $I^{\lor}$. We obtain the isomorphism of ${\mathbb{R}}$-vector spaces $\Phi:V_{{\mathbb{R}}}^{\lor}\to V^{\lor}$ as the composition of the inclusion $V_{{\mathbb{R}}}^{\lor}\to V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}$ and the projection $V_{{\mathbb{R}}}^{\lor}\otimes{\mathbb{C}}=V^{\lor}\oplus\overline{V^{\lor}}\to V^{\lor}$. It is explicitly described as
$\Phi(f)=\frac{1}{2}\bigl{(}f-\sqrt{-1}I^{\lor}(f)\bigr{)}.$
Remark 2.39
If we regard $V_{{\mathbb{R}}}^{\lor}$ with $I^{\lor}$ as a ${\mathbb{C}}$-vector space, then $\Phi$ is an isomorphism of ${\mathbb{C}}$-vector spaces.
We can check the following lemmas by direct computations.
Lemma 2.40
Let $g_{{\mathbb{R}}}$ be a non-degenerate symmetric bilinear form on $V_{{\mathbb{R}}}$. Let $\omega$ be a symplectic form on $V_{{\mathbb{R}}}$.
  • If $I$ is an isometry with respect to $g_{{\mathbb{R}}}$, we have $\Psi_{g_{{\mathbb{R}}}}\circ I=-I^{\lor}\circ\Psi_{g_{{\mathbb{R}}}}$. If $I$ is an isometry with respect to $\omega$, we have $\Psi_{\omega}\circ I=-I^{\lor}\circ\Psi_{\omega}$.
  • Suppose that $I$ is an isometry with respect to $g_{{\mathbb{R}}}$, and that $\omega(u,v)=g_{{\mathbb{R}}}(u,Iv)$ for any $u,v$. Then, we obtain $\omega^{\lor}(u^{\lor},v^{\lor})=-g_{{\mathbb{R}}}^{\lor}(u^{\lor},I^{\lor}v^{\lor})$. In particular, if $\omega(\cdot,I\cdot)$ is positive definite, $-\omega^{\lor}(\cdot,I^{\lor},\cdot)$ is positive definite.
2.7 Bilinear forms on the first cohomology group of a compact Riemann surface
Let $\Sigma$ be a compact Riemann surface. There exists a natural symplectic form $\omega_{\Sigma}$ on $H^{1}(\Sigma,{\mathbb{R}})$ given by
$\omega_{\Sigma}(a,b)=\int_{\Sigma}a\wedge b.$
There exists the Hodge decomposition:
$H^{1}(\Sigma,{\mathbb{R}})\otimes{\mathbb{C}}=H^{1}(\Sigma,\mathcal{O}_{\Sigma})\oplus H^{0}(\Sigma,K_{\Sigma}).$
The projections induce isomorphisms of ${\mathbb{R}}$-vector spaces
${\tt F}_{1}:H^{1}(\Sigma,{\mathbb{R}})\simeq H^{1}(\Sigma,\mathcal{O}_{\Sigma}),\quad\quad{\tt F}_{2}:H^{1}(\Sigma,{\mathbb{R}})\simeq H^{0}(\Sigma,K_{\Sigma}).$
2.7.1 Bilinear forms on $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$
Let $\omega_{1}$ be the real symplectic form on $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$ given by
$\omega_{1}(a_{1},a_{2})=-\omega_{\Sigma}({\tt F}_{1}^{-1}(a_{1}),{\tt F}_{1}^{-1}(a_{2})).$
There exist harmonic representative $\eta_{i}$ of $a_{i}$, i.e., $\eta_{i}$ are $(0,1)$-forms such that $\partial_{\Sigma}\eta_{i}=0$, and $[\eta_{i}]=a_{i}$ in $H^{1}(\Sigma,\mathcal{O}_{\Sigma})$. We have
${\tt F}_{1}^{-1}([\eta_{i}])=[\eta_{i}+\overline{\eta_{i}}].$
Therefore,
$\omega_{1}([\eta_{1}],[\eta_{2}])=-\int_{\Sigma}(\eta_{1}+\overline{\eta_{1}})\wedge(\eta_{2}+\overline{\eta_{2}})=-\int_{\Sigma}\bigl{(}\eta_{1}\overline{\eta_{2}}+\overline{\eta_{1}}\eta_{2}\bigr{)}.$
The associated positive definite symmetric bilinear form $g_{1{\mathbb{R}}}$ is given as follows:
$g_{1{\mathbb{R}}}([\eta_{1}],[\eta_{2}])=\omega_{1}([\sqrt{-1}\eta_{1}],[\eta_{2}])=-\sqrt{-1}\int_{\Sigma}(\eta_{1}\overline{\eta}_{2}-\overline{\eta_{1}}\eta_{2})=-\sqrt{-1}\int_{\Sigma}(\eta_{1}\overline{\eta}_{2}+\eta_{2}\overline{\eta_{1}}).$
The associated Hermitian metric $g_{1}=g_{1{\mathbb{R}}}+\sqrt{-1}\omega_{1}$ is given as follows:
$g_{1}([\eta_{1}],[\eta_{2}])=-2\sqrt{-1}\int_{\Sigma}\eta_{1}\overline{\eta}_{2}.$
2.7.2 Bilinear forms on $H^{0}(\Sigma,K_{\Sigma})$
Let $\omega_{2}$ be the real symplectic form on $H^{0}(\Sigma,K_{\Sigma})$ given by
$\omega_{2}(a_{1},a_{2})=\omega_{\Sigma}({\tt F}_{2}^{-1}(a_{1}),{\tt F}_{2}^{-1}(a_{2})).$
We may naturally regard $a_{i}$ as holomorphic $1$-forms on $\Sigma$. We have
${\tt F}_{2}^{-1}(a_{i})=a_{i}+\overline{a_{i}}.$
Therefore,
$\omega_{2}(a_{1},a_{2})=\int_{\Sigma}(a_{1}+\overline{a_{1}})\wedge(a_{2}+\overline{a_{2}})=\int_{\Sigma}\bigl{(}a_{1}\overline{a_{2}}+\overline{a_{1}}a_{2}\bigr{)}.$
The associated positive definite symmetric bilinear form $g_{2{\mathbb{R}}}$ is given as follows.
$g_{2{\mathbb{R}}}(a_{1},a_{2})=\omega_{2}(\sqrt{-1}a_{1},a_{2})=\sqrt{-1}\int_{\Sigma}(a_{1}\overline{a_{2}}-\overline{a_{1}}a_{2})=\sqrt{-1}\int_{\Sigma}(a_{1}\overline{a}_{2}+a_{2}\overline{a_{1}}).$
The associated hermitian metric $g_{2}$ is given as follows:
$g_{2}(a_{1},a_{2})=2\sqrt{-1}\int_{\Sigma}a_{1}\overline{a}_{2}.$
2.7.3 Duality
By the Serre duality, the following pairing is perfect.
$H^{1}(\Sigma,\mathcal{O}_{\Sigma})\otimes H^{0}(\Sigma,K_{\Sigma})\longrightarrow{\mathbb{C}},\quad\quad(a,b)\longmapsto-\int_{\Sigma}a\wedge b.$ (26)
By this perfect pairing, we may regard $H^{0}(\Sigma,K_{\Sigma})\simeq H^{1}(\Sigma,\mathcal{O}_{\Sigma})^{\lor}$. We also have the isomorphism $H^{1}(\Sigma,\mathcal{O}_{\Sigma})^{\lor}\simeq H^{1}(\Sigma,\mathcal{O}_{\Sigma})_{{\mathbb{R}}}^{\lor}$ as explained in §2.6.2. By the composition, we obtain the following isomorphism of ${\mathbb{R}}$-vector spaces:
$H^{0}(\Sigma,K_{\Sigma})_{{\mathbb{R}}}\simeq H^{1}(\Sigma,\mathcal{O}_{\Sigma})_{{\mathbb{R}}}^{\lor}.$ (27)
We have the bilinear forms $g_{1{\mathbb{R}}}^{\lor}$ and $\omega_{1}^{\lor}$ on $H^{1}(\Sigma,\mathcal{O}_{\Sigma})_{{\mathbb{R}}}^{\lor}$ as explained in §2.6.1.
Lemma 2.41
Under the isomorphism (27), we have $\omega_{2}=-\omega_{1}^{\lor}$ and $g_{2{\mathbb{R}}}=g_{1{\mathbb{R}}}^{\lor}$.
Proof   Let $\widetilde{\Psi}_{\omega_{1}}:H^{1}(\Sigma,\mathcal{O}_{\Sigma})_{{\mathbb{R}}}\simeq H^{0}(\Sigma,K_{\Sigma})_{{\mathbb{R}}}$ denote the composite of $\Psi_{\omega_{1}}$ and the inverse of (27). (See §2.6.1 for $\Psi_{\omega_{1}}$.) For harmonic $(0,1)$-forms $\eta_{i}$ $(i=1,2)$, the following holds:
$\frac{1}{2}\Bigl{(}\Psi_{\omega_{1}}(\eta_{1})-\sqrt{-1}I^{\lor}\Psi_{\omega_{1}}(\eta_{1})\Bigr{)}(\eta_{2})=-\int_{\Sigma}\overline{\eta}_{1}\eta_{2}.$
We obtain $\widetilde{\Psi}_{\omega_{1}}(\eta_{1})=-\overline{\eta}_{1}$.
Let $a_{i}\in H^{0}(\Sigma,K_{\Sigma})$ $(i=1,2)$. We obtain
$-\omega_{1}^{\lor}(a_{1},a_{2})=-\omega_{1}\bigl{(}\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{1}),\widetilde{\Psi}_{\omega_{1}}^{-1}(a_{2})\bigr{)}=-\omega_{1}(\overline{a}_{1},\overline{a}_{2})=\int_{\Sigma}\bigl{(}\overline{a}_{1}a_{2}+a_{1}\overline{a}_{2}\bigr{)}=\omega_{2}(a_{1},a_{2}).$
We obtain $g_{2{\mathbb{R}}}=g_{1{\mathbb{R}}}^{\lor}$ from Lemma 2.40.
2.8 Estimate for the family of the associated Laplacian operators
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. Let $(E,\theta)$ be a generically regular semisimple Higgs bundle of degree 00 whose spectral curve is smooth. We fix a flat metric $h_{\det(E)}$ of $\det(E)$. For each $t>0$, there exists a harmonic metric $h_{t}$ of $(E,t\theta)$. We recall that the sequence $h_{t}$ is convergent to the limiting configuration $h_{\infty}$ in the $C^{\infty}$-sense locally on $X^{\circ}$.
Let $\mathop{\rm End}\nolimits_{0}(E)$ denote the trace free part of $\mathop{\rm End}\nolimits(E)$. Let (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}$ denote the formal adjoint of ¯$\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{t}$ and $g_{X}$. We consider the operators $\Delta_{t}=(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{t}}\circ(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)$ on $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$. For any $s$ of $A^{0}(\mathop{\rm End}\nolimits_{0}(E))$, let $\|s\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of $s$ with respect to $h_{t}$ and $g_{X}$.
Proposition 2.42
There exists $C>0$ such that $\|\Delta_{t}(s)\|_{L^{2},h_{t}}\geq C\|s\|_{L^{2},h_{t}}$ for any $t\geq 1$ and $s\in A^{0}(\mathop{\rm End}\nolimits_{0}(E))$.
Proof   If $s\in A^{0}(\mathop{\rm End}\nolimits_{0}(E))$ satisfies $\Delta_{t}s=0$, it is standard to obtain (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)s=0$. Because $(E,t\theta)$ is stable, we obtain $s=0$.
Let $a_{t}>0$ denote the first eigenvalue of $\Delta_{t}$. It is enough to prove that $\inf_{t\geq 1}a_{t}>0$. Suppose that there exists a subsequence $t(i)\to\infty$ such that $a_{t(i)}\to 0$, and we shall deduce a contradiction. There exists a sequence $s_{i}\in A^{0}(\mathop{\rm End}\nolimits_{0}(E))$ such that (i) $\|s_{i}\|_{L^{2},h_{t(i)}}=1$, (ii) $\Delta_{t(i)}s_{i}=a_{t(i)}\cdot s_{i}$. We obtain
$\|t[\theta,s_{i}]\|_{L^{2},h_{t(i)}}+\|t[\theta^{\dagger}_{h_{t(i)}},s_{i}]\|_{L^{2},h_{t(i)}}+\|\overline{\partial}_{E}(s_{i})\|_{L^{2},h_{t(i)}}+\|\partial_{E,h_{t(i)}}(s_{i})\|_{L^{2},h_{t(i)}}\to 0.$
By taking a subsequence, we may assume that the sequence $s_{i}$ is weakly convergent in $L_{1}^{2}$ locally on $X^{\circ}$. Let $s_{\infty}$ denote the limit. The sequence $s_{i}$ $(i=1,2,\ldots)$ is convergent to $s_{\infty}$ in $L^{p}$ for any $p>1$ locally on $X^{\circ}$. We obtain $\theta(s_{\infty})=0$ and ¯$\overline{\partial}_{E}(s_{\infty})=0$. We also have $\|s_{\infty}\|_{L^{2},h_{\infty}}\leq 1$. Because $\mathop{\rm tr}\nolimits s_{i}=0$, we obtain $\mathop{\rm tr}\nolimits s_{\infty}=0$.
We set $b_{i}:=\max_{X}|s_{i}|_{h_{t(i)}}$. Because $\|s_{i}\|_{L^{2},h_{t(i)}}=1$, there exists $C_{1}>0$ such that $b_{i}\geq C_{1}$.
Lemma 2.43
The sequence $b_{i}$ is bounded.
Proof   We set $\widetilde{s}_{i}=b_{i}^{-1}s_{i}$ and $h_{i}=h_{t(i)}$. We take $P_{i}\in X$ such that $|\widetilde{s}_{i|P_{i}}|_{h_{i}}=1$. By taking a subsequence, we may assume that the sequence $P_{i}$ $(i=1,2,\ldots)$ is convergent. Let $P_{\infty}$ denote the limit.
Let $U$ be a relatively compact neighbourhood of $P_{\infty}$ in $X$ with a holomorphic coordinate $z$. Let (¯$(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}$ denote the formal adjoint of ¯$\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta$ with respect to $h_{i}$ and the Euclidean metric $dz\,d\overline{z}$. By Lemma 2.45 below, we obtain
$-\partial_{z}\partial_{\overline{z}}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2\bigl{|}(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)^{\ast}_{h_{i}}(\overline{\partial}_{E}+t\mathop{\rm ad}\nolimits\theta)(\widetilde{s}_{i})\bigr{|}_{h_{i}}\cdot|\widetilde{s}_{i}|_{h_{i}}=2a_{t(i)}g_{U}|\widetilde{s}_{i}|_{h_{i}}^{2}\leq 2a_{t(i)}g_{U}.$
Here, $g_{U}$ is an ${\mathbb{R}}_{>0}$-valued $C^{\infty}$-function on $U$ such that $g_{U}$ and $g_{U}^{-1}$ are bounded. There exists a bounded $C^{\infty}$-function $\beta_{U}$ on $U$ such that $\partial_{z}\partial_{\overline{z}}\beta_{U}=g_{U}$. We obtain
$-\partial_{z}\partial_{\overline{z}}\bigl{(}|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}\bigr{)}\leq 0.$
Let us consider the case where $P_{\infty}\in X^{\circ}$. We may assume that $U$ is relatively compact in $X^{\circ}$. There exists $\epsilon>0$ such that $B(P_{\infty},\epsilon)=\bigl{\{}P\in U\,\big{|}\,|z(P)-z(P_{\infty})|<\epsilon\bigr{\}}$ is relatively compact subset in $U$. By the mean value property of the subharmonic function $|\widetilde{s}_{i}|^{2}_{h_{i}}-2a_{t(i)}\beta_{U}$, we obtain
$(|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U})(P)\leq\frac{1}{\pi\epsilon^{2}}\int_{B(P_{\infty},\epsilon)}\bigl{(}|\widetilde{s}_{i}|^{2}_{h_{i}}-2a_{t(i)}\beta_{U}\bigr{)}\frac{|dz\,d\overline{z}|}{2}.$
Because $a_{t(i)}\to 0$, there exists $C_{4}>0$ such that the following holds:
$C_{4}\leq\int_{B(P_{\infty},\epsilon)}|\widetilde{s}_{i}|_{h_{i}}^{2}=b_{i}^{-2}\int_{B(P_{\infty},\epsilon)}|s_{i}|_{h_{i}}^{2}.$
Because $s_{i}$ is convergent to $s_{\infty}$ in $L^{2}$ on $U$, there exists $C_{5}>0$ such that $b_{i}\leq C_{5}$.
Let us consider the case $P_{\infty}\in D(\theta)$. Let $U^{\prime}$ be a relatively compact neighbourhood of $P_{\infty}$ in $U$ such that $U^{\prime}\cap D(\theta)=\{P_{\infty}\}$. By applying the maximum principle to the subharmonic function $|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}$, there exists $Q_{i}\in\partial U^{\prime}$ such that
$\bigl{(}|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}\bigr{)}(P_{i})\leq\bigl{(}|\widetilde{s}_{i}|_{h_{i}}^{2}-2a_{t(i)}\beta_{U}\bigr{)}(Q_{i}).$
Because $a_{t(i)}\to 0$, there exists $C_{6}>0$ such that
$|\widetilde{s}_{i|Q_{i}}|_{h_{i}}^{2}\geq C_{6}.$
By taking a subsequence, we may assume that the sequence $Q_{i}$ is convergent. Let $Q_{\infty}\in\partial U^{\prime}$ denote the limit. There exists $\epsilon>0$ such that $B(Q_{\infty},\epsilon)=\bigl{\{}P\in U\,\big{|}\,|z(P)-z(Q_{\infty})|<\epsilon\bigr{\}}$ is relatively compact in $U\setminus\{P_{\infty}\}$. By the mean value property of the subharmonic function $|\widetilde{s}_{i}|_{h_{i}}^{2}-a_{t(i)}\beta_{U}$, there exists $C_{7}>0$ such that
$C_{7}\leq\int_{B(Q_{\infty},\epsilon)}|\widetilde{s}_{i}|_{h_{i}}^{2}=b_{i}^{-2}\int_{B(Q_{\infty},\epsilon)}|s_{i}|_{h_{i}}^{2}.$
Because $s_{i}$ is convergent in $L^{2}$ on $B(Q_{\infty},\epsilon)$, the sequence $b_{i}$ is bounded. Thus, we obtain Lemma 2.43.
By the Lebesgue theorem and the boundedness of the sequence $b_{i}$, we obtain the following lemma.
Lemma 2.44
$s_{\infty}$ is a non-zero bounded section of $E$ on $X^{\circ}$.
Hence, $s_{\infty}$ is an automorphism of the filtered Higgs bundle $(\mathcal{P}^{h_{\infty}}_{\ast}E,\theta)$. (See [21] for filtered Higgs bundles, for example.) Because $\Sigma_{\theta}$ is connected, $(\mathcal{P}^{h_{\infty}}_{\ast}E,\theta)$ is stable, and hence we obtain $s_{\infty}=\alpha\cdot\mathop{\rm id}\nolimits_{E}$ for some $\alpha\in{\mathbb{C}}$. Because $\mathop{\rm tr}\nolimits s_{\infty}=0$, we obtain $\alpha=0$, i.e., $s_{\infty}=0$. It contradicts $s_{\infty}\neq 0$. Thus, we obtain a contradiction as desired, and the proof of Proposition 2.42 is completed.
We recall the following general lemma.
Lemma 2.45
Let $(E,\theta,h)$ be a harmonic bundle on $X$. Let $U$ be an open subset of $X$ with a holomorphic coordinate $z$. Let $f$ be the automorphism of $E_{|U}$ determined by $\theta=f\,dz$. Let $\nabla$ denote the Chern connection of $(E,h)$. Then, for any $s\in A^{0}(U,E)$, we have
$-\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=2\mathop{\rm Re}\nolimits\bigl{(}(-\nabla_{z}\nabla_{\overline{z}}+f_{h}^{\dagger}\circ f)s,s\bigr{)}-|\nabla_{z}s|_{h}^{2}-|\nabla_{\overline{z}}s|_{h}^{2}-|f(s)|_{h}^{2}-|f^{\dagger}_{h}(s)|_{h}^{2}.$ (28)
Proof   Note that $\bigl{(}\nabla_{z}\nabla_{\overline{z}}-\nabla_{\overline{z}}\nabla_{z}\bigr{)}+[f,f^{\dagger}_{h}]=0$. We obtain
$-\partial_{z}\partial_{\overline{z}}|s|_{h}^{2}=-h\bigl{(}\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}-h\bigl{(}s,\nabla_{\overline{z}}\nabla_{z}s\bigr{)}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =-h\bigl{(}\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}-h\bigl{(}s,\nabla_{z}\nabla_{\overline{z}}s\bigr{)}-h\bigl{(}s,[f,f^{\dagger}_{h}]s\bigr{)}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl{(}-\nabla_{z}\nabla_{\overline{z}}s,s\bigr{)}+2h\bigl{(}s,f^{\dagger}_{h}\circ f(s)\bigr{)}-\bigl{|}f(s)\bigr{|}_{h}^{2}-\bigl{|}f^{\dagger}_{h}(s)\bigr{|}_{h}^{2}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}\\ =2\mathop{\rm Re}\nolimits h\bigl{(}(-\nabla_{z}\nabla_{\overline{z}}+f^{\dagger}_{h}f)s,s\bigr{)}-\bigl{|}f(s)\bigr{|}_{h}^{2}-\bigl{|}f^{\dagger}_{h}(s)\bigr{|}_{h}^{2}-\bigl{|}\nabla_{\overline{z}}s\bigr{|}_{h}^{2}-\bigl{|}\nabla_{z}s\bigr{|}_{h}^{2}.$ (29)
Thus, we obtain (28).
3 Horizontal deformations
3.1 Line bundles on spectral curves
3.1.1 Spectral curves
We recall the spectral curves by following [1]. Let $X$ be a compact Riemann surface. In this section, we prefer to use the notation $K_{X}$ to denote the sheaf of holomorphic $1$-forms because we consider $K_{X}^{j/2}$. We set
$\mathbb{P}=\mathop{\rm Proj}\nolimits(\mathop{\rm Sym}\nolimits(\mathcal{O}_{X}\oplus K_{X}^{-1})),$
i.e., $\mathbb{P}$ denotes the projective completion of $K_{X}$. Let $\pi_{\mathbb{P}}:\mathbb{P}\to X$ denote the projection. Let $\mathcal{O}_{\mathbb{P}}(1)$ denote the tautological bundle. We have $\pi_{\mathbb{P}\ast}(\mathcal{O}_{\mathbb{P}}(1))=\mathcal{O}_{X}\oplus K_{X}^{-1}$. Let $y$ denote the section of $\mathcal{O}_{\mathbb{P}}(1)$ induced by the section $1$ of $\mathcal{O}_{X}\subset\mathcal{O}_{X}\oplus K_{X}^{-1}$. Let $x$ denote the section of $\pi_{\mathbb{P}}^{\ast}K_{X}\otimes\mathcal{O}_{\mathbb{P}}(1)$ induced by the section $1$ of $\mathcal{O}_{X}\subset K_{X}\otimes(\mathcal{O}_{X}\oplus K_{X}^{-1})$.
We set ${\boldsymbol{A}}_{X,n}=\bigoplus_{j=1}^{n}H^{0}(X,K_{X}^{j})$. For any $s=(s_{1},\ldots,s_{n})\in{\boldsymbol{A}}_{X,n}$, we obtain the section
$\xi_{s}=x^{n}+\sum_{j=1}^{n}(-1)^{j}s_{j}y^{j}x^{n-j}\in H^{0}\bigl{(}\mathbb{P},\pi_{\mathbb{P}}^{\ast}K_{X}^{n}\otimes\mathcal{O}_{\mathbb{P}}(n)\bigr{)}.$
Let $Z_{s}$ be the subscheme of $\mathbb{P}$ obtained as the 00-set of $\xi_{s}$. It is contained in $T^{\ast}X\subset\mathbb{P}$. Let ${\boldsymbol{A}}^{\prime}_{X,n}$ denote the subset of $s\in{\boldsymbol{A}}^{\prime}_{X,n}$ such that $Z_{s}$ are smooth. Let $\pi_{s}:Z_{s}\to X$ denote the projection. For each $Q\in Z_{s}$, let $r_{Q}$ denote the ramification index of $\pi_{s}$ at $Q$. We obtain the effective divisor $\widetilde{D}_{s}=\sum(r_{Q}-1)Q$. We have $K_{Z_{s}}=\pi^{\ast}(K_{X})(\widetilde{D}_{s})$. We recall the following.
Lemma 3.1
Let $g_{X}$ denote the genus of $X$.
  • For any $s\in{\boldsymbol{A}}^{\prime}_{X,n}$, the genus of $Z_{s}$ is $n^{2}(g_{X}-1)+1$.
  • $\deg(\widetilde{D}_{s})=2n(n-1)(g_{X}-1)$.
Let $Z\subset{\boldsymbol{A}}^{\prime}_{X,n}\times T^{\ast}X$ denote the universal curve. The natural morphism $Z\to{\boldsymbol{A}}^{\prime}_{X,n}$ is smooth.
3.1.2 Singular flat metrics
We consider $s\in{\boldsymbol{A}}^{\prime}_{X,n}$ and a holomorphic line bundle $L$ on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. We set $Z_{s}^{\circ}=Z_{s}\setminus|\widetilde{D}_{s}|$. Because
$\deg(L)-\sum_{Q}\frac{1}{2}(r_{Q}-1)=\deg(L)-\frac{1}{2}\deg(\widetilde{D}_{s})=\deg(L)-n(n-1)(g_{X}-1)=0,$
we obtain the following lemma.
Lemma 3.2
There exists a Hermitian metric $h_{L}$ of $L_{|Z_{s}^{\circ}}$ satisfying the following conditions.
  • The Chern connection $\nabla_{L}$ of $(L_{|Z_{s}^{\circ}},h_{L})$ is flat.
  • For each $Q\in Z_{s}$, there exists a frame $v_{Q}$ of $L$ around $Q$ such that $h_{L}(v_{Q},v_{Q})=|\zeta_{Q}|^{-r_{Q}+1}$, where $\zeta_{Q}$ is a local coordinate around $Q$ such that $\zeta_{Q}(Q)=0$, and $r_{Q}$ is the ramification index of $\pi_{s}$ at $Q$.
Such a Hermitian metric is unique up to the multiplication of positive constants.
Note that for each $Q\in Z_{s}$ there exists a neighbourhood $(Z_{s})_{Q}$ and a pairing $C_{Q}:L_{|(Z_{s})_{Q}}\otimes L_{|(Z_{s})_{Q}}\longrightarrow\mathcal{O}_{Z_{s}}((r_{Q}-1)Q)$ such that $C_{Q}$ is compatible with $h_{L}$, i.e., $|C_{Q}(v_{Q},v_{Q})|=|\zeta_{Q}|^{-r_{Q}+1}$ for $v_{Q}$ as above. We set $U(1)=\bigl{\{}a\in{\mathbb{C}}\,\big{|}\,|a|=1\bigr{\}}$. The monodromy representation induces a homomorphism $H_{1}(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}})\to U(1)$ denoted by $\kappa_{L}$.
Lemma 3.3
Let $L_{i}$ $(i=1,2)$ be holomorphic line bundles on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. Let $U_{j}$ $(j=1,\ldots,m)$ be simply connected open subsets of $X$ such that $U_{j}\cap U_{k}=\emptyset$ $(j\neq k)$ and that $|\widetilde{D}_{s}|\subset\bigcup U_{j}$. Let
$\iota:H_{1}\left(Z_{s}\setminus\bigcup U_{j},{\mathbb{Z}}\right)\longrightarrow H_{1}\left(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}}\right)$
denote the morphism induced by the inclusion. If $\kappa_{L_{1}}\circ\iota=\kappa_{L_{2}}\circ\iota$, there exists an isomorphism $L_{1}\simeq L_{2}$.
Proof   For any $Q\in|\widetilde{D}_{s}|$, let $\gamma_{Q}$ be a loop around $Q$. We obtain $\kappa_{L_{i}}(\gamma_{Q})=(-1)^{r_{Q}-1}$ by the definition of $h_{L_{i}}$. Hence, the assumption of Lemma 3.3 implies $\kappa_{L_{1}}=\kappa_{L_{2}}$. We obtain the isomorphism of the flat bundles $(L_{1|Z_{s}\setminus|\widetilde{D}_{s}|},\nabla_{L_{1}})\simeq(L_{2|Z_{s}\setminus|\widetilde{D}_{s}|},\nabla_{L_{2}})$. It extends to an isomorphism of holomorphic line bundles $L_{1}\simeq L_{2}$.
3.1.3 A special case
We fix a line bundle $K_{X}^{1/2}$ and an isomorphism $(K_{X}^{1/2})^{2}\simeq K_{X}$. We have
$\deg(\pi_{s}^{\ast}K_{X}^{(n-1)/2})=n(n-1)(g_{X}-1)=\frac{1}{2}\deg(\widetilde{D}_{s}).$ (30)
Proposition 3.4
The image of $\kappa_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}:H_{1}(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}})\to U(1)$ is contained in {±$\{\pm 1\}$.
Proof   Recall that $Z_{s}$ is the 00-set of $\xi_{s}\in H^{0}(\mathbb{P},\pi_{\mathbb{P}}^{\ast}K_{X}^{n}\otimes\mathcal{O}_{\mathbb{P}}(n))$, and that the line bundle $\mathcal{O}_{\mathbb{P}}(n)$ is trivialized by the section $\pi_{\mathbb{P}}^{\ast}(y)$ on a neighbourhood of $Z_{s}$. (See §3.1.1 for the section $y$.) Hence, there exists a natural isomorphism between $\pi_{s}^{\ast}K_{X}^{n}$ and the normal bundle of $Z_{s}\subset T^{\ast}X$. By the symplectic structure of $T^{\ast}X$, we obtain the isomorphism $\pi_{s}^{\ast}K_{X}^{n}\simeq K_{Z_{s}}$. Hence, there exists a natural isomorphism
$C_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}:\pi_{s}^{\ast}(K_{X}^{(n-1)/2})\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})\simeq\pi_{s}^{\ast}(K_{X}^{n-1})\simeq K_{Z_{s}}\otimes\pi_{s}^{\ast}(K_{X}^{-1})\simeq\mathcal{O}_{Z_{s}}(\widetilde{D}_{s}).$ (31)
For any local section $u$ of $\pi_{s}^{\ast}K_{X}^{(n-1)/2}$, we define
$h_{\pi_{s}^{\ast}K_{X}^{(n-1)/2}}(u,u)=|C_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}(u,u)|.$
In this way, we obtain a singular flat metric $h_{\pi_{s}^{\ast}K_{X}^{(n-1)/2}}$ as in Lemma 3.2. Because $h_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}$ is compatible with the non-degenerate symmetric bilinear product, we obtain the claim of Proposition 3.4. (See [23, Lemma 2.7].)
3.1.4 Some representation
Let $L$ be a holomorphic line bundle on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. We obtain the holomorphic line bundle $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$ of degree 00. It is equipped with the unique unitary flat connection $\nabla_{\widehat{L}}$ whose $(0,1)$-part equals the holomorphic structure of $\widehat{L}$. The monodromy representation of $\nabla_{\widehat{L}}$ induces a homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$.
Remark 3.5
Let $(K_{X}^{1/2})^{\prime}$ be another square root of $K_{X}$. We obtain the holomorphic line bundle $M=(K_{X}^{1/2})^{-1}\otimes(K_{X}^{1/2})^{\prime}$ of degree 00 on $X$. It is equipped with a unitary flat connection $\nabla_{M}$ which induces $\kappa_{M}:H_{1}(X,{\mathbb{Z}})\to\{\pm 1\}$. It induces $\pi_{s}^{\ast}(\kappa_{M}):H_{1}(Z_{s},{\mathbb{Z}})\to\{\pm 1\}$. Let $\widehat{\kappa}^{\prime}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$ be the homomorphism obtained from $L$ and $(K_{X}^{1/2})^{\prime}$. Then, we have $\widehat{\kappa}_{L}=\widehat{\kappa}^{\prime}_{L}\otimes\pi_{s}^{\ast}(\kappa_{M})$.
There exists a natural isomorphism $L\simeq\widehat{L}\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})$. For $\gamma\in H_{1}(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}})$, we have
$\kappa_{L}(\gamma)=\widehat{\kappa}_{L}(\gamma)\cdot\kappa_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}(\gamma)\in U(1).$
3.2 Horizontal deformation of line bundles
3.2.1 Horizontal family of line bundles
Let $B$ be a simply connected Hausdorff space with a base point $b_{0}$. Let $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ be a continuous map. We set $Z_{b}:=Z_{\varphi(b)}$. Let $L_{b}$ $(b\in B)$ be holomorphic line bundles on $Z_{b}$ of degree $n(n-1)(g_{X}-1)$. For any $b_{1},b_{2}\in B$, we choose a continuous path $\gamma$ in $B$ such that $\gamma(0)=b_{1}$ and $\gamma(1)=b_{2}$, and then we obtain the isomorphisms $H_{1}(Z_{b_{1}},{\mathbb{Z}})\simeq H_{1}(Z_{b_{2}},{\mathbb{Z}})$ along the path $\varphi\circ\gamma$, which is independent of the choice of $\gamma$.
Definition 3.6
We say that the family $L_{b}$ $(b\in B)$ is horizontal if $\widehat{\kappa}_{L_{b}}:H_{1}(Z_{b},{\mathbb{Z}})\to U(1)$ are independent of $b\in B$. This condition is independent of the choice of $K_{X}^{1/2}$.
The following lemma is obvious.
Lemma 3.7
Let $N$ be any holomorphic line bundle on $X$ with an isomorphism $N^{2}\simeq K_{X}$. Then, the family $\pi_{\varphi(b)}^{\ast}(N^{n-1})$ $(b\in B)$ is horizontal.
The following lemma is well known.
Lemma 3.8
Let $L$ be any holomorphic line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$. There exists a horizontal family $L_{b}$ $(b\in B)$ such that $L_{b_{0}}\simeq L$. If $L_{b}^{\prime}$ $(b\in B)$ is another horizontal family such that $L^{\prime}_{b_{0}}\simeq L$, then there exists $L_{b}\simeq L_{b}^{\prime}$ for each $b\in B$.
Proof   We set $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$, which is a holomorphic line bundle of degree 00. We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$. It induces homomorphisms $H_{1}(Z_{b},{\mathbb{Z}})\to U(1)$ and unitary flat line bundles $(\widehat{L}_{b},\nabla_{\widehat{L}_{b}})$ on $Z_{b}$ $(b\in B)$. Let $\widehat{L}_{b}$ denote the underlying holomorphic line bundle. We set $L_{b}=\widehat{L}_{b}\otimes\pi_{\varphi(b)}^{\ast}(K_{X}^{(n-1)/2})$. Then, $L_{b}$ $(b\in B)$ is a horizontal family.
Let $L^{\prime}_{b}$ $(b\in B)$ be another horizontal family such that $L^{\prime}_{b_{0}}\simeq L$. Because $\widehat{\kappa}_{L_{b}}=\widehat{\kappa}_{L^{\prime}_{b}}$ for any $b\in B$, there exist isomorphisms $L_{b}\otimes\pi_{\varphi(b)}^{\ast}(K_{X}^{(n-1)/2})^{-1}\simeq L^{\prime}_{b}\otimes\pi_{\varphi(b)}^{\ast}(K_{X}^{(n-1)/2})^{-1}$, and hence $L_{b}\simeq L_{b}^{\prime}$.
3.2.2 Complex analytic property of horizontal families
Let $B$ be a multi-disc with the base point $b_{0}=(0,\ldots,0)$. Let $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ be a holomorphic map. We set $Z_{B}:=Z\times_{{\boldsymbol{A}}^{\prime}_{X,n}}B$. The natural morphism $q_{0}:Z_{B}\to B$ is smooth. We use the notation $\widetilde{D}_{b}$ and $\pi_{b}$ instead of $\widetilde{D}_{\varphi(b)}$ and $\pi_{\varphi(b)}$, respectively. Let $\pi_{B}:Z_{B}\to B\times X$ be the induced morphism. For any holomorphic line bundle $\mathcal{L}$ on $Z_{B}$, we set $\mathcal{L}_{b}=\mathcal{L}_{|Z_{b}}$ $(b\in B)$. If the family $\mathcal{L}_{b}$ $(b\in B)$ is horizontal, we say that $\mathcal{L}$ is horizontal.
Lemma 3.9
Let $L$ be any holomorphic line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$. Then, there exists a horizontal holomorphic line bundle $\mathcal{L}$ on $Z_{B}$ such that $\mathcal{L}_{b_{0}}\simeq L$. If $\mathcal{L}^{\prime}$ is another horizontal holomorphic line bundle on $Z_{B}$ such that $\mathcal{L}^{\prime}_{b_{0}}\simeq L$, then there exists a holomorphic isomorphism $\mathcal{L}\simeq\mathcal{L}^{\prime}$.
Proof   We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$ as in the proof of Lemma 3.8. Because $Z_{B}$ is homotopy equivalent to $Z_{b_{0}}$, we obtain $\widehat{\kappa}_{L}^{\prime}:H_{1}(Z_{B},{\mathbb{Z}})\to U(1)$. There exists a unitary flat line bundle $(\widehat{\mathcal{L}},\nabla_{\widehat{\mathcal{L}}})$ on $Z_{B}$ which induces the homomorphism $\widehat{\kappa}_{L}^{\prime}$. Let $q_{2}:B\times X\to X$ denote the projection. We obtain the line bundle $(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2})$ on $Z_{B}$. We set
=^$\mathcal{L}=\widehat{\mathcal{L}}\otimes(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2}).$
Then, $\mathcal{L}$ is horizontal, and there exists an isomorphism $\mathcal{L}_{b_{0}}\simeq L$.
Let $\mathcal{L}^{\prime}$ be a horizontal holomorphic line bundle with an isomorphism $\mathcal{L}^{\prime}_{b_{0}}\simeq L$. We set $\mathcal{M}=\mathcal{L}^{\prime}\otimes\mathcal{L}^{-1}$. By Lemma 3.8, $\mathcal{M}_{b}\simeq\mathcal{O}_{Z_{b}}$ for any $b\in B$. Hence, $q_{0\ast}(\mathcal{M})$ is a holomorphic line bundle on $B$. Because $B$ is a multi-disc, there exists an isomorphism $q_{0\ast}(\mathcal{M})\simeq\mathcal{O}_{B}$. A nowhere vanishing section of $q_{0\ast}(\mathcal{M})$ induces an isomorphism $\mathcal{L}\simeq\mathcal{L}^{\prime}$.
Remark 3.10
We shall also explain another construction of a horizontal holomorphic line bundle in terms of locally defined symmetric pairings in §3.2.5.
3.2.3 Horizontal property and symmetric pairings
We continue to use the notation in §3.2.2. Let $\mathcal{R}_{\pi_{B}}$ denote the ramification divisor of $\pi_{B}$ on $Z_{B}$. (See §3.2.8.) Let $H$ be a closed complex analytic hypersurface of $Z_{B}$ such that $H\to B$ is proper and finite, and that $|\mathcal{R}_{\pi_{B}}|\subset H$. For any $b\in H$, we set $H_{b}:=H\cap Z_{b}$. We assume the following.
Condition 3.11
There exist open subsets $(Z_{B})_{0}$ and $(Z_{B})_{Q}$ $(Q\in H_{b_{0}})$ such that the following holds.
  • $Z_{B}=(Z_{B})_{0}\cup\bigcup_{Q\in H_{b_{0}}}(Z_{B})_{Q}$.
  • $(Z_{B})_{0}$, $(Z_{B})_{Q}$ and $(Z_{B})_{0}\cap(Z_{B})_{Q}$ are Stein.
  • $\overline{(Z_{B})_{Q}}\cap\overline{(Z_{B})_{Q^{\prime}}}=\emptyset$ if $Q\neq Q^{\prime}$.
  • The boundaries $\partial(Z_{B})_{0}$ and $\partial(Z_{B})_{Q}$ are submanifolds of $Z_{B}$, and the projections $\partial(Z_{B})_{0}\to B$ and $\partial(Z_{B})_{Q}\to B$ are proper and fiber bundles.
  • There exists a bi-holomorphic isomorphism $(Z_{B})_{0}\simeq B\times((Z_{B})_{0}\cap Z_{b_{0}})$ over $B$.
  • Each $(Z_{B})_{Q}\cap Z_{b_{0}}$ is bi-holomorphic to an open disc. Each $(Z_{B})_{0}\cap(Z_{B})_{Q}\cap Z_{b_{0}}$ is homeomorphic to an annulus $\{1/2<|z|<1\}$.
Remark 3.12
We shall explain a way to obtain such an open covering in §3.2.7.
We set $(Z_{b})_{0}:=(Z_{B})_{0}\cap Z_{b}$, $(Z_{b})_{Q}:=(Z_{B})_{Q}\cap Z_{b}$ and $(Z_{b})_{0,Q}:=(Z_{B})_{0}\cap(Z_{B})_{Q}\cap Z_{b}$. We also set $(Z_{B})_{0,Q}=(Z_{B})_{0}\cap(Z_{B})_{Q}$. We note the following lemma which follows from the conditions.
Lemma 3.13
For any $Q\in H_{b_{0}}$ and $b\in B$, $(Z_{b})_{Q}$ is homeomorphic to an open disc, and $(Z_{b})_{0,Q}$ is homeomorphic to an annulus.
Let $p_{0}:(Z_{B})_{0}\to(Z_{b_{0}})_{0}$ denote the morphism induced by the isomorphism $(Z_{B})_{0}\simeq B\times(Z_{b_{0}})_{0}$ and the projection $B\times(Z_{b_{0}})_{0}\to(Z_{b_{0}})_{0}$. We shall prove the following proposition in §3.2.43.2.6.
Proposition 3.14
Let $\mathcal{L}$ be a holomorphic line bundle on $Z_{B}$ such that $\deg\mathcal{L}_{b}=n(n-1)(g_{X}-1)$ for any $b\in B$. Let $h_{\mathcal{L}_{b_{0}}}$ denote the singular flat metric of $\mathcal{L}_{b_{0}}$ as in §3.1.2. Then, $\mathcal{L}$ is horizontal if and only if there exist an isomorphism $\varphi_{0}:p_{0}^{\ast}(\mathcal{L}_{|(Z_{b_{0}})_{0}})\simeq\mathcal{L}_{|(Z_{B})_{0}}$ and non-degenerate symmetric pairings $C_{Q}:\mathcal{L}_{|(Z_{B})_{Q}}\otimes\mathcal{L}_{|(Z_{B})_{Q}}\simeq\mathcal{O}_{Z_{B}}(\mathcal{R}_{\pi_{B}})_{|(Z_{B})_{Q}}$ $(Q\in H_{b_{0}})$ such that the following holds.
  • $C_{Q|(Z_{b_{0}})_{Q}}$ is compatible with $h_{\mathcal{L}_{b_{0}}|(Z_{b_{0}})_{Q}}$. Moreover, on $(Z_{B})_{0,Q}$, we have $p_{0}^{\ast}(C_{Q|(Z_{b_{0}})_{0,Q}})=C_{Q}$ under the isomorphism $\varphi_{0}$.
Remark 3.15
As we shall see in §3.2.4, for each $b\in B$, there exists a singular flat metric $h_{\mathcal{L}_{b}}$ of $\mathcal{L}_{b}$ such that $h_{\mathcal{L}_{b}|(Z_{b})_{Q}}$ is compatible with $C_{Q|(S_{b})_{Q}}$.
3.2.4 The “if” part of Proposition 3.14
Let us prove the “if” part. Let $b\in B$. Let $h_{0,b}$ be the flat Hermitian metric of $\mathcal{L}_{b|(Z_{b})_{0}}$ induced by $h_{\mathcal{L}_{b_{0}}|(Z_{b_{0}})_{0}}$ and the isomorphism $\varphi_{0}$. Let $h_{Q,b}$ be the singular flat Hermitian metric of $\mathcal{L}_{b|(Z_{b})_{Q}}$ compatible with $C_{Q|(Z_{b})_{Q}}$. By the third condition, we obtain a singular flat Hermitian metric $h_{b}$ of $\mathcal{L}_{b}$ from $h_{0,b}$ and $h_{Q,b}$ $(Q\in H_{b_{0}})$. It is a singular flat Hermitian metric $h_{\mathcal{L}_{b}}$ as in Lemma 3.2.
We remark the following lemma as a consequence of Proposition 3.4.
Lemma 3.16
For $b\in B$, let $\kappa_{b}:H_{1}((Z_{b})_{0},{\mathbb{Z}})\to U(1)$ be the representation induced by the Chern connection of $\pi_{b}^{\ast}(K_{X}^{(n-1)/2})$ and the singular metric $h_{\pi_{b}^{\ast}(K_{X}^{(n-1)/2})}$. Then, $\kappa_{b}$ are independent of $b\in B$.
Let $L_{b}$ be a holomorphic line bundle on $Z_{b}$ such that $\widehat{\kappa}_{L_{b}}$ is equal to $\widehat{\kappa}_{\mathcal{L}_{b_{0}}}$. (See §3.1.4 for the representation $\widehat{\kappa}_{L_{b}}$.) Let
$\iota_{b}:H_{1}\Bigl{(}Z_{b}\setminus\bigcup(Z_{b})_{Q},{\mathbb{Z}}\Bigr{)}\longrightarrow H_{1}\bigl{(}Z_{b},{\mathbb{Z}}\bigr{)}$
denote the homomorphism induced by the inclusion. By Lemma 3.16 and the construction of $h_{\mathcal{L}_{b}}$, we obtain $\kappa_{L_{b}}\circ\iota=\kappa_{\mathcal{L}_{b}}\circ\iota$. Hence, by Lemma 3.3, $L_{b}$ is isomorphic to $\mathcal{L}_{b}$, i.e., $\mathcal{L}$ is horizontal.
3.2.5 Construction of horizontal family via symmetric pairings
Let $L$ be a holomorphic line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$. We obtain the holomorphic line bundle $\mathcal{L}_{0}=p_{0}^{\ast}(L_{|(Z_{b_{0}})_{0}})$ on $(Z_{B})_{0}$. We set $\mathcal{L}_{Q}=\mathcal{O}_{(Z_{B})_{Q}}$. We fix an isomorphism $\mathcal{L}_{Q|(Z_{b_{0}})_{Q}}\simeq L_{|(Z_{b_{0}})_{Q}}$.
Lemma 3.17
For each $Q$, there exists an isomorphism $\varphi_{1,Q}:\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$ such that $\varphi_{1,Q|(Z_{b_{0}})_{0,Q}}$ equals the identity.
Proof   We use Oka-Grauert principle (see [13, §1.1.2, Theorem B’]). There exists a topological isomorphism $\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. Because $(Z_{B})_{0,Q}$ is Stein, there exists a holomorphic isomorphism $\varphi^{\prime}_{1,Q}:\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. The restriction $\varphi^{\prime}_{1,Q|(Z_{b_{0}})_{0,Q}}$ can be regarded as a holomorphic function $\alpha:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}^{\ast}$. Both $(Z_{b_{0}})_{0,Q}$ and ${\mathbb{C}}^{\ast}$ are homotopic to $S^{1}=\bigl{\{}a\in{\mathbb{C}}\,\big{|}\,|a|=1\bigr{\}}$. Because $(Z_{B})_{0,Q}$ is Stein, we can modify $\varphi^{\prime}_{1,Q}$ such that the induced map $H_{1}((Z_{b_{0}})_{0,Q},{\mathbb{Z}})\to H_{1}({\mathbb{C}}^{\ast},{\mathbb{Z}})$ is 00. Then, there exists $\beta:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}$ such that $\alpha=\exp(\beta)$. Because $(Z_{B})_{0,Q}$ is Stein, there exists $\widetilde{\beta}:(Z_{B})_{0,Q}\to{\mathbb{C}}$ whose restriction to $(Z_{b_{0}})_{0,Q}$ equals $\beta$. Then, $\varphi_{1,Q}=\exp(-\widetilde{\beta})\varphi^{\prime}_{1,Q}$ satisfies the desired condition.
Because $(Z_{B})_{Q}$ is Stein, there exists a non-degenerate symmetric bilinear product
$\widetilde{C}_{Q}:\mathcal{L}_{Q}\otimes\mathcal{L}_{Q}\simeq\mathcal{O}_{Z_{B}}(\mathcal{R}_{\pi_{B}})_{|(Z_{B})_{Q}}.$
Let $h_{L}$ be the singular flat Hermitian metric of $L$ as in §3.1.2. There exists a non-degenerate symmetric pairing
$C_{Q}:L_{|(Z_{b_{0}})_{Q}}\otimes L_{|(Z_{b_{0}})_{Q}}\simeq\mathcal{O}_{Z_{b_{0}}}(\mathcal{R}_{\pi_{b_{0}}})_{|(Z_{b_{0}})_{Q}}$
compatible with $h_{L}$. We obtain the nowhere vanishing holomorphic function $g_{Q}$ on $(Z_{B})_{0,Q}$ determined by
$p_{0}^{\ast}C_{Q}=g_{Q}\cdot\varphi_{1,Q}^{\ast}\widetilde{C}_{Q}.$
Lemma 3.18
There exists a holomorphic function $g_{Q}^{1/2}$ on $(Z_{B})_{0,Q}$ such that $(g_{Q}^{1/2})^{2}=g_{Q}$.
Proof   The restriction of $g_{Q}$ to $(Z_{b_{0}})_{0,Q}$ has a square root. Then, it is easy to see that $g_{Q}$ also has a square root.
We obtain the following isomorphisms:
$\varphi_{2,Q}=g_{Q}^{-1/2}\varphi_{1,Q}:\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}.$
We obtain a holomorphic line bundle $\Psi(L)$ on $Z_{B}$ by gluing $\mathcal{L}_{0}$ and $\mathcal{L}_{Q}$ $(Q\in H_{b_{0}})$ via the isomorphisms $\varphi_{2,Q}$. Because we have already proved “the if part” of Proposition 3.14 in §3.2.4, we obtain that $\Psi(L)$ is horizontal. Because $(g^{1/2}_{Q})_{|(Z_{b_{0}})_{0,Q}}$ extends to a nowhere vanishing holomorphic function on $(Z_{b_{0}})_{Q}$, we can check that $\Psi(L)_{b_{0}}$ is isomorphic to $L$.
3.2.6 The “only if” part of Proposition 3.14
From $\mathcal{L}_{b_{0}}$, we obtain the horizontal holomorphic line bundle $\Psi(\mathcal{L}_{b_{0}})$ such that $\Psi(\mathcal{L}_{b_{0}})_{b_{0}}\simeq\mathcal{L}_{b_{0}}$ by applying the construction in §3.2.5. By the uniqueness in Lemma 3.9, $\Psi(\mathcal{L}_{b_{0}})\simeq\mathcal{L}$. By the construction of $\Psi(\mathcal{L}_{b_{0}})$, there exist an isomorphism $\varphi_{0}:p_{0}^{\ast}\bigl{(}\Psi(\mathcal{L}_{b_{0}})_{|(Z_{b_{0}})_{0}}\bigr{)}\simeq\Psi(\mathcal{L}_{b_{0}})_{|(Z_{B})_{0}}$ and non-degenerate symmetric pairings $C_{Q}:\Psi(\mathcal{L}_{b_{0}})_{|(Z_{B})_{Q}}\otimes\Psi(\mathcal{L}_{b_{0}})_{|(Z_{B})_{Q}}\simeq\mathcal{O}_{Z_{B}}(\mathcal{R}_{\pi_{B}})_{|(Z_{B})_{Q}}$ $(Q\in H_{b_{0}})$ satisfying the condition in Proposition 3.14. Thus, the “only if” part of Proposition 3.14 is proved.
3.2.7 Complement
We set $D_{b}=\pi_{b}(|\widetilde{D}_{b}|)$. (See §3.2.2 for $\widetilde{D}_{b}$.) For each $P\in D_{b_{0}}$, let $X_{P}$ be a simply connected open neighbourhood of $P$ in $X$ such that $X_{P}\cap D_{b_{0}}=\{P\}$. Let $X_{P}^{\prime}$ be a simply connected relatively compact open neighbourhood of $P$ in $X_{P}$.
Condition 3.19
$D_{b}\subset\bigcup X_{P}^{\prime}$ for any $b\in B$.
We set $X_{1}=X\setminus\bigcup_{P\in D_{b_{0}}}\overline{X^{\prime}_{P}}$. We set $H=\pi_{B}^{-1}\bigl{(}\pi_{B}(|\mathcal{R}_{\pi_{B}}|)\bigr{)}\subset Z_{B}$. We set $(Z_{B})_{0}=\pi_{B}^{-1}(B\times X_{1})$. For each $Q\in\pi_{b_{0}}^{-1}(D_{b_{0}})$, let $(Z_{B})_{Q}$ denote the connected component of $\pi_{B}^{-1}(B\times X_{\pi_{b_{0}}(Q)})$ which contains $Q$.
Lemma 3.20
Condition 3.11 is satisfied for $(Z_{B})_{0}$ and $(Z_{B})_{Q}$ $(Q\in H_{b_{0}})$.
Proof   We can check that $(Z_{B})_{0}$, $(Z_{B})_{Q}$ and $(Z_{B})_{0,Q}$ are Stein by using [12, §V.1 Theorem 1]. The natural morphisms $a:(Z_{B})_{0}\to B\times X_{1}$ and $b:B\times(Z_{b_{0}})_{0}\to B\times X_{1}$ are covering maps over $B$. Because they are the same maps over $\{b_{0}\}\times X_{1}$, there exists the unique homeomorphism $c:(Z_{B})_{0}\longrightarrow B\times(Z_{b_{0}})_{0}$ such that $a=b\circ c$. It is easy to see that $c$ and $c^{-1}$ are holomorphic. The other conditions can also be checked easily.
3.2.8 Appendix: Ramification divisor
Let $a_{i}:\mathcal{X}_{i}\to B$ be smooth morphisms of complex manifolds such that $\dim\mathcal{X}_{1}=\dim\mathcal{X}_{2}=\dim B+1$. Let $\varphi:\mathcal{X}_{1}\to\mathcal{X}_{2}$ be a proper morphism of complex manifolds over $B$, i.e., $a_{2}\circ\varphi=a_{1}$. Let $K_{\mathcal{X}_{i}}$ $(i=1,2)$ denote the canonical line bundle of $\mathcal{X}_{i}$. There exists the effective divisor $\mathcal{R}_{\varphi}$ on $\mathcal{X}_{1}$ such that the natural morphism $\varphi^{\ast}K_{\mathcal{X}_{2}}\to K_{\mathcal{X}_{1}}$ induces an isomorphism
$K_{\mathcal{X}_{1}}\simeq\varphi^{\ast}(K_{\mathcal{X}_{2}})\otimes\mathcal{O}_{\mathcal{X}_{1}}(\mathcal{R}_{\varphi}).$
The divisor $\mathcal{R}_{\varphi}$ is called the ramification divisor of $\varphi$.
In general, for a complex manifold $Y$, let $\Omega^{1}_{Y}$ denote the sheaf of holomorphic $1$-forms on $Y$. Let $\Omega^{1}_{\mathcal{X}_{i}/B}$ denote the cokernel of $a_{i}^{\ast}\Omega^{1}_{B}\to\Omega^{1}_{\mathcal{X}_{i}}$, which is a holomorphic line bundle on $\mathcal{X}_{i}$. The natural morphism $\varphi^{\ast}\Omega^{1}_{\mathcal{X}_{2}/B}\longrightarrow\Omega^{1}_{\mathcal{X}_{1}/B}$ induces an isomorphism
$\Omega^{1}_{\mathcal{X}_{1}/B}\simeq\varphi^{\ast}\Omega^{1}_{\mathcal{X}_{2}/B}\otimes\mathcal{O}_{\mathcal{X}_{1}}(\mathcal{R}_{\varphi}).$
Let $f:B^{\prime}\to B$ be any morphism of complex manifolds. We set $\mathcal{X}_{i}^{\prime}=\mathcal{X}_{i}\times_{B}B^{\prime}$. We obtain the induced morphisms $a_{i}^{\prime}:\mathcal{X}_{i}^{\prime}\to B^{\prime}$ and $\varphi^{\prime}:\mathcal{X}_{1}^{\prime}\to\mathcal{X}_{2}^{\prime}$. We also obtain the induced morphism $f_{1}:\mathcal{X}^{\prime}_{1}\to\mathcal{X}_{1}$. The ramification divisor is compatible with the base change, i.e., $f_{1}^{\ast}(\mathcal{R}_{\varphi})=\mathcal{R}_{\varphi^{\prime}}$.
3.3 Horizontal deformations of Higgs bundles with smooth spectral curves
3.3.1 Horizontal family of Higgs bundles with smooth spectral curves
Let $B$ be a simply connected Hausdorff space. Let $(E_{b},\theta_{b})$ $(b\in B)$ be a family of Higgs bundles of rank $n$ of degree 00 on $X$ such that the following holds.
  • Each $\Sigma_{\theta_{b}}$ is smooth. The induced map $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ is continuous.
According to Hitchin and Beauville-Narasimhan-Ramanan, for each $b\in B$, there exists a line bundle $L_{b}$ on $\Sigma_{\theta_{b}}=Z_{b}$ of degree $n(n-1)(g_{X}-1)$ with an isomorphism $\pi_{\varphi(b)\ast}(L_{b})\simeq(E_{b},\theta_{b})$. Such an $L_{b}$ is unique up to an isomorphism.
Definition 3.21
The family $(E_{b},\theta_{b})$ $(b\in B)$ is called horizontal if the family $L_{b}$ $(b\in B)$ is horizontal.
The following lemma is well known and an obvious translation of Lemma 3.8.
Lemma 3.22
Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ whose spectral curve is smooth. Then, there exists a horizontal family $(E_{b},\theta_{b})$ $(b\in B)$ such that $(E_{b_{0}},\theta_{b_{0}})\simeq(E,\theta)$. For another such family $(E^{\prime}_{b},\theta^{\prime}_{b})$ $(b\in B)$, there exist isomorphisms $(E_{b},\theta_{b})\simeq(E^{\prime}_{b},\theta^{\prime}_{b})$.
3.3.2 Complex analytic property of horizontal families
Let $B$ and $\varphi$ be as in §3.2.2.
Definition 3.23
Let $(\mathcal{E},\theta)$ be a relative Higgs bundle on $B\times X$ such that $\Sigma_{\mathcal{E},\theta}=Z_{B}$. We set $(\mathcal{E}_{b},\theta_{b}):=(\mathcal{E},\theta)_{|\{b\}\times X}$. Assume that $\deg(\mathcal{E}_{b})=0$. We say that $(\mathcal{E},\theta)$ is horizontal if the family $(\mathcal{E}_{b},\theta_{b})$ $(b\in B)$ is horizontal.
Lemma 3.24
Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ such that $\Sigma_{\theta}=Z_{b_{0}}$. There exists a horizontal relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $B\times X$ such that $(\mathcal{E}_{b_{0}},\theta_{\mathcal{E},b_{0}})=(E,\theta)$. If $(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$ be another such relative Higgs bundle, then there exists an isomorphism $(\mathcal{E},\theta_{\mathcal{E}})\simeq(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$.
Proof   Let $L$ be a line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$ corresponding to $(E,\theta)$. By Lemma 3.8, there exists a horizontal line bundle $\mathcal{L}$ on $Z_{B}$ such that $\mathcal{L}_{b_{0}}\simeq L$. According to [11, Chapter 2, §4.2], $\mathcal{E}=\pi_{B\ast}\mathcal{L}$ is a locally free $\mathcal{O}_{B\times X}$-module. It is naturally equipped with the relative Higgs field $\theta_{\mathcal{E}}$, and $(\mathcal{E}_{b_{0}},\theta_{\mathcal{E},b_{0}})\simeq(E,\theta)$. The second claim also follows from Lemma 3.8.
3.3.3 Horizontal property and symmetric pairings
Let $X_{P}$, $X_{P^{\prime}}$ and $X_{1}$ be as in §3.2.7. Let $(\mathcal{E},\theta_{\mathcal{E}})$ be a relative Higgs bundle on $B\times X$ such that (i) $\Sigma_{\mathcal{E},\theta_{\mathcal{E}}}=Z_{B}$, (ii) $\deg(\mathcal{E}_{b},\theta_{\mathcal{E},b})=0$ for any $b\in B$.
Let $q_{2}:B\times X_{1}\to X_{1}$ denote the projection. Let $\mathcal{L}_{1}$ be the line bundle on $\pi_{B}^{-1}(B\times X_{1})$ which induces $(\mathcal{E},\theta_{\mathcal{E}})_{|B\times X_{1}}$. There exists the holomorphic isomorphism $\rho_{1}:B\times\pi_{b_{0}}^{-1}(X_{1})\simeq\pi_{B}^{-1}(B\times X_{1})$ over $B\times X_{1}$.
We obtain the following proposition as a consequence of Proposition 3.14 and Lemma 3.20.
Proposition 3.25
$(\mathcal{E},\theta_{\mathcal{E}})$ is horizontal if and only if there exist a holomorphic isomorphism $\varphi:q_{2}^{\ast}(\mathcal{E}_{b_{0}|X_{1}})\simeq\mathcal{E}_{|B\times X_{1}}$ and non-degenerate symmetric pairings $C_{P}$ of $(\mathcal{E},\theta_{\mathcal{E}})_{|B\times X_{P}}$ $(P\in D_{b_{0}})$ such that the following holds.
  • $\varphi$ is induced by a holomorphic isomorphism $\rho_{1}^{\ast}(\mathcal{L}_{1|\pi_{b_{0}}^{-1}(X_{1})})\simeq\mathcal{L}_{1|\pi_{B}^{-1}(B\times X_{1})}$.
  • Each $C_{P|\{b_{0}\}\times X_{P}}$ is compatible with the limiting configuration of $(\mathcal{E}_{b_{0}},\theta_{\mathcal{E},b_{0}})$. Moreover, on $B\times(X_{1}\cap X_{P})$, we have $q_{2}^{\ast}C_{P|\{b_{0}\}\times(X_{1}\cap X_{P})}=C_{P|B\times(X_{1}\cap X_{P})}$ under the isomorphism $\varphi$.
3.4 Horizontal infinitesimal deformations
Suppose that $X$ is compact. Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ which is generically regular semisimple with smooth spectral curve.
Definition 3.26
  • An infinitesimal deformation $(E^{[1]},\theta^{[1]})$ of $(E,\theta)$ is called horizontal if there exists a relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $S\times X$ such that (i) $(\mathcal{E},\theta_{\mathcal{E}})$ is horizontal, (ii) there exists $s_{0}\in S$ and $v\in T_{s_{0}}S$ such that $\iota_{v}^{\ast}(\mathcal{E},\theta_{\mathcal{E}})\simeq(E^{[1]},\theta^{[1]})$. (See §2.3.2 for the notation.)
  • An element of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is called horizontal if it corresponds to a horizontal infinitesimal deformation. Let $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\subset H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ denote the subset of horizontal elements.
Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 00 and rank $n$ on $X$. There exists the Hitchin fibration $\Phi_{H}:\mathcal{M}_{H}\to{\boldsymbol{A}}_{X,n}$. We obtain the point $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ corresponding to $(E,\theta)$. Let $\mathcal{U}$ be any simply connected open neighbourhood of $\Phi_{H}(E,\theta)$. According to Lemma 3.24, there exists a morphism of complex manifolds $\Psi:\mathcal{U}\to\mathcal{M}_{H}$ such that (i) $\Phi_{H}\circ\Psi=\mathop{\rm id}\nolimits_{\mathcal{U}}$, (ii) the family of Higgs bundles $\Psi(x)$ $(x\in\mathcal{U})$ is horizontal, (iii) $\Psi(\Phi_{H}(E,\theta))=(E,\theta)$.
The tangent space $T_{(E,\theta)}\mathcal{M}_{H}$ is identified with $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. We obtain the following lemma by definition.
Lemma 3.27
$T_{(E,\theta)}\Psi(\mathcal{U})\subset T_{(E,\theta)}\mathcal{M}_{H}$ is identified with $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$. In particular, $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ is a ${\mathbb{C}}$-vector subspace of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ such that $2\dim H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}=\dim H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$.
3.4.1 Čech description of horizontal infinitesimal deformations
For each $P\in D(\theta)$, let $(X_{P},z_{P})$ denote a holomorphic coordinate neighbourhood around $P$ as in §2.1.1. We set $(E_{P},\theta_{P}):=(E,\theta)_{|X_{P}}$. Let $C_{P}$ be a non-degenerate symmetric pairing of $(E_{P},\theta_{P})$ compatible with $h_{\infty}$ on $X_{P}\setminus\{P\}$. Let $X_{P}^{\prime}$ be a relatively compact open neighbourhood of $P$ in $X_{P}$. We set $X_{0}=X\setminus\bigcup_{P\in D(\theta)}\overline{X_{P}^{\prime}}$. We obtain an open covering $X=X_{0}\cup\bigcup_{P\in D(\theta)}X_{P}$. We assume $X_{P}\cap X_{P^{\prime}}=\emptyset$ $(P\neq P^{\prime})$.
We may regard $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X_{0}}$ as a subsheaf of $(\mathop{\rm End}\nolimits(E)\otimes K_{X})_{|X_{0}}$ by the splitting (18). We may also regard $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta)_{|X_{0}\cap X_{P}}$ as a subsheaf of $\bigl{(}\mathop{\rm Hom}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})\otimes K_{X}\bigr{)}_{|X_{0}\cap X_{P}}$.
Then, we define a subcomplex $\mathfrak{A}^{\bullet}$ of $\mathop{\rm Tot}\nolimits\check{C}^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ as follows:
$\mathfrak{A}^{0}:=\bigoplus_{P\in D(\theta)}H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P})),$
$\mathfrak{A}^{1}:=H^{0}(X_{0},\mathop{\rm Cok}\nolimits\mathop{\rm ad}\nolimits\theta)\oplus\bigoplus_{P\in D(\theta)}H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})\otimes K_{X})\oplus\bigoplus_{P\in D(\theta)}H^{0}(X_{0}\cap X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P})),$
$\mathfrak{A}^{2}:=\bigoplus_{P\in D(\theta)}H^{0}(X_{0}\cap X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})\otimes K_{X}).$
We obtain the morphism
$\Upsilon_{0}:H^{1}\bigl{(}\mathfrak{A}^{\bullet}\bigr{)}\longrightarrow H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta)).$
By the isomorphism $\mathop{\rm Cok}\nolimits(\theta)\simeq\pi_{\ast}K_{\Sigma_{\theta}}$ in Proposition 2.4, we obtain the following natural morphism:
$\Upsilon_{1}:H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))\to H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}}).$ (32)
Theorem 3.28
The composite $\Upsilon_{1}\circ\Upsilon_{0}$ is an isomorphism. The image of $\Upsilon_{0}$ equals the space of isomorphism classes of horizontal infinitesimal deformations $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$.
Proof   Let us prove that $\Upsilon_{1}\circ\Upsilon_{0}$ is injective. Let $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$ for $a\in H^{1}(\mathfrak{A}^{\bullet})$. Let $(\varphi_{0},\varphi_{P},\psi_{P})\in\mathfrak{A}^{1}$ be a representative of $a$. Because $\Upsilon_{1}\circ\Upsilon_{0}(a)=0$, we obtain that $\varphi_{0}=0$ and $[\varphi_{P}]=0$ in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits\theta_{P})$ $(P\in D(\theta))$. There exist $\nu_{P}\in H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P}))$ such that $\mathop{\rm ad}\nolimits(\theta_{P})\nu_{P}=\varphi_{P}$. Because we have $\varphi_{P|X_{P}\cap X_{0}}=\mathop{\rm ad}\nolimits\theta_{P}(\psi_{P})$, we obtain $\nu_{P|X_{P}\cap X_{0}}=\psi_{P}$. Hence, we obtain $a=0$ in $H^{1}(\mathfrak{A}^{\bullet})$.
We obtain $\dim H^{1}(\mathfrak{A}^{\bullet})\leq\dim H^{0}(\Sigma,K_{\Sigma})=\frac{1}{2}\dim H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. By Proposition 3.25, any horizontal infinitesimal deformation is contained in the image of $\Upsilon_{0}$. The dimension of the space of the isomorphism classes of horizontal deformations equals $\dim{\boldsymbol{A}}=\frac{1}{2}\dim H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. Hence, we obtain that the image of $\Upsilon_{0}$ equals the space of the isomorphism classes of horizontal infinitesimal deformations. Because $\dim H^{1}(\mathfrak{A}^{\bullet})=\dim H^{0}(\Sigma,K_{\Sigma})$, $\Upsilon_{1}\circ\Upsilon_{0}$ is an isomorphism.
Corollary 3.29
$\Upsilon_{0}$ is injective, and $\Upsilon_{1}$ is surjective.
We define the isomorphism $\iota^{{\tt h}}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ by
$\iota^{{\tt h}}(\nu)=\Upsilon_{0}\circ\bigl{(}\Upsilon_{1}\circ\Upsilon_{0}\bigr{)}^{-1}(\nu).$
We also regard $\iota^{{\tt h}}$ as a morphism $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ in a natural way.
Let us give an explicit description of $(\Upsilon_{1}\circ\Upsilon_{0})^{-1}(\nu)\in H^{1}(\mathfrak{A}^{\bullet})$ for $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. As in §2.2.5, we have
$F_{\nu}\in H^{0}(X,\mathop{\rm End}\nolimits(E)\otimes K_{\Sigma_{\theta}}(D)).$
By Lemma 2.10, there exists a section $g_{P}\in H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E_{P};C_{P})(D))$ on $X_{P}$ such that
$\varphi_{P}:=F_{\nu|X_{P}}-\mathop{\rm ad}\nolimits(\theta)(g_{P})\in H^{0}(X_{P},\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E_{P};C_{P})\otimes K_{X_{P}}).$
We set $\psi_{P}:=g_{P|X_{P}\cap X_{0}}$ and $\varphi_{0}:=F_{\nu|X_{0}}$. Then, we obtain a $1$-cocycle $a=(\varphi_{0},\varphi_{P},\psi_{P})$ in $\mathfrak{A}^{1}$. The following lemma is clear by the construction.
Lemma 3.30
The cohomology class $[a]\in H^{1}(\mathfrak{A}^{\bullet})$ is $(\Upsilon_{1}\circ\Upsilon_{0})^{-1}(\nu)$. Its image in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is $\iota^{{\tt h}}(\nu)$.
Remark 3.31
We shall study the harmonic representative of $\iota^{{\tt h}}(\nu)$ later.
3.4.2 $C^{\infty}$-description of horizontal infinitesimal deformations
Let $\nu$ be a holomorphic $1$-form on $\Sigma_{\theta}$. We obtain the following proposition from Theorem 3.28.
Corollary 3.32
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ be a $1$-cocycle such that $\rho(\tau)=\nu$. (See Lemma 2.11 for the map $\rho$.) We also assume the following conditions.
  • $\tau_{|X\setminus\bigcup X_{P}^{\prime}}=F_{\nu|X\setminus\bigcup X_{P}^{\prime}}$.
  • $\tau_{|X_{P}}\in A^{1}(\mathop{\rm Def}\nolimits(E_{P},\theta_{P};C_{P}))$ for each $P\in D(\theta)$.
Then, $[\tau]=\iota^{{\tt h}}(\nu)$ holds in $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$. In particular, $[\tau]\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$. Conversely, any element of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ is described in this way.
3.4.3 Lagrangian property of the horizontal subspaces
For any Higgs bundle $(E,\theta)$ of rank $n$ of degree 00 on $X$, let $\omega_{\mathcal{M}_{H},(E,\theta)}$ denote the alternative bilinear form on $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ defined as follows [14, (6.5)]:
$\omega_{\mathcal{M}_{H},(E,\theta)}(\tau_{1},\tau_{2})=-\int_{X}\mathop{\rm Tr}\nolimits(\tau_{1}\wedge\tau_{2}).$ (33)
It induces the alternative bilinear form on $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$, which is also denoted by $\omega_{\mathcal{M}_{H},(E,\theta)}$. According to Hitchin, the family of the alternative bilinear forms $\bigcup_{(E,\theta)\in\mathcal{M}_{H}}\omega_{\mathcal{M}_{H},(E,\theta)}$ induces a holomorphic symplectic form $\omega_{\mathcal{M}_{H}}$ of $\mathcal{M}_{H}$.
Corollary 3.33
Let $(E,\theta)\in\mathcal{M}_{H}^{\prime}$. Then, the horizontal subspace $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is Lagrangian with respect to $\omega_{\mathcal{M}_{H},(E,\theta)}$.
Proof   For any $\iota^{{\tt h}}(\nu_{i})\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ $(i=1,2)$, there exists a representative $\tau_{i}\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfying the conditions in Corollary 3.32. On $X\setminus\bigcup X_{P}^{\prime}$, we have $\tau_{i}=\tau_{i}^{1,0}$, and hence $\tau_{1}\wedge\tau_{2}=0$. On $X_{P}$, $\tau_{i}^{1,0}$ (resp. $\tau_{i}^{0,1}$) are symmetric (anti-symmetric) with respect to $C_{P}$. We recall that for a symmetric $(n\times n)$-matrix $A$ and an anti-symmetric $(n\times n)$-matrix $B$, we obtain $\mathop{\rm tr}\nolimits(AB)=0$ because
$\mathop{\rm tr}\nolimits(AB)=\mathop{\rm tr}\nolimits(BA)=\mathop{\rm tr}\nolimits\bigl{(}{}^{t}\!(BA)\bigr{)}=\mathop{\rm tr}\nolimits\bigl{(}{}^{t}\!A{}^{t}\!B\bigr{)}=-\mathop{\rm tr}\nolimits(AB).$
Hence, we obtain $\mathop{\rm tr}\nolimits(\tau_{1}\wedge\tau_{2})=\mathop{\rm tr}\nolimits\bigl{(}\tau_{1}^{1,0}\wedge\tau_{2}^{0,1}\bigr{)}+\mathop{\rm tr}\nolimits\bigl{(}\tau_{1}^{0,1}\wedge\tau_{2}^{1,0}\bigr{)}=0$ on $X_{P}$. Hence, we obtain $\omega_{\mathcal{M}_{H},(E,\theta)}(\iota^{{\tt h}}(\nu_{1}),\iota^{{\tt h}}(\nu_{2}))=\omega_{\mathcal{M}_{H},(E,\theta)}(\tau_{1},\tau_{2})=0$.
4 Comparison of the Hitchin metric and the semi-flat metric
4.1 The semi-flat metric
4.1.1 Vertical infinitesimal deformations
We recall the following. (See §2.3.2 for relative Higgs bundles and the induced infinitesimal deformation.)
Definition 4.1
  • A family of Higgs bundles $(E_{s},\theta_{s})$ $(s\in S)$ is called vertical if the spectral curves $\Sigma_{\theta_{s}}\subset T^{\ast}X$ are independent of $s$.
  • A relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $S\times X$ is called vertical if the associated family $(\mathcal{E},\theta_{\mathcal{E}})_{|\{s\}\times X}$ $(s\in S)$ is vertical.
  • An infinitesimal deformation of $(E,\theta)$ is called vertical if it is induced by a vertical relative Higgs bundle. Let $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\subset H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ denote the subset of vertical infinitesimal deformations.
Suppose that $(E,\theta)$ is generically regular semisimple and that the spectral curve $\Sigma_{\theta}$ is smooth. The natural morphism $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}=\mathcal{H}^{0}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta)\bigr{)}\longrightarrow\mathop{\rm Def}\nolimits(E,\theta)$ induces a morphism
$\iota^{{\tt v}}:H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\longrightarrow H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta)).$ (34)
Lemma 4.2
The morphism (34) is injective. The image equals $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$.
Proof   Because $\pi_{\ast}\mathcal{O}_{\Sigma_{\theta}}=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,\theta))$, we obtain the first claim. Let $U\subset{\mathbb{C}}$ be a neighbourhood of 00. For any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$, there exists a holomorphic line bundle $\mathcal{L}$ on $U\times\Sigma_{\theta}$ which induces $a$ as the infinitesimal part. Then, the second claim is clear.
Any $a\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ is represented by a $(0,1)$-form $\nu$ such that the support of $\nu$ is compact in $\Sigma_{\theta}\setminus\Sigma_{\theta|D(\theta)}$. We obtain the $C^{\infty}$-section $F^{\prime}_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X^{\circ}$ as in §2.2.7 whose support is compact. It naturally extends to a $C^{\infty}$-section $F_{\nu}$ of $\mathop{\rm End}\nolimits(E)\otimes\Omega^{0,1}$ on $X$. Because $[\theta,F_{\nu}]=0$, it is a $1$-cocycle in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$. The image $\iota^{{\tt v}}(a)\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ is represented by $F_{\nu}$.
Remark 4.3
We shall study the harmonic representative of $\iota^{{\tt v}}(a)$ later.
As we recalled in §3.4.3, there exists the complex symplectic structure $\omega_{\mathcal{M}_{H}}$ of $\mathcal{M}_{H}$. The following is well known.
Lemma 4.4
$H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ is a Lagrangian subspace with respect to $\omega_{\mathcal{M}_{H},(E,\theta)}$.
Proof   For $\iota^{{\tt v}}(\nu_{i})\in H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$, let $F_{\nu_{i}}$ be representatives as above. Then, $F_{\nu_{1}}\wedge F_{\nu_{2}}=0$. Hence, $\omega_{\mathcal{M}_{H},(E,\theta)}(\iota^{{\tt v}}(\nu_{1}),\iota^{{\tt v}}(\nu_{2}))=0$.
4.1.2 The induced perfect pairings
There exists the following decomposition.
$H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))=H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\oplus H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}.$
Because $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ is Lagrangian, we obtain the following induced perfect pairings.
$H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\otimes H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\longrightarrow{\mathbb{C}}$ (35)
Lemma 4.5
Under the isomorphisms $\iota^{{\tt v}}:H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\simeq H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ and $\Upsilon_{1}:H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\simeq H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, the perfect pairing (35) is identified with the perfect pairing (26).
Proof   There exists a $(0,1)$-form $\nu$ on $\Sigma_{\theta}$ such that (i) it represents $a$, (ii) the support of $\nu$ is compact in $\Sigma_{\theta}\setminus\Sigma_{\theta|D(\theta)}$. Then, $\iota^{{\tt v}}(a)$ is represented by $F_{\nu}$. For $b\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, there exists a closed $1$-form $\tau$ in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ such that $\rho(\tau)=b$. (See Lemma 2.11 for $\rho$.) We have
$\omega_{\mathcal{M}_{H},(E,\theta)}(\iota^{{\tt v}}(a),[\tau])=-\int_{X}\mathop{\rm Tr}\nolimits(F_{\nu}\cdot\tau)=-\int_{\Sigma_{\theta}}a\wedge b.$
Thus, we obtain Lemma 4.5.
4.1.3 Hermitian metrics on $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$
Let $g_{1}$ and $g_{2}$ be the Hermitian metrics of $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ as in §2.7. There exists the decomposition
$H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))=H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\oplus H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\simeq H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}}).$
Let $g_{1}\oplus g_{2}$ denote the Hermitian metric of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$ induced by $g_{i}$ $(i=1,2)$.
4.1.4 The semi-flat metric
Note that $\Phi_{H}:\mathcal{M}_{H}^{\prime}\to{\boldsymbol{A}}_{X,n}^{\prime}$ is an algebraic integrable system. For each $s\in{\boldsymbol{A}}_{X,n}^{\prime}$, we obtain the spectral curve $\Sigma_{s}\subset T^{\ast}X$, and $\Phi_{H}^{-1}(s)$ is identified with the moduli space $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ of holomorphic line bundles on $\Sigma_{s}$ of degree $d=n(n-1)(g_{X}-1)$. By choosing $K_{X}^{1/2}$, we obtain the Hitchin section $\Psi$ of $\Phi_{H}$. By using the Hitchin section, we may regard $\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s})$ as the quotient of $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ by a lattice. The space $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$ is equipped with the real symplectic form $\omega_{s}$ which induces a positive symmetric bilinear form on $H^{1}(\Sigma_{s},\mathcal{O}_{\Sigma_{s}})$. The cohomology class $[\omega_{s}]\in H^{2}(\mathop{\rm Pic}\nolimits_{d}(\Sigma_{s}),{\mathbb{R}})$ is a positive polarization in the sense of [8, §3]. According to [8, Theorem 3.8], we obtain a hyperkähler metric $g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}$, which is called the semi-flat metric. Let $g_{\mathop{\rm sf}\nolimits}$ denote the associated Kähler metric with respect to the natural complex structure of $\mathcal{M}^{\prime}_{H}$ as an open subset of the moduli space of Higgs bundles.
Proposition 4.6
Under the identification $T_{(E,\theta)}\mathcal{M}^{\prime}_{H}\simeq H^{1}(X,\mathop{\rm End}\nolimits(E,\theta))$, the metric $(g_{\mathop{\rm sf}\nolimits})_{(E,\theta)}$ equals $g_{1}\oplus g_{2}$.
Proof   Let $\omega_{1}$ and $\omega_{2}$ be the real symplectic forms on $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ and $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ as in §2.7. Let us recall the construction of the semi-flat metric at $T_{(E,\theta)}\mathcal{M}$ by following [8, §2 and §3].
Let $\omega_{\mathcal{M}_{H}}$ denote the complex symplectic structure of $\mathcal{M}_{H}$ in [14, (6.5)]. Under the isomorphism $T_{(E,\theta)}\mathcal{M}_{H}\simeq H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))$, $(\omega_{\mathcal{M}_{H}})_{(E,\theta)}$ equals the symplectic form $\omega_{\mathcal{M}_{H},(E,\theta)}$ in §4.1.2. The tangent space $T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}_{X,n}$ is identified with the dual space of the space of the isomorphism classes of vertical infinitesimal deformations:
$T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}_{X,n}\simeq H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\simeq\bigl{(}H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\bigr{)}^{\lor}=H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})^{\lor}.$ (36)
We obtain the positive symplectic form $\omega_{\Phi_{H}(E,\theta)}=-\omega_{1}^{\lor}$ on $T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}_{X,n}$. In this way, we obtain the Kähler form $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}=\{\omega_{s}\}_{s\in{\boldsymbol{A}}^{\prime}_{X,n}}$ on ${\boldsymbol{A}}^{\prime}_{X,n}$. The decomposition $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))=H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}\oplus H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}$ induces the flat connection on the tangent bundle of $\mathcal{M}_{H}^{\prime}$. They give the special Kähler structure on ${\boldsymbol{A}}^{\prime}_{X,n}$. (See [8, §3].) We obtain the special Kähler metric $g_{3}$ of ${\boldsymbol{A}}^{\prime}_{X,n}$ induced by $\omega_{{\boldsymbol{A}}^{\prime}_{X,n}}$. Let $g_{3,{\mathbb{R}}}$ denote the real part of $g_{3}$.
Let $g_{3,{\mathbb{R}}|(E,\theta)}$ denote the induced positive definite symmetric bilinear form of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\simeq T_{\Phi_{H}(E,\theta)}{\boldsymbol{A}}^{\prime}_{X,n}$. We obtain the positive definite symmetric bilinear form $g_{3,{\mathbb{R}}|(E,\theta)}^{\lor}$ of $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt v}}$ by the duality. By definition of $g_{\mathop{\rm sf}\nolimits}$, we have
$(g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}})_{(E,\theta)}=g_{3,{\mathbb{R}}|(E,\theta)}^{\lor}\oplus g_{3,{\mathbb{R}}|(E,\theta)}.$
Under the identification $H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))^{{\tt h}}\simeq H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, $g_{3,{\mathbb{R}}|(E,\theta)}$ is the positive definite symmetric bilinear form associated with $\omega_{2}=-\omega_{1}^{\lor}$ by Lemma 2.41, we obtain $g_{3,{\mathbb{R}}|(E,\theta)}=g_{2,{\mathbb{R}}}$. Because $g_{1,{\mathbb{R}}}=g_{2,{\mathbb{R}}}^{\lor}$ by Lemma 2.41, we obtain $g_{3,{\mathbb{R}}|(E,\theta)}^{\lor}=g_{1,{\mathbb{R}}}$.
4.2 The Hitchin metric and the comparison with the semi-flat metric
4.2.1 The Hitchin metric
Let $(E,\theta)\in\mathcal{M}_{H}$. Let $h$ be a harmonic metric of $(E,\theta)$. Let $\mathop{\rm Harm}\nolimits^{1}(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$ denote the space of harmonic $1$-forms of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$. There exist the following isomorphisms:
$T_{(E,\theta)}\mathcal{M}_{H}\simeq H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))\simeq\mathop{\rm Harm}\nolimits^{1}(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h).$
For $\tau\in\mathop{\rm Harm}\nolimits^{1}(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$, we set
$g_{H,{\mathbb{R}}|(E,\theta)}(\tau,\tau)=2\sqrt{-1}\int_{X}\bigl{(}h\bigl{(}\tau^{1,0},\tau^{1,0}\bigr{)}-h\bigl{(}\tau^{0,1},\tau^{0,1}\bigr{)}\bigr{)}=2\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl{(}\tau^{1,0}\wedge(\tau^{1,0})^{\dagger}_{h}-\tau^{0,1}\wedge(\tau^{0,1})^{\dagger}_{h}\bigr{)}.$
We obtain the positive definite symmetric product $g_{H,{\mathbb{R}}|(E,\theta)}$ of the ${\mathbb{R}}$-vector space $T_{(E,\theta)}\mathcal{M}_{H}$, and the Hermitian metric $g_{H|(E,\theta)}$ of $T_{(E,\theta)}\mathcal{M}_{H}$. According to Hitchin, they induce the hyperkähler metric $g_{H,{\mathbb{R}}}$ and the Kähler metric $g_{H}$ of $\mathcal{M}_{H}$, respectively. (See §4.2.3 below for a more precise comparison with the Riemannian metric constructed in [14].) We have $g_{H,{\mathbb{R}}}=\mathop{\rm Re}\nolimits g_{H}$. The metrics $g_{H}$ and $g_{H,{\mathbb{R}}}$ are called the Hitchin metric.
4.2.2 Main Theorem
Let $s$ be the $C^{\infty}$-automorphism of the complex vector bundle $T\mathcal{M}_{H}^{\prime}$ determined by $g_{H}=g_{\mathop{\rm sf}\nolimits}\cdot s$. Similarly, let $s_{{\mathbb{R}}}$ be the $C^{\infty}$-automorphism of the real vector bundle $T\mathcal{M}_{H}^{\prime}$ determined by $g_{H,{\mathbb{R}}}=g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}\cdot s_{{\mathbb{R}}}$. We obtain the following theorem from Proposition 4.21, Proposition 4.31 and Proposition 4.32 below.
Theorem 4.7
For any compact subset $K\subset\mathcal{M}_{H}^{\prime}$, there exist positive constants $B_{i}>0$ $(i=1,2)$ such that $\bigl{|}(s-\mathop{\rm id}\nolimits)_{(E,t\theta)}\bigr{|}_{g_{\mathop{\rm sf}\nolimits}}\leq B_{1}\exp(-B_{2}t)$ and $\bigl{|}(s_{{\mathbb{R}}}-\mathop{\rm id}\nolimits)_{(E,t\theta)}\bigr{|}_{g_{\mathop{\rm sf}\nolimits,{\mathbb{R}}}}\leq B_{1}\exp(-B_{2}t)$ as $t\to\infty$ for any $(E,\theta)\in K$.
Though the statements in §4.34.4 are given for a fixed Higgs bundle $(E,\theta)$, the estimates can be given locally uniformly on $\mathcal{M}^{\prime}_{H}$.
4.2.3 Appendix
It is well known that $g_{H,{\mathbb{R}}}$ in §4.2.1 equals the Riemannian metric of $\mathcal{M}_{H}$ introduced in [14, §6]. We include an explanation for the convenience of readers. Let $(E,\theta,h)$ be a harmonic bundle. Let $\nabla_{h}=\overline{\partial}_{E}+\partial_{E,h}$ denote the Chern connection of $(E,h)$. Let $\theta^{\dagger}_{h}$ denote the adjoint of $\theta$ with respect to $h$. We obtain the following equations for $\tau^{1,0}+\tau^{0,1}\in A^{1}(\mathop{\rm End}\nolimits(E))$ as the linearization of the Hitchin equation for a pair of a holomorphic structure and a Higgs field:
$\partial_{E,h}\tau^{0,1}+[\theta^{\dagger}_{h},\tau^{1,0}]-\Bigl{(}\partial_{E,h}\tau^{0,1}+[\theta^{\dagger}_{h},\tau^{1,0}]\Bigr{)}^{\dagger}_{h}=0,$ (37)
¯$\overline{\partial}_{E}\tau^{1,0}+[\theta,\tau^{0,1}]=0.$ (38)
Let $\mathfrak{u}(E,h)\subset\mathop{\rm End}\nolimits(E)$ denote the subbundle of endomorphisms of $E$ which are anti-self-adjoint with respect to $h$. The linearization $d_{1}:A^{0}(\mathfrak{u}(E,h))\to A^{1}(\mathop{\rm End}\nolimits(E))$ of the action of the gauge group is given by
$d_{1}(a)=[\theta,a]+\overline{\partial}_{E}(a).$
We consider the real symmetric bilinear forms on $A^{0}(\mathfrak{u}(E,h))$ and $A^{1}(\mathop{\rm End}\nolimits(E))$ induced by $h$ and a conformal metric $g_{X}$ of $X$. The formal adjoint $d_{1}^{\ast}:A^{1}(\mathop{\rm End}\nolimits(E))\to A^{0}(\mathfrak{u}(E,h))$ is given by
$d_{1}^{\ast}(\tau)=-\sqrt{-1}\Lambda_{g_{X}}\Bigl{(}\partial_{E,h}\tau^{0,1}+[\theta^{\dagger}_{h},\tau^{1,0}]+\bigl{(}\partial_{E,h}\tau^{0,1}+[\theta^{\dagger}_{h},\tau^{1,0}]\bigr{)}^{\dagger}_{h}\Bigr{)}.$ (39)
The system of equations [14, (6.1)] is equivalent to the system of equations (37), (38) and $d_{1}^{\ast}\tau=0$. It is equivalent to the system of equations (38) and $(\partial_{E,h}+\mathop{\rm ad}\nolimits\theta^{\dagger}_{h})\tau=0$, i.e., the condition for $\tau$ to be a harmonic $1$-form. Because the (real) Hitchin metric at $(E,\theta)\in\mathcal{M}_{H}$ is defined as the restriction of the $L^{2}$-metric to the space of the solutions of [14, (6.1)], it equals $g_{H,{\mathbb{R}}|(E,\theta)}$.
4.3 Harmonic representatives of horizontal infinitesimal deformations
Let $\iota^{{\tt h}}_{t}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\longrightarrow H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$ denote the composite of the following morphisms:
$H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\stackrel{{\scriptstyle c_{1}}}{{\simeq}}H^{0}(\Sigma_{t\theta},K_{\Sigma_{t\theta}})\stackrel{{\scriptstyle c_{2}}}{{\longrightarrow}}H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta)).$
Here, $c_{1}$ is induced by $\Sigma_{\theta}\simeq\Sigma_{t\theta}$, and $c_{2}$ is the morphism $\iota^{{\tt h}}$ in (3.4.1) for $(E,t\theta)$. Let $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. Let us study the harmonic $1$-form ${\tt H}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that $[{\tt H}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$. Let $\|\nu\|$ denote the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$. Let $g_{X}$ denote a conformal metric of $X$.
4.3.1 Asymptotically harmonic representative on $X^{\circ}$
We recall $X^{\circ}=X\setminus D(\theta)$. For simplicity of the description, $D(\theta)$ is also denoted by $D$. We obtain $F_{\nu}\in H^{0}(X^{\circ},\mathop{\rm End}\nolimits(E)\otimes K_{X})$ as in §2.2.7. Let $N(D)$ be a neighbourhood of $D$.
Lemma 4.8
There exist positive constants $B_{i}$ $(i=0,1,2)$ such that the following holds on $X\setminus N(D)$ for any $t\geq 1$:
$|F_{\nu}|_{h_{t},g_{X}}\leq B_{0}\|\nu\|,\quad\quad\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}F_{\nu}\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\nu\|.$
Proof   We obtain the first from [20, Corollary 2.6], and the second from [20, Corollary 2.5].
Lemma 4.9
There exist positive constants $B_{i}$ $(i=1,2)$ such that the following holds:
$\left|\sqrt{-1}\int_{N(D)}h_{t}(F_{\nu},F_{\nu})-\sqrt{-1}\int_{\Sigma_{\theta|N(D)}}\nu\wedge\overline{\nu}\right|\\ \leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (40)
Proof   We have
$\sqrt{-1}\int_{\Sigma_{\theta|N(D)}}\nu\wedge\overline{\nu}=\sqrt{-1}\int_{N(D)}\mathop{\rm tr}\nolimits\bigl{(}F_{\nu}\cdot F_{\overline{\nu}}\bigr{)}.$
We also have
$\sqrt{-1}\int_{N(D)}h_{t}(F_{\nu},F_{\nu})=\sqrt{-1}\int_{N(D)}\mathop{\rm tr}\nolimits\bigl{(}F_{\nu}\cdot(F_{\nu})^{\dagger}_{h_{t}}\bigr{)}.$
By [20, Corollary 2.6], there exist $B_{i}>0$ $(i=3,4)$ such that
$\left|(F_{\nu}^{\dagger})_{h_{t}}-F_{\overline{\nu}}\right|_{h_{t}}\leq B_{3}\exp(-B_{4}t)\|\nu\|.$
Hence, we obtain (40).
We obtain the $1$-cocycle $(\mathop{\rm tr}\nolimits F_{\nu})\mathop{\rm id}\nolimits_{E}\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$.
Lemma 4.10
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{g_{X}}((\mathop{\rm tr}\nolimits F_{\nu})\mathop{\rm id}\nolimits_{E})=0$.
4.3.2 Asymptotically harmonic representatives around the discriminant
Let $U$ be an open subset of $X$ with a holomorphic coordinate $z$ such that the following holds.
  • The coordinate induces the holomorphic isomorphism $U\simeq\{|z|<2\}$.
  • $D(\theta)\cap U\subset\{|z|<1/4\}$.
  • Each connected component of $\Sigma_{\theta|U}$ is simply connected.
We fix a base point of each connected component of $\Sigma_{\theta|U}$. We set $(E_{U},\theta_{U})=(E,\theta)_{|U}$. Note that there exists a non-degenerate symmetric product $C_{U}$ of $(E_{U},\theta_{U})$ compatible with the limiting configuration $h_{\infty}$. Let $h_{U,t}$ be the harmonic metric of $(E_{U},t\theta_{U})$ compatible with $C_{U}$. There exists a holomorphic function $\alpha_{U}$ on $\Sigma_{\theta|U}$ such that $d\alpha_{U}=\nu_{|\Sigma_{\theta|U}}$ and that $\alpha_{U}$ is 00 at the base points of the connected components of $\Sigma_{\theta|U}$. We obtain $F_{\alpha_{U}}\in\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E_{U},\theta_{U};C_{U}))$ as in §2.2.7. We set
${\tt H}_{U}(\nu,t):=(\partial_{E_{U},h_{U,t}}+\mathop{\rm ad}\nolimits t(\theta_{U})^{\dagger}_{h_{U,t}})F_{\alpha_{U}}.$
By Proposition 2.30, we obtain the following.
Lemma 4.11
${\tt H}_{U}(\nu,t)$ is a harmonic $1$-form of $(\mathop{\rm End}\nolimits(E_{U}),t\mathop{\rm ad}\nolimits\theta_{U},h_{U,t})$ on $U$, i.e.,
(¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U}){\tt H}_{U}(\nu,t)=0,\quad(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U}(\nu,t)=0.$
Also, ${\tt H}_{U}(\nu,t)$ is contained in $A^{1}(\mathop{\rm Def}\nolimits(E_{U},\theta_{U};C_{U}))$. Moreover, we have $\rho({\tt H}_{U}(\nu,t))=\pi_{\ast}(\nu)_{|U}$, which implies $\mathop{\rm tr}\nolimits({\tt H}_{U}(\nu,t))=\mathop{\rm tr}\nolimits\pi_{\ast}(\nu)_{|U}$. (See §2.2.4 for $\mathop{\rm tr}\nolimits\pi_{\ast}(\nu)$.)
For $0<r\leq 2$, we set $U(r):=\{|z|<r\}$.
Lemma 4.12
For any $21/4<r_{2}<r_{1}<$ 2, there exist positive constants $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ for any $t\geq 1$:
$|F_{\alpha_{U}}|\leq B_{0}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)-F_{\nu}|_{h_{U,t}}\leq B_{1}e^{-B_{2}t}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)|_{h_{U,t}}\leq B_{3}\|\nu\|.$
Proof   We obtain the first from [20, Corollary 2.6], the second from [20, Corollary 2.5, Proposition 2.10], and the third from Lemma 4.8.
Lemma 4.13
There exits a unique element $\sigma_{U}(\nu,t)\in A^{0}(U\cap X^{\circ},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C_{U}))$ such that
(¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})\sigma_{U}(\nu,t)={\tt H}_{U}(\nu,t)-F_{\nu}.$
For any $21/4<r_{2}<r_{1}<$ 2, there exist $B_{i}>0$ $(i=1,2)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ such that
$|\sigma_{U}(\nu,t)|_{h_{U,t}}\leq B_{1}e^{-B_{2}t}\|\nu\|.$
Proof   We obtain the first claim from Proposition 2.12. The second follows from Lemma 4.12.
Remark 4.14
Constants $B_{i}$ of the lemmas in this section are not necessarily the same.
For any $20<r<$ 2, let $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}$ denote the $L^{2}$-norm of ${\tt H}_{U}(\nu,t)$ on $U(r)$ with respect to $h_{U,t}$. Note that it is independent of the choice of a conformal metric of $X$.
Lemma 4.15
For any $20<r<$ 2, there exists $B>0$ such that $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}\leq B\|\nu\|$ for any $t\geq 1$.
Proof   We note the following equality:
${\tt H}_{U}(\nu,t)=(\partial_{E_{U},h_{U,t}}+\mathop{\rm ad}\nolimits t(\theta_{U})^{\dagger}_{h_{U,t}})F_{\alpha_{U}}=(\overline{\partial}_{E_{U}}+\partial_{E_{U},h_{U,t}}+\mathop{\rm ad}\nolimits t\theta_{U}+\mathop{\rm ad}\nolimits t(\theta_{U})^{\dagger}_{h_{U,t}})F_{\alpha_{U}}.$
By Proposition 2.27, we obtain the following equality:
$\sqrt{-1}\int_{U(r)}\Bigl{(}h_{U,t}({\tt H}_{U}(\nu,t)^{1,0},{\tt H}_{U}(\nu,t)^{1,0})-h_{U,t}({\tt H}_{U}(\nu,t)^{0,1},{\tt H}_{U}(\nu,t)^{0,1})\Bigr{)}=\\ \sqrt{-1}\int_{\partial U(r)}\Bigl{(}h_{U,t}(F_{\alpha_{U}},{\tt H}_{U}(\nu,t)^{1,0})-h_{U,t}(F_{\alpha_{U}},{\tt H}_{U}(\nu,t)^{0,1})\Bigr{)}.$ (41)
We obtain the claim of Lemma 4.15 from Lemma 4.12.
Lemma 4.16
For any $20<r<$ 2, there exists $B^{\prime}>0$ such that $|{\tt H}_{U}(\nu,t)|_{h_{U,t},g_{X}}\leq B^{\prime}t\|\nu\|$ on $U(r)$ for any $t\geq 1$.
Proof   We set $Y(t)=\{|w|<2t\}\subset{\mathbb{C}}$. We define the isomorphism $\varphi_{t}:Y(t)\to U$ by $\varphi_{t}(w)=t^{-1}w$. We obtain the harmonic bundles $(\widetilde{E}_{U,t},\widetilde{\theta}_{U,t},\widetilde{h}_{U,t})=\varphi_{t}^{\ast}(E_{U},t\theta_{U},h_{U,t})$ on $Y(t)$. Let $f_{U,t}$ be the endomorphism of $\widetilde{E}_{U,t}$ defined by $\widetilde{\theta}_{U,t}=f_{U,t}\,dw$. Then, there exists $B_{1}>0$ such that any eigenvalue $\alpha$ of $f_{U,t|Q}$ satisfy $|\alpha|\leq B_{1}$ for any $t\geq 1$ and any $Q\in Y(t)$. Let $10<r_{1}<r_{2}<$ 1. By Lemma 4.15, we have
$\|\varphi_{t}^{-1}{\tt H}_{U}(\nu,t)_{|Y(r_{2}t)}\|_{L^{2},\widetilde{h}_{U,t}}\leq B\|\nu\|.$
Let $g_{0}=dw\,d\overline{w}$ be the Euclidean metric of $Y(t)$. By Proposition 2.35, there exists $B_{2}>0$ such that
$|\varphi_{t}^{-1}{\tt H}_{U}(\nu,t)|_{\widetilde{h}_{U,t},g_{0}}\leq B_{2}\|\nu\|$
on $Y(r_{1}t)$. It implies the claim of the lemma.
Corollary 4.17
There exist $B_{1},B_{2}>0$ such that the following holds for any $t\geq 1$:
|(¯$\bigl{|}(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{t},g_{X}}{\tt H}_{U,t}\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\nu\|.$ (42)
Proof   As remarked in Lemma 4.11, we have (¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U,t}=0$. By the estimate $h_{U,t}-h_{t}=O(e^{-\beta t})$ with respect to $h_{U,t}$ for some $\beta>0$ in [23, Theorem 7.17], we obtain (42) from Lemma 4.16.
Lemma 4.18
There exist $B_{i}>0$ $(i=1,2)$ such that the following holds.
$\left|\sqrt{-1}\int_{U(1)}\Bigl{(}h_{t}({\tt H}_{U}(\nu,t)^{1,0},{\tt H}_{U}(\nu,t)^{1,0})-h_{t}({\tt H}_{U}(\nu,t)^{0,1},{\tt H}_{U}(\nu,t)^{0,1})\Bigr{)}-\sqrt{-1}\int_{\Sigma_{\theta|U(1)}}\nu\wedge\overline{\nu}\right|\\ \leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (43)
Proof   By [23, Theorem 7.17], there exists $\beta>0$ such that $h_{U,t}-h_{t}=O(e^{-\beta t})$ on $U(1)$ as $t\to\infty$. Hence, by Lemma 4.15, there exist $B_{i}>0$ $(i=3,4)$ such that the following holds:
$\left|\sqrt{-1}\int_{U(1)}h_{t}\bigl{(}{\tt H}_{U}(\nu,t)^{p,q},{\tt H}_{U}(\nu,t)^{p,q}\bigr{)}-\sqrt{-1}\int_{U(1)}h_{U,t}\bigl{(}{\tt H}_{U}(\nu,t)^{p,q},{\tt H}_{U}(\nu,t)^{p,q}\bigr{)}\right|\\ \leq B_{3}\exp(-B_{4}t)\|\nu\|^{2}.$
Recall (41). By Lemma 4.12, there exist $B_{i}>0$ $(i=5,6)$ such that
$\left|\sqrt{-1}\int_{\partial U(1)}\Bigl{(}h_{U,t}(F_{\alpha_{U}},{\tt H}_{U}(\nu,t)^{0,1})\Bigr{)}\right|\leq B_{5}\exp(-B_{6}t)\|\nu\|^{2}.$
We have
$\sqrt{-1}\int_{\partial U(1)}h_{U,t}(F_{\alpha_{U}},{\tt H}_{U}(\nu,t)^{1,0})=\sqrt{-1}\int_{\partial U(1)}\mathop{\rm tr}\nolimits\bigl{(}F_{\alpha_{U}}\cdot({\tt H}_{U}(\nu,t)^{1,0})^{\dagger}_{h_{U,t}}\bigr{)}.$
We also have
$\sqrt{-1}\int_{\Sigma_{\theta|U(1)}}\nu\wedge\overline{\nu}=\sqrt{-1}\int_{\partial\Sigma_{\theta|U(1)}}\alpha_{U}\wedge\overline{\nu}=\sqrt{-1}\int_{\partial U(1)}\mathop{\rm tr}\nolimits\bigl{(}F_{\alpha_{U}}\cdot F_{\overline{\nu}}\bigr{)}.$
By [20, Corollary 2.6] and Lemma 4.12, there exist $B_{i}>0$ $(i=7,8)$ such that the following holds on a neighbourhood of $\partial U(1)$:
$\bigl{|}({\tt H}_{U}(\nu,t)^{1,0})^{\dagger}_{h_{U,t}}-F_{\overline{\nu}}\bigr{|}_{h_{U,t}}\leq B_{7}\exp(-B_{8}t)\|\nu\|.$
Then, we obtain (43).
4.3.3 Asymptotically harmonic representatives on $X$
Let $U_{j}$ $(j=1,\ldots,m)$ be open subsets of $X$ equipped with a holomorphic coordinate $z_{j}$ satisfying the conditions in §4.3.2, such that $D(\theta)\subset\bigcup_{j=1}^{m}U_{j}$ and that $U_{j}\cap U_{k}=\emptyset$ $(j\neq k)$. We set $r_{1}=1$ and $r_{2}=\frac{1}{2}$. Let $\chi_{j}:X\to[0,1]$ be a $C^{\infty}$-function such that $\chi_{j}=1$ on $U_{j}(r_{2})$ and $\chi_{j}=0$ on $X\setminus U_{j}(r_{1})$. There exists a closed $1$-form ${\tt H}^{\prime}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that the following holds.
  • On $X\setminus\bigcup_{j=1}^{m}U_{j}(r_{1})$, we have ${\tt H}^{\prime}(\nu,t)=F_{\nu}$.
  • On $U_{j}(r_{1})\setminus\overline{U}_{j}(r_{2})$, we have ${\tt H}^{\prime}(\nu,t)=F_{\nu}+(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\bigl{(}\chi_{j}\cdot\sigma_{U_{j}}(\nu,t)\bigr{)}$.
  • On $U_{j}(r_{2})$, we have ${\tt H}^{\prime}(\nu,t)={\tt H}_{U_{j}}(\nu,t)$.
Lemma 4.19
  • (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt H}^{\prime}(\nu,t)=0$, and $\rho_{t}({\tt H}^{\prime}(\nu,t))=\pi_{\ast}(\nu)$. In particular, $\mathop{\rm tr}\nolimits({\tt H}^{\prime}(\nu,t))=\mathop{\rm tr}\nolimits\pi_{\ast}(\nu)$. Here, $\rho_{t}$ denotes the morphism $\rho$ (see Lemma 2.11) for $(E,t\theta)$.
  • There exist $B_{i}>0$ $(i=1,2)$ such that $|{\tt H}^{\prime}(\nu,t)-F_{\nu}|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)$ on $X\setminus\bigcup U_{j}(r_{2})$ and that $|{\tt H}^{\prime}(\nu,t)-{\tt H}_{U_{j}}(\nu,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)$ on $U_{j}(r_{1})$.
  • There exists $C>0$ such that $|{\tt H}^{\prime}(\nu,t)|_{h_{t},g_{X}}\leq Ct\|\nu\|$ and $\|{\tt H}^{\prime}(\nu,t)\|_{L^{2},h_{t}}\leq C\|\nu\|$ for any $t\geq 1$.
  • There exist $B_{i}>0$ $(i=3,4)$ such that |(¯$\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)_{h_{t},g_{X}}^{\ast}{\tt H}^{\prime}(\nu,t)\bigr{|}_{h_{t}}\leq B_{3}\exp(-B_{4}t)$ on $X$ for any $t\geq 1$.
Proof   We obtain the first two claims by the construction. We obtain the third claim from the estimates for $F_{\nu}$ and ${\tt H}_{\nu}$. The fourth claim follows from Lemma 4.8 and Corollary 4.17.
4.3.4 Harmonic representatives
We note that $\mathop{\rm tr}\nolimits\Bigl{(}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}({\tt H}^{\prime}(\nu,t))\Bigr{)}=0$. There exists a unique $\gamma(\nu,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ such that $\mathop{\rm tr}\nolimits\gamma(\nu,t)=0$ and
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
Lemma 4.20
There exist $B_{i}>0$ $(i=1,2)$ such that the following holds on $X$ for any $t\geq 1$:
$|\gamma(\nu,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\nu\|.$ (44)
Proof   By Proposition 2.42 and Lemma 4.19, there exist $B_{i}>0$ $(i=11,12)$ such that the following holds for any $t\geq 1$:
$\|\gamma(\nu,t)\|_{L^{2},h_{t},g_{X}}\leq B_{11}\exp(-B_{12}t)\|\nu\|.$
Let $P$ be any point of $X$, which is not necessarily contained in $D(\theta)$. Let $(X_{P},z_{P})$ be a holomorphic coordinate neighbourhood around $P$ such that $z(P)=0$ and that $X_{P}\simeq\{|z_{P}|<2\}$. We set $Y(t)=\bigl{\{}w\in{\mathbb{C}}\,\big{|}\,|w|<2t\bigr{\}}$. For any $t\geq 1$, let $\varphi_{t}:Y(t)\to X_{P}$ be defined by $\varphi_{t}(w)=t^{-1}w$. We obtain the harmonic bundle $\varphi_{t}^{\ast}(E,t\theta,h_{t})$ on $Y(t)$. By applying Proposition 2.36 to the section $\varphi_{t}^{\ast}\gamma(\nu,t)$, we obtain (44).
We set
${\tt H}(\nu,t)={\tt H}^{\prime}(\nu,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t).$
Then, ${\tt H}(\nu,t)$ is a harmonic $1$-form satisfying $[{\tt H}(\nu,t)]=[{\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}(\nu)$.
4.3.5 Estimate of the pairings
Proposition 4.21
There exist $B_{i}>0$ $(i=1,2)$ such that
$\Bigl{|}g_{H}(\iota^{{\tt h}}_{t}(\nu),\iota^{{\tt h}}_{t}(\nu))-g_{\mathop{\rm sf}\nolimits}(\iota^{{\tt h}}_{t}(\nu),\iota^{{\tt h}}_{t}(\nu))\Bigr{|}\leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (45)
Proof   The estimate (45) is equivalent to
$\left|2\sqrt{-1}\int_{X}\Bigl{(}h_{t}({\tt H}(\nu,t)^{1,0},{\tt H}(\nu,t)^{1,0})-h_{t}({\tt H}(\nu,t)^{0,1},{\tt H}(\nu,t)^{0,1})\Bigr{)}-2\sqrt{-1}\int_{\Sigma_{\theta}}\nu\overline{\nu}\right|\leq B_{1}\exp(-B_{2}t)\|\nu\|^{2}.$ (46)
We set $X_{1}=X\setminus\bigcup_{j=1}^{m}U_{j}(1)$. By Lemma 4.20, there exist $B_{i}>0$ $(i=3,4)$ such that the following holds:
$\sup_{X_{1}}\bigl{|}{\tt H}(\nu,t)-F_{\nu}\bigr{|}_{h_{t},g_{X}}+\sum_{j=1}^{m}\sup_{U_{j}(1)}\bigl{|}{\tt H}(\nu,t)-{\tt H}_{U_{j}}(\nu,t)\bigr{|}_{h_{t},g_{X}}\leq B_{3}\exp(-B_{4}t)\|\nu\|.$
Then, we obtain (46) from Lemma 4.9 and Lemma 4.18.
4.4 Harmonic representatives of vertical infinitesimal deformations
Let $\iota^{{\tt v}}_{t}:H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$ denote the composite of the following morphisms:
$H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\stackrel{{\scriptstyle c_{1}}}{{\simeq}}H^{1}(\Sigma_{t\theta},\mathcal{O}_{\Sigma_{t\theta}})\stackrel{{\scriptstyle c_{2}}}{{\longrightarrow}}H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta)).$
Here, $c_{1}$ is induced by the isomorphism $\Sigma_{\theta}\simeq\Sigma_{t\theta}$, and $c_{2}$ is induced by $\pi_{t\ast}\mathcal{O}_{\Sigma_{t\theta}}=\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E,t\theta))$.
Let $\tau$ be a harmonic $(0,1)$-form on $\Sigma_{\theta}$, i.e., $\tau$ is a $(0,1)$-form such that $\partial_{\Sigma_{\theta}}\tau=0$. It induces $[\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. Conversely, any element of $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ can be expressed uniquely in this way. We shall study the harmonic representative of ${\tt V}(\tau,t)$ of $\iota_{t}^{{\tt v}}([\tau])$. Let $\|\tau\|$ denote the $L^{2}$-norm of $\tau$ on $\Sigma_{\theta}$.
4.4.1 Preliminary
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exist the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
We have the holomorphic vector bundles $E^{\dagger}_{t}=(E,\partial_{E,h_{t}})$ with the Higgs field $\theta_{h_{t}}^{\dagger}$ on $X^{\dagger}$. The spectral curve of $(E,t\theta_{h_{t}}^{\dagger})$ equals $(\Sigma_{t\theta})^{\dagger}\subset(T^{\ast}X)^{\dagger}=T^{\ast}(X^{\dagger})$. Let $\pi^{\dagger}_{t}:(\Sigma_{t\theta})^{\dagger}\to X^{\dagger}$ denote the projection. There exists a holomorphic line bundle $L^{\dagger}_{t}$ on $(\Sigma_{t\theta})^{\dagger}$ with an isomorphism $(\pi_{t}^{\dagger})_{\ast}L^{\dagger}_{t}\simeq E^{\dagger}_{t}$.
There exist the following natural morphisms:
$\mathop{\rm tr}\nolimits:(\pi^{\dagger}_{t})_{\ast}\mathcal{O}_{\Sigma_{t\theta}^{\dagger}}=(\pi^{\dagger}_{1})_{\ast}\mathcal{O}_{\Sigma_{\theta}^{\dagger}}\longrightarrow\mathcal{O}_{X^{\dagger}},\quad\quad\mathop{\rm tr}\nolimits:(\pi^{\dagger}_{t})_{\ast}K_{\Sigma_{t\theta}^{\dagger}}=(\pi^{\dagger}_{1})_{\ast}K_{\Sigma_{\theta}^{\dagger}}\longrightarrow K_{X^{\dagger}}.$
4.4.2 Asymptotically harmonic representatives on $X^{\circ}$
We obtain $F_{\tau}\in A^{0,1}(X^{\circ},\mathop{\rm End}\nolimits(E))$ as in §2.2.7. For simplicity of the description, $D(\theta)$ is also denoted by $D$. Let $N(D)$ be a neighbourhood of $D$.
Lemma 4.22
There exist positive constants $B_{i}$ $(i=0,1,2)$ such that the following holds on $X\setminus N(D)$ for any $t\geq 1$:
$|F_{\tau}|_{h_{t},g_{X}}\leq B_{0}\|\tau\|,\quad\quad\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}F_{\tau}\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\tau\|.$
Proof   We obtain the first from [20, Corollary 2.6], and the second from [20, Proposition 2.10].
The following lemma is obvious by the construction.
Lemma 4.23
We have $\mathop{\rm tr}\nolimits F_{\tau}=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits F_{\tau}=0$.
4.4.3 Asymptotically harmonic $1$-forms around the discriminant
Let $(U,z)$ be as in §4.3.2. We choose a base point of each connected component of $\Sigma_{\theta|U}$. There exists an anti-holomorphic function $\beta_{U}$ such that ¯$\overline{\partial}\beta_{U}=\tau$ on $\Sigma_{\theta|U}$ and that $\beta_{U}=0$ at the base points of the connected components.
We may regard $\beta_{U}$ as a holomorphic function on $\Sigma_{\theta|X_{P}}^{\dagger}$. We obtain the endomorphism $F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{U}}$ of $E^{\dagger}_{t}$. It satisfies
$(\partial_{E,h_{t}}+\mathop{\rm ad}\nolimits t\theta^{\dagger}_{h_{t}})F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{U}}=0.$
As in Proposition 2.30, we obtain the following harmonic $1$-form of $(E_{t},t\theta^{\dagger}_{h_{t}},h_{t})$ on $U^{\dagger}$:
${\tt V}_{U}(\tau,t):=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{U}}.$
By Corollary 2.25, ${\tt V}_{U}(\tau,t)$ is also a harmonic $1$-form of $(E,t\theta,h_{t})$ on $U$. In particular, the following holds on $U$:
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}_{U}(\tau,t)=0,\quad\quad(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}_{U}(\tau,t)=0.$
Lemma 4.24
For any $21/4<r_{2}<r_{1}<$ 2, there exist $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$:
$\bigl{|}F_{(E^{\dagger}_{t},t\theta^{\dagger}_{h_{t}}),\beta_{U}}\bigr{|}_{h_{t}}\leq B_{0}\|\tau\|,\quad\bigl{|}{\tt V}_{U}(\tau,t)-F_{\tau}\bigr{|}_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|,\quad\bigl{|}{\tt V}_{U}(\tau,t)\bigr{|}_{h_{t},g_{X}}\leq B_{3}\|\tau\|.$
Proof   We obtain the first from [20, Corollary 2.6], the second from [20, Corollary 2.5, Corollary 2.6, Proposition 2.10], and the third from Lemma 4.22.
For any $20<r<$ 2, let $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of ${\tt V}_{U}(\tau,t)$ on $U(r)$ with respect to $h_{t}$. We obtain the following lemma by using Proposition 2.27 and Lemma 4.24, as in the proof of Lemma 4.15.
Lemma 4.25
For any $20<r<$ 2, there exists $B>0$ such that $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}\leq B\|\tau\|$ for any $t\geq 1$.
The following lemma is similar to Lemma 4.16.
Lemma 4.26
For any $20<r<$ 2, there exists $B^{\prime}>0$ such that $|{\tt V}_{U}(\tau,t)|_{h_{t},g_{X}}\leq B^{\prime}t\|\tau\|$ on $U(r)$ for any $t\geq 1$.
By Proposition 2.30, we obtain the following lemma.
Lemma 4.27
$\mathop{\rm tr}\nolimits({\tt V}_{U}(\tau,t))=\mathop{\rm tr}\nolimits(\pi_{1}^{\dagger})_{\ast}(\tau)$.
Remark 4.28
From the holomorphic function $\overline{\beta}_{U}$ on $\Sigma_{\theta|U}$, we obtain the holomorphic endomorphism $F_{\overline{\beta}_{U}}$ of $(E,t\theta)_{|U}$. We have $(F_{\overline{\beta}_{U}})^{\dagger}_{h_{t}}=F_{(E^{\dagger}_{t},t\theta^{\dagger}_{t}),\beta_{U}}$.
4.4.4 Asymptotically harmonic representatives on $X$
Let $(U_{j},z_{j})$ and $\chi_{j}$ $(j=1,\ldots,m)$ be as in §4.3.3. Let $\tau^{\prime}$ be the $(0,1)$-form on $\Sigma_{\theta}$ defined as follows:
$\tau^{\prime}:=\tau-\sum_{j}\overline{\partial}\bigl{(}\pi^{\ast}(\chi_{j})\cdot\beta_{U_{j}}\bigr{)}.$
We obtain a $1$-cocycle $F_{\tau^{\prime}}\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that $[F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}([\tau])$. We set
${\tt V}^{\prime}(\tau,t)=F_{\tau^{\prime}}+\sum_{j=1}^{m}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\Bigl{(}\chi_{j}\cdot F_{(E^{\dagger}_{h_{t}},t\theta^{\dagger}_{h_{t}}),\beta_{U_{j}}}\Bigr{)}.$
By the construction, the following holds.
Lemma 4.29
  • (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}^{\prime}(\tau,t)=0$ and $[{\tt V}^{\prime}(\tau,t)]=\iota^{{\tt v}}_{t}([\tau])$.
  • We have $\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=0$.
  • There exists $B_{i}>0$ $(i=1,2)$ such that $|{\tt V}^{\prime}(\tau,t)-F_{\tau}|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $X\setminus\bigcup U_{j}(r_{2})$ and that $|{\tt V}^{\prime}(\tau,t)-{\tt V}_{U_{j}}(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $U_{j}(r_{1})$.
  • There exists $B>0$, which is independent of $t$, such that $|{\tt V}^{\prime}(\tau,t)|_{h_{t},g_{X}}\leq Bt\|\tau\|$ and that $\|{\tt V}^{\prime}(\tau,t)\|_{L^{2},h_{t}}\leq B\|\tau\|$.
  • There exist $B_{i}>0$ $(i=1,2)$ such that |(¯$\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$.
4.4.5 Harmonic representatives
There uniquely exists $\gamma(\tau,t)\in A^{0}(\mathop{\rm End}\nolimits(E))$ satisfying $\mathop{\rm tr}\nolimits\gamma(\tau,t)=0$ and
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t).$
The following lemma is similar to Lemma 4.20.
Lemma 4.30
There exists $B_{i}$ $(i=1,2)$ such that the following holds on $X$ for any $t\geq 1$:
$|\gamma(\tau,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|.$ (47)
We set
${\tt V}(\tau,t):={\tt V}^{\prime}(\tau,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\tau,t).$
Then, ${\tt V}(\tau,t)$ is a harmonic $1$-form satisfying $[{\tt V}(\tau,t)]=\iota_{t}^{{\tt v}}([\tau])$.
4.4.6 Estimate of the pairings
Proposition 4.31
There exists $B_{i}$ $(i=1,2)$ such that
$\Bigl{|}g_{H}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt v}}_{t}(\tau))-g_{\mathop{\rm sf}\nolimits}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt v}}_{t}(\tau))\Bigr{|}\leq B_{1}\exp(-B_{2}t)\|\tau\|^{2}.$ (48)
Proof   The estimate (48) is equivalent to the following:
$\left|2\sqrt{-1}\int_{X}\Bigl{(}h_{t}({\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0})-h_{t}({\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1})\Bigr{)}-2(-\sqrt{-1})\int_{\Sigma_{\theta}}\tau\overline{\tau}\right|\leq B_{1}\exp(-B_{2}t)\|\tau\|^{2}.$ (49)
Because $\partial\tau=0$, we have
$-2\sqrt{-1}\int_{\Sigma_{\theta}}\tau\overline{\tau}=-2\sqrt{-1}\int_{\Sigma_{\theta}}\tau^{\prime}\overline{\tau}.$
Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}(\tau,t)=0$, we have
$2\sqrt{-1}\int_{X}\Bigl{(}h_{t}({\tt V}(\tau,t)^{1,0},{\tt V}(\tau,t)^{1,0})-h_{t}({\tt V}(\tau,t)^{0,1},{\tt V}(\tau,t)^{0,1})\Bigr{)}=-2\sqrt{-1}\int_{X}h_{t}(F_{\tau^{\prime}},{\tt V}(\tau,t)^{0,1})\\ =-2\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl{(}F_{\tau^{\prime}}\cdot({\tt V}(\tau,t)^{0,1})^{\dagger}_{h_{t}}\bigr{)}.$ (50)
Then, we obtain (49) from Lemma 4.29 and Lemma 4.30.
Proposition 4.32
There exist $B_{1},B_{2}>0$ such that the following holds for any $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ and $\tau\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$:
$\bigl{|}g_{H}(\iota^{{\tt v}}_{t}(\tau),\iota^{{\tt h}}_{t}(\nu))\bigr{|}\leq B_{1}\exp(-B_{2}t)\|\nu\|\|\tau\|.$ (51)
Proof   The estimate (51) is equivalent to the following:
$\left|2\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt H}(\nu,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt H}(\nu,t)^{0,1}\bigr{)}\Bigr{)}\right|\leq B_{1}\exp(-B_{2}t)\|\nu\|\|\tau\|.$ (52)
Because (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}(\nu,t)=0$, we have
$2\sqrt{-1}\int_{X}\Bigl{(}h_{t}\bigl{(}{\tt V}(\tau,t)^{1,0},{\tt H}(\nu,t)^{1,0}\bigr{)}-h_{t}\bigl{(}{\tt V}(\tau,t)^{0,1},{\tt H}(\nu,t)^{0,1}\bigr{)}\Bigr{)}=2\sqrt{-1}\int_{X}h_{t}(F_{\tau^{\prime}},{\tt H}(\nu,t)^{0,1})\\ =2\sqrt{-1}\int_{X}\mathop{\rm tr}\nolimits\bigl{(}F_{\tau^{\prime}}\cdot({\tt H}(\nu,t)^{0,1})^{\dagger}_{h_{t}}\bigr{)}.$ (53)
We obtain (52) from Lemma 4.19 and Lemma 4.20.